### v1

# Diffusion approximation of the stationary distribution of a two-level single server queue

Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study in the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation. It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [1, 2, 11, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
Keywords: single server queue, level dependent arrival and service, stationary distribution, heavy traffic, diffusion approximation, Palm distribution, process limit, stochastic integral equation.
11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan & School of Data Science, Chinese University of Hong Kong, Shenzhen, China.
1 Introduction
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length. Namely, customers arrive subject to a renewal process as long as its queue length (including a customer being served) is less than threshold $\ell_{0}$, while it is changed to another renewal process otherwise. The service rate is changed similarly but in a slightly different way. That is, the service rate is constant when the queue length is not greater than the threshold, while it is another constant otherwise. We assume that customers are served in the FCFS manner, and the amounts of the work which are brought by them are independent and identically distributed and independent of the arrival process. We refer to this model as a 2-level $GI/G/1$ queue.
This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service rate. Here, the system is an internet server, and its energy consumption depends on service rate. Then, the problem is to find an optimal level to minimize the operating energy of the server and customer’s waiting cost.
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. It may be considered to approximate using a quasi-birth-and-death process, and it is not hard to numerically compute its performance. However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential. Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of this 2-level $GI/G/1$ queues under a certain scaling in heavy traffic. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
There is another motivation of this study. In general, theoretical study of diffusion approximation is hard for state-dependent queues. For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see Mandelbaum and Pats [9], Yamada [13]), while the arrival process and service rate discontinuously change in the present model. Thus, theoretical study for its diffusion approximation is challenging.
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue (e.g., see Kingman [8]). Furthermore, Miyazawa [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture is easily verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [1, 2, 11, 12]. This verification, Theorem 3.1, requires a certain limitation on system parameters or the exponentially distributed workload brought by a customer. The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
In this paper, we only study the limit of the scaled stationary distribution, but, as our results are related to the diffusion approximation, it is interesting to identify a reflected Brownian motion as a process limit for the 2-level $GI/G/1$ queue. This is a hard question because this Brownian motion discontinuously changes its drift and variance at the fixed state. We intuitively discuss what can be expected about this process limit in Section 5.1.
This paper is made up by five sections. We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1. A main result, Theorem 3.1, and the exponentially distributed case are presented in Section 3.2. In Section 4, we prove Theorem 3.1. Finally, in Section 5, we intuitively discuss two topics. In Section 5.1, we consider a process-limit for the 2-level $GI/G/1$ queue in heavy traffic. In Section 5.2, we intuitively verify the conjecture in Remark 3.1.
2 Framework for the 2-level GI/G/1 queues
We take the stochastic framework introduced in [12] to study the 2-level $GI/G/1$ queue. Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining service time, respectively, at time $t$. Define
$X(\cdot)=\{(L(t),R_{e}(t),R_{s}(t));t\geq 0\},$
Then, it is easy to see that $X(\cdot)$ is a Markov process with state space +×$\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and the 2-level $GI/G/1$ queue is described by $X(\cdot)$, where $\mathbb{Z}_{+}$ is the set of nonnegative integers and $\mathbb{R}_{+}=[0,\infty)$. It can be assumed that $X(t)$ is right-continuous and has left-limits at any point of time. Let $N_{e}(\cdot)\equiv\{N_{e}(t);t\geq 0\}$ and $N_{s}(\cdot)\equiv\{N_{s}(t);t\geq 0\}$ be the counting processes generated by the vanishing instants of $R_{e}(t)$ and $R_{s}(t)$, respectively. Note that those counting processes can be viewed as point processes which are integer-valued random measures on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ whose atoms are the counting instants of $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where ${\mathcal{B}}(\mathbb{R}_{+})$ is the Borel field on $\mathbb{R}_{+}$. We denote those point processes by $N_{e}$ and $N_{s}$ respectively.
According to Miyazawa [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
(2.a) Measurability: $A\in{\mathcal{F}}_{t}\;\mbox{ implies }\;\Theta_{s}^{-1}A\in{\mathcal{F}}_{s+t},\qquad s,t\geq 0,$
(2.b) Consistency: $X(t)\circ\Theta_{s}(\omega)=X(s+t)(\omega),\qquad s,t\geq 0,\omega\in\Omega,$
$N(t)\circ\Theta_{s}(\omega)=N(s+t)(\omega)-N(s)(\omega),\quad s,t\geq 0,\omega\in\Omega,$
for N()=Ne() and $\mbox{for $N(\cdot)=N_{e}(\cdot)$ and $N(\cdot)=N_{s}(\cdot)$},$
where $W\circ\Theta_{s}(\omega)=W(\Theta_{s}(\omega))$ for random variable $W$, and $\Theta_{s}^{-1}A=\{\omega\in\Omega;\Theta_{s}(\omega)\in A\}$.
3 A sequence of the 2-level GI/G/1 queues
We consider the scaling limit of the stationary distribution of the 2-level $GI/G/1$ queue in heavy traffic. To this end, we introduce a sequence of this queueing models indexed by $r\in(0,1]$. Since we consider their countable sequence, one may consider $r=1/m$ for positive integer $m$, but we keep $r$ just for convenience.
3.1 Assumptions
For the $r$-th model, let $\ell^{(r)}_{0}$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$. We refer to this number as the queue length. Let $t^{(r)}_{e,n}$ be the $n$-th arrival time of a customer, and let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customer. We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$. Thus, the inter-arrival times are independent, but they have may two different distributions depending on the queue length. Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are random variables which may have different distributions. As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers. We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable. We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$. Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively. Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
$\frac{d}{dt}R^{(r)}_{s}(t)=\left\{\begin{array}[]{ll}-c^{(r)}_{1},&L^{(r)}(t)\leq\ell^{(r)}_{0},\\ -c^{(r)}_{2},&L^{(r)}(t)>\ell^{(r)}_{0}.\end{array}\right.$
These dynamics is depicted in the sample path of $L^{(r)}(t)$ below.
We will use the following notations.
$m^{(r)}_{e,i}=\mathbb{E}(T^{(r)}_{e,i}),\qquad\lambda^{(r)}_{i}=1/m^{(r)}_{e,i},\qquad\sigma^{(r)}_{e,i}=\mathbb{E}[(T^{(r)}_{e,i}-m^{(r)}_{e,i})^{2}],\qquad i=1,2,$
$m^{(r)}_{s}=\mathbb{E}(T^{(r)}_{s}),\qquad\mu^{(r)}=1/m^{(r)}_{s},\qquad\sigma^{(r)}_{s}=\mathbb{E}[(T^{(r)}_{s}-m^{(r)}_{s})^{2}],$
$\rho^{(r)}_{i}=\lambda^{(r)}_{i}/(c^{(r)}_{i}\mu^{(r)}),\qquad i=1,2.$
For the $r$-th model, at time $t$, let $R^{(r)}_{e}(t)$ be the remaining arrival time, and let $R^{(r)}_{s}(t)$ be the remaining workload of a customer being served. Define $X^{(r)}(\cdot)\equiv\{X^{(r)}(t);t\geq 0\}$ by
$X^{(r)}(t)=(L^{(r)}(t),R^{(r)}_{e}(t),R^{(r)}_{s}(t)),$
and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively. Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov. Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with respect to the time-shift operator $\Psi$. We assume the following assumptions.
  • (3.a)
    $\{\mathbb{E}[(T^{(r)}_{e,i})^{3}];r\in(0,1]\}$ for $i=1,2$ and $\{\mathbb{E}[(T^{(r)}_{s})^{3}];r\in(0,1]\}$ are bounded.
  • (3.b)
    For $i=1,2$, $m^{(r)}_{e,i},\lambda^{(r)}_{i},\sigma^{(r)}_{e,i},c^{(r)}_{i}$ converge to $m_{e,i},\lambda_{i},\sigma_{e,i},c_{i}>0$, respectively, as $r\downarrow 0$. Similarly, $m^{(r)}_{s},\mu^{(r)},\sigma^{(r)}_{s}$ converge to $m_{s},\mu,\sigma_{s}$, respectively, as $r\downarrow 0$.
  • (3.c)
    $c^{(r)}_{i}\mu^{(r)}-\lambda^{(r)}_{i}=r\mu^{(r)}b_{i}+o(r)$ as $r\downarrow 0$, and $b_{2}>0$.
  • (3.d)
    $r\ell^{(r)}_{0}=\ell_{0}+o(r)$, and $\ell^{(r)}_{0}\geq 2$.
  • (3.e)
    $\rho^{(r)}_{2}<1$.
Note that (3.a) implies that
  • (3.a’)
    $\{(T^{(r)}_{e,i})^{2};r\in(0,1]\}$ for $i=1,2$ and $\{(T^{(r)}_{s})^{2};r\in(0,1]\}$ are uniformly integrable.
Furthermore, (3.b) and (3.c) imply that $1-\rho^{(r)}_{i}=rb_{i}/c_{i}+o(r)$, $\rho_{i}\equiv\frac{\lambda_{i}}{c_{i}\mu}=1$ and $\lambda_{i}=c_{i}\mu$. Hence, $\lambda_{1}=\lambda_{2}$ if and only if $c_{1}=c_{2}$. It is easy to see that $X^{(r)}(t)$ has the stationary distribution by (3.e). This distribution is denoted by $\nu^{(r)}$.
We take countably many $r$’s for a limiting operation. For such a sequence, there exists a stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator $\Theta_{\bullet}$ on which $X^{(r)}(\cdot)$ is a stationary process. In what follows, $\mathbb{P}$ is always this probability measure. By the stability condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution. Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$. For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$. For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$, and let $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper. We note the following simple but important facts.
Lemma 3.1.
Under the stationary assumption, $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
Proof.
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N_{e}(t)]\leq\mathbb{E}[N_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N_{e}(1)]\leq\mathbb{E}[N_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$. Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$. ∎
3.2 Main results
We are now ready to present a main result, which will be proved in Section 4. Recall that $L^{(r)}$ is a random variable subject to the stationary distribution $\nu^{(r)}$ of the $r$-th 2-level $GI/G/1$ queue. We decompose $\nu^{(r)}$ into marginal measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}>\ell^{(r)}_{0}),$
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}),$
Theorem 3.1.
Assume the assumptions (3.a)–(3.e), and let
$\sigma_{i}^{2}=\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2},\qquad\beta_{i}=\frac{2b_{i}}{c_{i}\sigma_{i}^{2}},\qquad i=1,2.$ (3.1)
For each $r\in(0,1]$, define $\widetilde{\nu}_{i|cd}$ as the probability measures on ${\mathcal{B}}(\mathbb{R}_{+})$ which have the density functions $h_{i|cd}$ for $i=1,2$ which are
$h_{1|cd}(x)=1(0\leq x\leq\ell_{0})\times\left\{\begin{array}[]{ll}\displaystyle 1/\ell_{0},&b_{1}=0,\\ \displaystyle\frac{e^{\beta_{1}\ell_{0}}\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}e^{-\beta_{1}x},&b_{1}\not=0,\end{array}\right.$ (3.4)
$h_{2|cd}(x)=1(x>\ell_{0})\beta_{2}e^{-\beta_{2}(x-\ell_{0})}.$ (3.5)
  • (i)
    For each $i=1,2$, if $\nu^{(r)}_{i}(\mathbb{R}_{+})$ is bounded away from 00 as $r\downarrow 0$, then $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$, namely,
    $\lim_{r\downarrow 0}\nu_{i|cd}^{(r)}(B)=\widetilde{\nu}_{icd}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.6)
  • (ii)
    Define probability measure $\widetilde{\nu}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ and constant $\widetilde{\alpha}_{e}$ as
    $\widetilde{\nu}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{2},\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}),$ (3.7)
    $\widetilde{\alpha}_{e}=\lambda_{1}\widetilde{A}_{1}+\lambda_{2}\widetilde{A}_{2},$ (3.8)
    where $\widetilde{A}_{i}$ for $i=1,2$ are given by
    $\widetilde{A}_{1}=\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)},\;&b_{1}\not=0,\end{cases}\qquad\widetilde{A}_{2}=1-\widetilde{A}_{1}.$ (3.9)
    If either $\lambda_{1}=\lambda_{2}$ (equivalently, $c_{1}=c_{2}$) or $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,1]$, then $\alpha^{(r)}_{e}$ converges to $\widetilde{\alpha}_{e}$ and $\nu^{(r)}$ weakly converges to $\widetilde{\nu}$, namely
    $\lim_{r\downarrow 0}\nu^{(r)}(B)=\widetilde{\nu}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.10)
  • (iii)
    If $\lambda_{1}>\lambda_{2}$ ($\lambda_{1}<\lambda_{2}$), then $\nu^{(r)}_{1|cd}$ ($\nu^{(r)}_{2|cd}$) weakly converges to $\widetilde{\nu}_{1|cd}$ ($\widetilde{\nu}_{2|cd}$, respectively), as $r\downarrow 0$.
Remark 3.1.
Note that, under one of the conditions of (ii) of this theorem, we have
$\lim_{r\downarrow 0}\nu^{(r)}_{i}(\mathbb{R}_{+})=\widetilde{A}_{i},\qquad i=1,2.$ (3.11)
Conversely, if (3.11) holds, then we have the weak convergence (3.10) from (i) because $\widetilde{A}_{i}>0$. Thus, (3.11) is equivalent to (3.10). We conjecture to have (3.10) without any condition in (ii). There are two supports for this conjecture. One is from a process limit for the 2-level $GI/G/1$ queue in heavy traffic. Another is from the intuitive verification of (3.11). They are detailed in Sections 5.1 and 5.2, repectively.
Let $S^{(r)}(n)$ be the sojourn time of the $n$-th arriving customer in service, then the traffic intensity $\rho^{(r)}$ can be defined under the stationary setting as
$\rho^{(r)}=\alpha^{(r)}_{e}\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}S^{(r)}_{k}.$ (3.12)
Hence, by Little’s law, we have $\rho^{(r)}=\mathbb{P}(L^{(r)}>0)=1-\mathbb{P}(L^{(r)}=0)$. In diffusion applications of a single server queue, the mean queue length is estimated through $(1-\rho^{(r)})\mathbb{E}[L^{(r)}]$ as $r\downarrow 0$. The following corollary computes this limit, whose details are given in Appendix A.1.
Corollary 3.1.
Under either one of the assumptions in (ii) of Theorem 3.1,
$\lim_{r\downarrow 0}(1-\rho^{(r)})\mathbb{E}[L^{(r)}]=\lim_{r\downarrow 0}\mathbb{P}(L^{(r)}=0)\mathbb{E}[L^{(r)}]$
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{2}\not=0.\end{cases}$ (3.13)
Let consider Theorem 3.1 when the inter-arrival times and service workload are all exponentially distributed. We call this case exponential.
Proposition 3.1.
Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{M}$ which is defined as $\widetilde{\nu}_{M}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{1}(M)+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{2}(M)$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
$\widetilde{A}_{1}(M)=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
and $\widetilde{A}_{2}(M)=1-\widetilde{A}_{1}(M)$.
Remark 3.2.
Since $\sigma_{i}^{2}=2$ and $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$ in the exponential case, $\widetilde{\nu}_{M}$ and $\widetilde{A}_{i}(M)$ agree with $\widetilde{\nu}$ and $\widetilde{A}_{i}$ in Theorem 3.1, respectively, for $i=1,2$.
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its independent proof in Appendix A.2. This proof is straightforward because $L^{(r)}(\cdot)$ is a birth-and-death process, and helpful to intuitively convince (ii) of Theorem 3.1.
4 Proofs
In this section we prove Theorem 3.1. Compared with the exponential case, a major difficulty comes from the fact that $Z^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions. This method is called a BAR approach, and has been used in [1, 2, 11]. Recently, this approach is further studied in [12].
The proof of Theorem 3.1 is composed of three parts. We first introduce Palm distributions and the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
4.1 Palm distribution and BAR
Recall that we are working on stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}$. Furthermore, by the stability assumption (3.e), we can assume that this framework is stationary, that is,
$\mathbb{P}(\Theta_{t}^{-1}A)=\mathbb{P}(A),\qquad A\in{\mathcal{F}}.$
[…]
$\mathbb{P}^{(r)}_{v}(A)=(\alpha^{(r)}_{v})^{-1}\mathbb{E}\left[\int_{(0,1]}1_{\Theta_{u}^{-1}A}N^{(r)}_{v}(du)\right],\qquad A\in{\mathcal{F}},v=e,s,$ (4.1)
where $1_{A}$ is the indicator function of the event $A\in{\mathcal{F}}$.
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [12]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S\equiv\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
${\mathcal{H}}f({\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f({\boldsymbol{x}})-(c_{1}1(1\leq x_{1}\leq\ell_{0})+c_{2}1(x_{1}>\ell_{0}))\frac{\partial^{+}}{\partial x_{3}}f({\boldsymbol{x}}),\qquad{\boldsymbol{x}}\in S,$
where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$, and define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
$L^{(r)}_{1}(t)=L(t-)+1(\Delta N_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
where $n_{e,1}(t)=\sum_{n=1}^{N_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$. We call $X^{(r)}_{1}(t)$ an intermediate state. Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N_{s}(t)))$ when $\Delta N_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N_{s}(t)=0$.
We further define difference operators $\Delta_{e}$ and $\Delta_{s}$ as
$\Delta_{e}f(X^{(r)}(t))=f(X^{(r)}_{1}(t))-f(X^{(r)}(t-)),$
$\Delta_{s}f(X^{(r)}(t))=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)),$
Then, from the elementary observation of the time evolution of $X^{(r)}(t)$, we have
$f(X^{(r)}(t)=f(X^{(r)}(0))+\int_{0}^{t}{\mathcal{H}}f(X^{(r)}(u))du+\sum_{u\in(0,t]}[\Delta_{e}f(X^{(r)}(u))+\Delta_{s}f(X^{(r)}(u))],$
where the summation is well defined because $\Delta_{e}f(X^{(r)}(u))$ and $\Delta_{s}f(X^{(r)}(u))$ vanish except for finitely many $u\in(0,t]$. Taking the expectation of the above equation by $\mathbb{P}$, we have a stationary equation:
$\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$
$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]=0,\qquad f\in C_{b}^{p},$ (4.2)
The formula (4.1) is a special case of the rate conservation law (e.g., see [10]), and referred as a basic adjoint relation ship, BAR in short.
Remark 4.1.
In using Palm distributions, we must be careful about the following two points.
  • (i)
    $X^{(r)}(0-)$ is not defined for $X^{(r)}(\cdot)=\{X^{(r)}(t);t\geq 0\}$, but well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. Thus, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined.
  • (ii)
    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N_{e}$ is counted at time 00 and just before when $N_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [12] for its details. We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to explicitly refer to the intermediate state $X^{(r)}_{1}(0)$. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
We note basic facts on $\alpha^{(r)}_{e}$ and $L^{(r)}(\cdot)$ under the Palm distributions.
Lemma 4.1.
Under conditions (3.a) and (3.c), $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$, and
$\mathbb{P}^{(r)}_{e}[(L^{(r)}(0-)=k]=\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=k],\qquad k\geq 0,$ (4.3)
$\alpha^{(r)}_{e}=\frac{\lambda^{(r)}_{1}\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)<\ell^{(r)}_{0})+\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)\geq\ell^{(r)}_{0})},$ (4.4)
$\alpha^{(r)}_{e}=c^{(r)}_{1}\mu^{(r)}\mathbb{P}(1\leq L^{(r)}\leq\ell^{(r)}_{0})+c^{(r)}_{2}\mu^{(r)}\mathbb{P}(L^{(r)}>\ell^{(r)}_{0}).$ (4.5)
Proof.
For each integer $k\geq 0$, let $f({\boldsymbol{x}})=x_{1}\wedge(k+1)$ for ${\boldsymbol{x}}\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N_{e}(u)-N_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N_{s}(u)-N_{s}(u-)=1)$, we have
$\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]$
$\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
Apply this to (4.1) yields
$\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$. Because of this equality, the above formula also yields (4.3). We next let $f({\boldsymbol{x}})=x_{2}$, and apply (4.1), then
$-1+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[T^{(r)}_{e,1}1(L^{(r)}(0-)<\ell^{(r)}_{0})+T^{(r)}_{e,2}1(L^{(r)}(0-)\geq\ell^{(r)}_{0})\right]=0$
because $(R^{(r)}_{e})^{\prime}(t)=-1$. This proves (4.4). Similarly, let $f({\boldsymbol{x}})=x_{3}$, then
$(R^{(r)}_{s})^{\prime}(t)=-c_{1}^{(r)}1(1\leq L^{(r)}(0)\leq\ell^{(r)}_{0})-c_{2}^{(r)}1(L^{(r)}(0)>\ell^{(r)}_{0}),$ (4.6)
and $\mathbb{E}^{(r)}_{s}[\Delta_{s}R^{(r)}_{s}(0)]=\mathbb{E}[T^{(r)}_{s}]$. Hence, (4.1) yields (4.5). ∎
Since it follows from (4.4) that
$\min(\lambda^{(r)}_{1},\lambda^{(r)}_{2})\leq\alpha^{(r)}_{e}\leq\max(\lambda^{(r)}_{1},\lambda^{(r)}_{2}),\qquad r\in(0,1],$ (4.7)
we have the following corollary.
Corollary 4.1.
$\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{s}$. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
4.2 Preliminary lemmas
We will prove Theorem 3.1 using the moment generating functions $\varphi_{i}(\theta_{2})$ and $\varphi_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$. Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
$\varphi_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
$\varphi_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
To this end, we introduce test functions for the BAR (4.1). We first define functions $w^{(r)}_{i}$ and $g^{(r)}_{i,\theta}$ for $i=1,2$ and $\theta\in\mathbb{R}$ as
$w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{0}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{0}),\qquad z\in\mathbb{Z}_{+},$
$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{1}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r)},\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
where $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are the unique solutions of the following equations for each $\theta,\theta\in\mathbb{R}$.
$e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1.$ (4.8)
We note the following lemma, which is obtained from Lemma 5.8 of [2].
Lemma 4.2.
Under the assumptions (3.a)–(3.d), we have the following expansions as $r\downarrow 0$.
$\eta_{i}^{(r)}(r\theta)=\lambda^{(r)}_{i}r\theta+\frac{1}{2}(\lambda^{(r)}_{i})^{3}(\sigma^{(r)}_{e,i})^{2}r^{2}\theta^{2}+o(r^{2}),\qquad i=1,2,$ (4.9)
$\zeta^{(r)}(r\theta)=-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}r^{2}\theta^{2}+o(r^{2}).$ (4.10)
Furthermore, by , for $i=1,2$, there are constant $d_{e,i},d_{s},a>0$ such that, for $r\in(0,1]$ and $\theta\in\mathbb{R}$ satisfying $r|\theta|<a$,
$\left|\eta^{(r)}_{i}(r\theta)(u_{1}\wedge 1/r)+\zeta^{(r)}(r\theta)(u_{2}\wedge 1/r)\right|\leq|\theta|\left(d_{e,i}(ru_{1}\wedge 1)+d_{s}(ru_{2}\wedge 1)\right).$ (4.11)
By this lemma, $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ is bounded by $e^{|\theta_{i}|(d_{e,i}+d_{s})}$ for sufficiently small $r$ for each $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, and therefore, we have the following facts.
Corollary 4.2.
Let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$, then $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ converges to $1$ as $r\downarrow 0$, and
$\lim_{r\downarrow 0}\mathbb{E}\left[g^{(r)}_{i,r\theta_{i}}(R^{(r)})\right]=1,\qquad\lim_{r\downarrow 0}\mathbb{E}_{u}\left[g^{(r)}_{i,r\theta_{i}}(R^{(r)})\right]=1,\quad u=e,s.$ (4.12)
We define test function $f^{(r)}_{{\boldsymbol{\theta}}}$ for the BAR as
$f^{(r)}_{{\boldsymbol{\theta}}}(z,x_{1},x_{2})=\sum_{i=1,2}e^{\theta_{i}z}g^{(r)}_{i,\theta_{i}}(x_{1},x_{2})w^{(r)}_{i}(z),$ (4.13)
and auxiliary functions $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ as
$\psi^{(r)}_{i}(\theta_{i})=\mathbb{E}\left[e^{r\theta_{i}L^{(r)}}g^{(r)}_{i,r\theta_{i}}(R^{(r)})w^{(r)}_{i}(L^{(r)})\right],\qquad\theta_{1}\in\mathbb{R},\theta_{2}\leq 0.$ (4.14)
Then, by Lemma 4.2 and Corollary 4.2, we have
$\psi^{(r)}_{i}(\theta_{i})<\infty,\quad\mathbb{E}_{u}\left[f^{(r)}_{r{\boldsymbol{\theta}}}(L^{(r)},R^{(r)})\right]<\infty,\quad u=e,s,\quad\mbox{uniformly in $r\in(0,1]$},$ (4.15)
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations bellow for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{{\boldsymbol{\theta}}}$. Then,
$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\sum_{i=1,2}\eta_{i}^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})$
$\qquad+\sum_{i=1,2}c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})-c^{(r)}_{1}\zeta^{(r)}(r\theta_{1})\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)}))\right].$ (4.16)
Since $\lambda^{(r)}_{i}-c^{(r)}_{i}\mu^{(r)}=-b_{i}\mu^{(r)}r+o(r)$, it follows from Lemma 4.2, $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ and $\lambda_{i}=c_{i}\mu$ that
$\eta^{(r)}_{i}(r\theta_{i})+c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})$ $=-b_{i}r^{2}\mu\theta_{i}+\frac{1}{2}c_{i}(\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2})r^{2}\mu\theta_{i}^{2}+o(r^{2})$
$=-\frac{1}{2}c_{i}\sigma_{i}^{2}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o(r^{2}).$
Hence, from (4.2), we have
$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\left(-\beta_{i}+\theta_{i}\right)\mu\theta_{i}r^{2}\psi^{(r)}_{i}(\theta_{i})$
$\hskip 64.58313pt+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}),$ (4.17)
We next compute the Palm expectation terms in (4.1). In what follows, we use the following simplified notations.
$\widehat{T}^{(r)}_{e,i}\equiv T^{(r)}_{e,i}(0)\wedge 1/r,\quad i=1,2,\qquad\widehat{T}^{(r)}_{s}\equiv T^{(r)}_{s}(0-)\wedge 1/r,$
$\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$
Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L(^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L(^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
Hence, from the definitions of $X^{(r)}_{1}(0)$ and $\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))$, we have
$\mathbb{E}^{(r)}_{e}[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{e}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad=\mathbb{E}^{(r)}_{e}\Big{[}e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0))1(L^{(r)}_{1}(0)=\ell^{(r)}_{0}+1)$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})\Big{]}$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.18)
where the third equality is obtained by applying (4.8). Similarly, we have
$\mathbb{E}^{(r)}_{s}[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{s}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0)))1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0)=\ell^{(r)}_{0}+1)\Big{]}$
$\quad=e^{r\theta_{1}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad=(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\qquad+e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.19)
Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]\right)$. Then, by (4.2) and (4.2),
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.20)
Thus, applying (4.2), (4.2) and (4.2) to (4.1), we have a pre-limit BAR for the 2-level $GI/G/1$ queue.
Lemma 4.3.
For each fixed $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, we have, as $r\downarrow 0$,
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o(r^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
$\quad-c_{1}\left(\mu r\theta_{1}+\mu^{3}\sigma_{s}^{2}r^{2}\theta_{1}^{2}/2+o(r^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})]=D^{(r)}(\theta_{1},\theta_{2}).$ (4.21)
This BAR is called an asymptotic BAR, and a starting point of our proof of Theorem 3.1. We will divide this BAR by $r^{2}$, and take its limit as $r\downarrow 0$, expanding the exponential functions in (4.3). To perform this expansion, we consider asymptotic behaviors on $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ as $r\downarrow 0$.
Lemma 4.4.
Under the assumption (3.a), for $k=0,1,2$ as $r\downarrow 0$,
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}\in B)],$ (4.22)
$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
and, for $k=1,2$,
$\mathbb{E}[(\widehat{T}^{(r)}_{u})^{k}]=\mathbb{E}[(T^{(r)}_{u})^{k}]+o(r^{3-k}),\qquad u=(e,1),(e,2),s,$ (4.24)
$\mathbb{E}[(\widehat{R}^{(r)}_{v})^{k}]=\mathbb{E}[(R^{(r)}_{v})^{k}]+o(r^{2-k}),\qquad v=e,s,$ (4.25)
$\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]=\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}]+o(r^{3-k}),$ (4.26)
$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}]=\mathbb{E}^{(r)}_{e}[(R^{(r)}_{s}(0-))^{k}]+o(r^{3-k}).$ (4.27)
In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable.
This lemma is easily proved using the BAR (4.1), but its proof is lengthy, so we defer it to Appendix A.3. Because of the finiteness of the third moments in the assumption (3.a), the following corollary is immediate from Lemma 4.4.
Corollary 4.3.
For each $k=1,2$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
Note that the weak limit of $\nu^{(r)}$ should be obtained from (4.3). However, we will see that it requires the relation between the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, which is not easy to get from (4.3). To relax this difficulty, we prepare two lemmas below.
Lemma 4.5.
Under the assumption (3.a), we have, for $k=1,2,3$ and as $r\downarrow 0$,
$c^{(r)}_{1}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(1\leq L^{(r)}\leq\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.28)
$c^{(r)}_{2}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.29)
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}\leq\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,1})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]\Big{)},$ (4.30)
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)},$ (4.31)
Proof.
For $k\geq 1$, we first prove (4.5). We apply $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,L^{(r)}(t)\leq\ell^{(r)}_{0}).$
Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ and $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, respectively. ∎
Define, for $i=1,2$,
$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}].$ (4.32)
Then, $\Delta^{(r)}_{i}$ is well defined and finite by Corollary 4.3, and we have
Lemma 4.6.
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-r\mu b_{1}\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.33)
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]+o(r).$ (4.34)
Proof.
For $k=1$, multiply both sides of (4.5) and (4.5) by $\mu^{(r)}$ and $\lambda^{(r)}_{1}$, respectively, and take their difference. Since $c^{(r)}_{1}\mu^{(r)}-\lambda^{(r)}_{1}=r\mu^{(r)}b_{1}+o(r)$ by the assumption (3.c), $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.22) and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.23), we have
$(r\mu^{(r)}b_{1}+o(r))\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]-c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1. This proves (4.33). Similarly, (4.34) is obtained from (4.5) and (4.5), ∎
We will use Lemma 4.6 to determine the limits of $\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]$ and $\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]$, equivalently, $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, as $r\downarrow 0$ under some conditions.
4.3 Proof of Theorem 3.1
We first prove (i). We are now ready to take the limit of the BAR of (4.3) as $r\downarrow 0$. For this, we will use the following elementary expansions.
Lemma 4.7.
For each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$,
$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}+o(r),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r)+o(r),\quad i=1,2,$ (4.35)
$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+o(r),$ (4.36)
$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r).$ (4.37)
Lemma 4.8.
For each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$,
$e^{-\eta^{(r)}_{i}(r\theta_{i})\widehat{R}^{(r)}_{e}}=1-r\lambda_{i}\theta_{i}\widehat{R}^{(r)}_{e}-2^{-1}r^{2}\lambda_{i}^{3}\sigma_{e,i}^{2}r\theta_{i}^{2}(\widehat{R}^{(r)}_{e})^{2}+o(r^{2}),$ (4.38)
$e^{-\zeta^{(r)}(r\theta_{i})\widehat{R}^{(r)}_{s}}=1+\mu r\theta_{i}\widehat{R}^{(r)}_{s}-2^{-1}r^{i}\mu^{3}\sigma_{s}^{2}\theta_{i}^{2}(\widehat{R}^{(r)}_{s})^{2}+o(r^{2}),\qquad i=1,2,$ (4.39)
$e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}$
$\hskip 34.44434pt=r(\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}+2^{-1}r^{2}(\lambda_{2}^{3}\sigma_{e,2}^{2}r\theta_{2}^{2}-\lambda_{1}^{3}\sigma_{e,1}^{2}r\theta_{1}^{2})(\widehat{R}^{(r)}_{e})^{2}+o(r^{2}).$ (4.40)
$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}}$
$\hskip 34.44434pt=\mu r(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}-2^{-1}r^{2}\mu^{3}\sigma_{s}^{2}(\theta_{2}^{2}-\theta_{1}^{2})(\widehat{R}^{(r)}_{s})^{2})+o(r^{2}),$ (4.41)
These lemmas are proved in Appendix A.4. From BAR (4.3) in Lemma 4.3, we aim to determine the limiting measures $\widetilde{\nu}_{1}$ and $\widetilde{\nu}_{2}$. This will be done step by step.
(1st Step)   We expand (4.3) using Lemmas 4.7 and 4.8. For this, we need the asymptotic behaviors concerning $\widehat{R}^{(r)}_{e}$ and $\widehat{R}^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{s}$ and $\mathbb{P}^{(r)}_{e}$, respectively. To see them, we first put $\theta_{2}=\theta_{1}=\theta$ in (4.3), then
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o(r^{2})\big{)}\psi^{(r)}_{i}(\theta)$
$\quad=c_{1}\left(\mu r\theta+\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o(r^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta),$ (4.42)
where applying (4.35), (4.8) and Corollary 4.3 to (4.2) yields
$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+\theta r+o(r))\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r(\lambda_{2}-\lambda_{1})\theta(1+\theta r)\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r^{2})$
$\quad=\alpha^{(r)}_{e}e^{\theta\ell_{0}}r(\lambda_{2}-\lambda_{1})\theta\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r^{2}).$
Hence, dividing both sides of (4.3) by $r$ and letting $r\downarrow 0$, we have
$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
Since $e^{r\theta\ell_{0}}$ can be arbitrarily small, this implies that
$\lim_{r\downarrow 0}\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=0.$
and therefore, by Corollary 4.2, we have
$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}=0]=0.$ (4.43)
We next put $\theta_{1}=0$ and $\theta_{2}=\theta$ in (4.3). In this case,
$\frac{1}{2}c_{2}\sigma_{2}^{2}\big{(}(\beta_{2}-\theta)\mu\theta r^{2}+o(r^{2})\big{)}\psi^{(r)}_{2}(\theta)+D^{(r)}(0,\theta)=0,$
and, from (4.2),
$D^{(r)}(0,\theta)=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}-1)\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
$\quad+\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(1-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])$
$\quad+\alpha^{(r)}_{e}(\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]-e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})])$
$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})])$
$=\alpha^{(r)}_{e}r\theta e^{r\theta\ell^{(r)}_{0}}\big{(}\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda_{2}(1+r\theta)\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{)}+O(r^{2}),$
where the 3rd equation is obtained by (4.3), and the last one is by Corollary 4.3 and Lemma 4.8. Since $D^{(r)}(0,\theta)=O(r^{2})$ from (4.3), we have
$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\qquad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
because $\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are nonnegative. These results lead to the following asymptotic results, which strengthens Corollary 4.3.
Lemma 4.9.
$\lim_{r\downarrow 0}\mathbb{P}^{(r)}_{e}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,\quad\lim_{r\downarrow 0}\mathbb{P}^{(r)}_{s}[L^{(r)}(0-)=\ell^{(r)}_{0}]=0,$ (4.45)
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,\quad\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=0.$ (4.46)
Proof.
(4.45) is immediate from Lemma 4.6 by applying (4.43) and (4.44). As for (4.46), we only prove its first equation because its second one is similarly proved. By the Schwartz’s inequality,
$\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}=\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3/2}(\widehat{R}^{(r)}_{s})^{1/2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}$
$\leq\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})].$ (4.47)
From (4.5) for $k=3$ and Lemma 4.4, we have
$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad\leq\mathbb{E}[(T^{(r)}_{s})^{3}]+3c_{1}\mathbb{E}[(R^{(r)}_{s})^{2}1(R^{(r)}_{s}\geq 1/r,L^{(r)}\geq 1)]/\alpha^{(r)}$
$\quad=\mathbb{E}[(T^{(r)}_{s})^{3}]+\frac{c_{1}}{\min_{i=1,2}c_{i}}\mathbb{E}[(T^{(r)}_{s})^{3}1(T^{(r)}_{s}\geq 1/r)],$
which is uniformly bounded for $r\in(0.1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
(2nd Step)   In this step, we first expand the exponential terms in $D^{(r)}(\theta_{1},\theta_{2})$ of (4.2), using Lemmas 4.7 and 4.8 in the step 1, then
$r^{-1}D^{(r)}(\theta_{1},\theta_{2})$
$=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}-e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(r^{-1}+\mu\theta_{2}\widehat{R}^{(r)}_{s}-2^{-1}r\mu^{3}\sigma_{s}^{2}\theta_{2}^{2}(\widehat{R}^{(r)}_{s})^{2})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(\mu(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}-2^{-1}r\mu^{3}\sigma_{s}^{2}(\theta_{2}^{2}-\theta_{1}^{2})(\widehat{R}^{(r)}_{s})^{2})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r))$
×$\hskip 51.6665pt\times\mathbb{E}^{(r)}_{s}[(r^{-1}-\lambda_{1}\theta_{1}\widehat{R}^{(r)}_{e}-2^{-1}r\lambda_{1}^{3}\sigma_{e,1}^{2}\theta_{1}^{2}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}(1+\theta_{2}r)+o(r))\mathbb{E}^{(r)}_{s}[((\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}$
$\hskip 51.6665pt+2^{-1}r(\lambda_{2}^{3}\sigma_{e,2}^{2}r\theta_{2}^{2}-\lambda_{1}^{3}\sigma_{e,1}^{2}r\theta_{1}^{2})(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})]+o(r).$ (4.48)
We next apply Lemma 4.9 to this formula, then
$r^{-1}D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}-e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(r^{-1}+\mu\theta_{2}\widehat{R}^{(r)}_{s})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[\mu(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r))\mathbb{E}^{(r)}_{s}[(r^{-1}-\lambda_{1}\theta_{1}\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}(1+\theta_{2}r)+o(r))\mathbb{E}^{(r)}_{s}[(\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r)$
$=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}\theta_{2}-e^{\theta_{1}\ell_{0}}\theta_{1})(\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}])$
$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}\lambda_{2}\theta_{2}-e^{\theta_{1}\ell_{0}}\lambda_{1}\theta_{1})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+o(r).$ (4.49)
Recall that
$\Delta^{(r)}_{i}=\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\hskip 43.05542pt-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}],\qquad i=1,2,$
then (4.3) can be written as
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}r\big{(}e^{\theta_{2}\ell_{0}}\theta_{2}\Delta^{(r)}_{2}-e^{\theta_{1}\ell_{0}}\theta_{1}\Delta^{(r)}_{1}\big{)}+o(r^{2}).$ (4.50)
Substituting this into BAR (4.3) and dividing by $r$, we have
$\frac{1}{2}\sum_{i=1,2}r\mu c_{i}\sigma_{i}^{2}\big{(}\beta_{i}-\theta_{i}\big{)}\theta_{i}\psi^{(r)}_{i}(\theta_{i})$
$\quad=c_{1}\mu\theta_{1}\mathbb{P}[L^{(r)}=0]+\alpha^{(r)}_{e}\big{(}e^{\theta_{2}\ell_{0}}\theta_{2}\Delta^{(r)}_{2}-e^{\theta_{1}\ell_{0}}\theta_{1}\Delta^{(r)}_{1}\big{)}+o(r).$ (4.51)
Substituting $\theta_{2}=0$ or $\theta_{1}=0$, this asymptotic BAR is split into two equations:
$\frac{1}{2}r\mu c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-\alpha^{(r)}_{e}e^{\theta_{1}\ell_{0}}\Delta^{(r)}_{1}+o(r),$ (4.52)
$\frac{1}{2}r\mu c_{2}\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\Delta^{(r)}_{2}+o(r).$ (4.53)
(3rd Step)   We compute $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ from (4.52) and (4.53) separately. First consider (4.53). Note that (4.34) yields
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (4.54)
Hence, from (4.53), we have
$\frac{1}{2}r\mu c_{2}\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})$ $=\frac{1}{2}e^{\theta_{2}\ell_{0}}r\mu c_{2}\sigma_{2}^{2}\beta_{2}\varphi^{(r)}_{2}(0)+o(r),$
which is simply written as
$\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})$ $=e^{\theta_{2}\ell_{0}}\beta_{2}\varphi^{(r)}_{2}(0)+o(1).$ (4.55)
We next consider (4.52). From (4.33) of Lemma 4.6,
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-r\mu b_{1}\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r).$ (4.56)
Hence, from (4.52), we have
$r\mu c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})$
$\qquad=2c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0)]+2r\mu b_{1}e^{\theta_{1}\ell_{0}}\varphi^{(r)}_{1}(0)+o(r).$ (4.57)
This yields
$\mathbb{P}[L^{(r)}=0)]=r\frac{c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})-2b_{1}e^{\theta_{1}\ell_{0}}\varphi^{(r)}_{1}(0)}{2c_{1}(1-e^{\theta_{1}\ell_{0}})}+o(r).$ (4.58)
We separately consider (4.58) for $b_{1}=0$ and $b_{1}\not=0$.
(Case of $b_{1}=0$)   In this case, $\beta_{1}=b_{1}=0$, and therefore (4.58) yields
$\mathbb{P}[L^{(r)}=0)=r\lim_{\theta_{1}\to 0}\frac{r\sigma_{1}^{2}\theta_{1}\psi^{(r)}_{1}(\theta_{1})}{2(e^{\theta_{1}\ell_{0}}-1)}+o(r)=\frac{r\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)+o(r).$ (4.59)
Hence, for $b_{1}=0$, it follows from (4.56) and (4.59) that
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)+o(r),$ (4.60)
$\psi^{(r)}_{1}(\theta_{1})=\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\varphi^{(r)}_{1}(0)+o(r).$ (4.61)
(Case of $b_{1}\not=0$)   By substituting $\theta_{1}=\beta_{1}$ into (4.58), we have
$\mathbb{P}[L^{(r)}=0)]=\frac{rb_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\varphi^{(r)}_{1}(0)+o(r)$ (4.62)
and therefore, it follows from (4.56), (4.3) and (4.62) that
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\frac{r\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)+o(r),$ (4.63)
$\psi^{(r)}_{1}(\theta_{1})=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\varphi^{(r)}_{1}(0)+o(r).$ (4.64)
(4th Step)   We now complete the proof of Theorem 3.1. We first prove (i). Note that the bounded away assumption can be written as
$\liminf_{r\downarrow 0}\varphi^{(r)}_{i}(0)>0,\qquad i=1,2.$ (4.65)
Under this condition, for $\theta_{1}\in\mathbb{R}_{+}$ and $\theta_{2}\leq 0$,
$\lim_{r\downarrow 0}|\psi^{(r)}_{i}(\theta_{i})/\varphi^{(r)}_{i}(0)-\varphi^{(r)}_{i|cd}(\theta_{i})|=0.$ (4.66)
is equivalent to
$\lim_{r\downarrow 0}|(\psi^{(r)}_{i}(\theta_{1})-\varphi^{(r)}_{i}(\theta_{1}))|=0,$ (4.67)
since $\varphi^{(r)}_{i|cd}(\theta_{i})=\varphi^{(r)}_{i}(\theta_{i})/\varphi^{(r)}_{i}(0)$. Here, (4.67) holds because (4.12) implies that
$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{0})]\right|$
$\leq e^{|\theta_{1}|r\ell^{(r)}_{0}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
Hence, (4.66) holds. Applying it to (4.61) and (4.64) for $i=1$, we obtain
$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b\not=0,\end{cases}$ (4.68)
and, similarly, for $i=2$, (4.55) yields
$\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{0}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (4.69)
Since $\widetilde{\varphi}_{i|cd}$ is the moment generating function of $\widetilde{\nu}_{i|cd}$, (4.68) and (4.69) are equivalent to that $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$ for $i=1,2$, respectively. Thus, (i) is proved.
We next prove (ii). For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumption in (ii). First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$. In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{e}$ under $\mathbb{P}^{(r)}_{s}$, so, by Lemma 4.4 for $k=1$,
$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)$ $=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
Applying this fact to the definitions of $\Delta^{(r)}_{i}$ yields
$\lambda^{(r)}_{2}\Delta^{(r)}_{1}-\lambda^{(r)}_{1}\Delta^{(r)}_{2}$
$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (4.70)
Thus, if $b_{1}=0$, then it follows from (4.54) and (4.60) that
$\lambda_{2}\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Since $\mu c_{i}=\lambda_{i}$, this can be written as
$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$
Hence,
$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}\sigma_{1}^{2}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (4.71)
On the other hand, if $b_{1}\not=0$, then, from (4.54) and (4.63),
$\frac{\lambda_{2}\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Hence,
$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (4.72)
Thus, (4.65) is satisfied for $i=1,2$.
We next consider the case that $\lambda_{1}=\lambda_{2}$, equivalently, $c_{1}=c_{2}$. Then, from the definitions of $\Delta^{(r)}_{i}$, we have
$\Delta^{(r)}_{1}-\Delta^{(r)}_{2}=(\lambda^{(r)}_{1}-\lambda^{(r)}_{2})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=o(r),$ (4.73)
because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). On the other hand, if $b_{1}=0$, then (4.54) and (4.60) yield
$\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-\Delta^{(r)}_{2})=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r),$
Hence,
$\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\mu b_{2}(1-\varphi^{(r)}_{1}(0))+o(1)$
Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54) and (4.63) yield
$\frac{\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Again by $c_{1}=c_{2}$, this proves (4.72).
Let
$\widetilde{\varphi}_{i}(\theta)=\widetilde{\varphi}_{i|cd}(\theta)\widetilde{\varphi}_{i}(0),\qquad i=1,2,$
then $\widetilde{\varphi}_{i}$ is the moment generating function of $\widetilde{\nu}_{i}$, so (4.68) and (4.69) together with (4.71) and (4.71) prove that $\nu^{(r)}_{i}$ weakly converges to $\widetilde{\nu}_{i}$ for $i=1,2$. Thus, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}$ of (3.7). Similarly, by the assumption (3.c), (4.5), (4.59) and (4.62), we have
$\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\lim_{r\downarrow 0}(c^{(r)}_{1}\mu^{(r)}(\varphi^{(r)}_{1}(0)-\mathbb{P}[L^{(r)}=0])+c^{(r)}_{2}\mu^{(r)}\varphi^{(r)}_{2}(0))=\lambda_{1}\widetilde{\varphi}_{1}(0)+\lambda_{2}\widetilde{\varphi}_{2}(0).$
It remains to prove (iii). Assume that $\lambda_{1}>\lambda_{2}$, then $\Delta^{(r)}_{1}>\Delta^{(r)}_{2}$ by (4.73). For $b=0$, this yields
$\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
and therefore $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$. Similarly, for $b\not=0$, we also have $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$. Hence, (4.65) holds for $i=1$, so $\nu^{(r)}_{1|cd}$ weakly converges to $\widetilde{\nu}_{1|cd}$ as $r\downarrow 0$ by (i). Similarly, if $\lambda_{1}<\lambda_{2}$, then (4.65) holds for $i=2$, so $\nu^{(r)}_{2|cd}$ weakly converges to $\widetilde{\nu}_{2|cd}$.
5 Concluding remarks
In this section, we discuss two subjects, a process limit for the 2-level $GI/G/1$ queue in heavy traffic and a verification of (3.11) in Remark 3.1.
5.1 Process limit
Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic. From the assumptions (3.a)–(3.e), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation:
$\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})$
$\;\;+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)+\beta_{2}e^{\theta_{2}\ell_{0}}\theta\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\quad\theta\in\mathbb{R}\times\mathbb{R}_{-},$ (5.1)
which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
We like to identify this stationary distribution $\widetilde{\nu}$ by that of the reflecting diffusion process. Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$, and define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
$\sigma(x)=\sqrt{c_{1}}\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\sigma_{2}1(x>\ell_{0}),$
$b(x)=b_{1}1(x\leq\ell_{0})+b_{2}1(x>\ell_{0}),$
then one may expect that $Z(\cdot)$ satisfies the following stochastic integral equation.
[…]
where $W(\cdot)$ is the standard Brownian motion, $Y(\cdot)\equiv\{Y(t);t\geq 0\}$ is a non-deceasing process satisfying that $\int_{0}^{t}1(Z(u)>0)Y(du)=0$ for $t\geq 0$. If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [7] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$.
However, $Y(\cdot)$ may not be absolutely continuous. So far, assume that there is a weak solution of (5.2). For notational convenience, we continue to use the same notations for this weak solution. To apply Ito integration rule to (5.2), let us compute the quadratic variation $\big{[}Z\big{]}_{t}$ of $Z(\cdot)$. Define $M(\cdot)\equiv\{M(t);t\geq 0\}$ as
[…]
then $M(\cdot)$ is a martingale. Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$. Then, it follows from (5.2) that
[…]
Hence, by Ito integration rule, we have, for $f\in C^{2}_{b}(\mathbb{R}_{+})$, where $C^{2}_{b}(\mathbb{R}_{+})$ is the set of all bounded and twice differentiable function from $\mathbb{R}_{+}$ to $\mathbb{R}$ which have the derivatives from the right at the origin, we have
[…]
where $\Gamma$ is operator on the space $C^{2}_{b}(\mathbb{R}_{+})$ defined as
$\Gamma f(x)=-b(x)f^{\prime}(x)+\frac{1}{2}\sigma^{2}(x)f^{\prime\prime}(x),\qquad x\geq 0.$
In particular, let $f$ be
$f(x)=e^{\theta_{1}x}1(0\leq x\leq\ell_{0})+e^{\theta_{2}x}1(x>\ell_{0}),\qquad\theta_{i}\leq 0,i=1,2.$ (5.6)
If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.5) for this $f$ yields
$\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})+\theta_{1}\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$
which implies that $\widetilde{\varphi}_{1}(\theta_{1})$ is the moment generating function of the exponential distribution and $\widetilde{\varphi}_{2}(\theta_{2})=0$ for all $\theta_{2}\leq 0$. Obviously, this is absurd. Thus, there is no solution $Z(\cdot)$ for (5.2) if $Z(\cdot)$ is a stationary process.
What is a right stochastic integral equation for $Z(\cdot)$ Although $Z(\cdot)$ is not reflected at $\ell_{0}$, we may consider the discontinuous change of $\sigma(Z(u))$ at $Z(u)=\ell_{0}$ as something like reflection. This motivates us to define local times $L_{\ell_{0}}^{-}(t)$ and $L_{\ell_{0}}^{+}(t)$ at $\ell_{0}$ (e.g. see Theorem 22.5 of [6] for the definition of a local time) by
[…]
and replace (5.2) by
[…]
Then, applying Ito integration rule to this stochastic integral equation for $f$ of (5.6) and taking for $Z(\cdot)$ to be stationary yield (5.1). This suggests the following conjectures.
  • (a)
    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to $Z(\cdot)$ satisfying (5.9) as $r\downarrow 0$.
  • (b)
    Stochastic integral equation (5.9) has a unique weak solution.
Both are hard to verify, so we leave them for future work. Since $Z(\cdot)$ of (5.9) is specified by $b_{i},c_{i},\sigma_{i}^{2}$ for $i=1,2$ and $\ell_{0}$, its stationary distribution must be $\widetilde{\nu}$ of (3.7). This supports the conjecture in Remark 3.1 because of (ii) of Theorem 3.1.
5.2 Intuitive verification of (3.11)
We intuitively show (3.11). First assume that $b_{1}\not=0$. Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{0}$ from below or above. Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ after time $B^{(r)}_{2}$. Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, we have
$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1).$ (5.10)
On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below. Then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that
$\mathbb{E}[B^{(r)}_{1}]\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (5.11)
Since $B^{(r)}_{2}+B^{(r)}_{1}$ is the cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{\mathbb{E}[B^{(r)}_{1}]}{\mathbb{E}[B^{(r)}_{1}]+\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
Thus, for $b_{1}\not=0$, (3.11) is intuitively obtained for $i=1$, and it is similarly obtained for $i=2$. The case for $b_{1}=0$ is similarly derived. Thus, we have another support of (3.11).
Appendix
A.1 Proofs of Corollary 3.1
We only verify (3.1) for $b_{1}\not=0$ because it is similarly proved for $b_{1}=0$. First, from (4.62) and (3.9),
$\frac{1}{r}\mathbb{P}[L^{(r)}=0)]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
$=\frac{b_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1).$
On the other hand, from (3.7),
$\lim_{r\downarrow 0}\mathbb{E}[rL^{(r)}]$
$\quad=\frac{c_{1}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1-\beta_{1}\ell_{0})}{2b_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+\left(\ell_{0}+\frac{c_{2}\sigma_{2}^{2}}{2b_{2}}\right)\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}$
$\quad=\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))}.$
Hence, we have (3.1).
A.2 Proof of Proposition 3.1
Because of the exponentially distributed assumption, for each fixed $r$, $L^{(r)}(\cdot)$ is a continuous-time Markov chain with transition rates $q^{(r)}_{i,j}$ for $i,j\in\mathbb{Z}_{+}$ given by
$q^{(r)}_{0,0}=-c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{0,1}=\lambda^{(r)}_{1},$
$q^{(r)}_{i,i}=-(c^{(r)}_{1}\mu^{(r)}+\lambda^{(r)}_{1}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{1},\quad 1\leq i\leq\ell^{(r)}_{0}-1,$
$q^{(r)}_{i,i}=-(c^{(r)}_{1}\mu^{(r)}+\lambda^{(r)}_{2}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{2},\quad i=\ell^{(r)}_{0},$
$q^{(r)}_{i,i}=-(c^{(r)}_{2}\mu^{(r)}+\lambda^{(r)}_{2}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{2}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{2},\quad i\geq\ell^{(r)}_{0}+1,$
where all the other $q_{i,j}$’s vanish. Thus, $L^{(r)}(\cdot)$ is a birth-and-death process. By the stability condition (3.e), it has the stationary distribution. Denote it by $\pi^{(r)}_{n}=\mathbb{P}(L^{(r)}=n)$ for $n\geq 0$, then it is easy to see that
$\pi^{(r)}_{n}=\left\{\begin{array}[]{ll}\pi^{(r)}_{0}(\rho^{(r)}_{1})^{n},&0\leq n\leq\ell^{(r)}_{0},\\ \pi^{(r)}_{0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}(\rho^{(r)}_{2})^{n-\ell^{(r)}_{0}},&n>\ell^{(r)}_{0},\end{array}\right.$ (A.3)
where
$\pi^{(r)}_{0}=\left\{\begin{array}[]{ll}\displaystyle\frac{(1-\rho^{(r)}_{1})(1-\rho^{(r)}_{2})}{1-\rho^{(r)}_{2}+(\rho^{(r)}_{2}-\rho^{(r)}_{1})(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}},&\rho^{(r)}_{1}\not=1,\\ \displaystyle\frac{1-\rho^{(r)}_{2}}{(\ell^{(r)}_{0}+1)(1-\rho^{(r)}_{2})+\rho^{(r)}_{2}},&\rho^{(r)}_{1}=1.\end{array}\right.$ (A.6)
Hence, we can directly compute $\widetilde{\nu}_{i}(\mathbb{R}_{+})$ for $i=1,2$ through $\pi^{(r)}_{0}$ of (A.6), which yields
$\nu^{(r)}_{1}(\mathbb{R}_{+})=\sum_{n=0}^{\ell^{(r)}_{0}}\pi^{(r)}_{n}=\left\{\begin{array}[]{ll}\displaystyle\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}+1}}{1-\rho^{(r)}_{1}}\pi^{(r)}_{0},&\rho^{(r)}_{1}\not=1,\\ \displaystyle(\ell^{(r)}_{0}+1)\pi^{(r)}_{0},&\rho^{(r)}_{1}=1,\end{array}\right.$ (A.9)
$\nu^{(r)}_{2}(\mathbb{R}_{+})=\sum_{n=\ell^{(r)}_{0}+1}^{\infty}\pi^{(r)}_{n}=\frac{\rho^{(r)}_{2}}{1-\rho^{(r)}_{2}}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\pi^{(r)}_{0}.$ (A.10)
We now consider the asymptotic behaviors of $\pi^{(r)}_{0}$ and $\widetilde{\nu}_{i}(\mathbb{R}_{+})$ for $i=1,2$. Form the definition of $\rho^{(r)}_{i}$ and the assumptions (3.b) and (3.c),
$\rho^{(r)}_{i}=\frac{\lambda^{(r)}_{i}}{c^{(r)}_{i}\mu^{(r)}}=1-\frac{r\mu b_{i}+o(r)}{c^{(r)}_{i}\mu^{(r)}}=1-\frac{b_{i}}{c_{i}}r+o(r),\qquad r\downarrow 0.$
We first consider the case that $b_{1}\not=0$. Since $\beta_{i}=b_{i}/c_{i}$ by the exponential assumption,
$(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\big{(}1-\beta_{1}r+o(r)\big{)}^{r^{-1}\ell_{0}+o(1)}=e^{-\beta_{1}\ell_{0}}+o(1)),$ (A.11)
and therefore (A.6) yields
$\pi^{(r)}_{0}$ $=\frac{\beta_{1}\beta_{2}r^{2}+o(r^{2})}{\beta_{2}r+o(r)+(e^{-\beta_{1}\ell_{0}}+o(1))((\beta_{1}-\beta_{2})r+o(r))}$
$=\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))=\frac{b_{1}}{c_{1}}\frac{c_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}r+o(r),$ (A.12)
which is compatible with (4.59). Hence, (A.9) yield
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})$ $=\lim_{r\downarrow 0}\frac{1-e^{-\beta_{1}\ell_{0}}+o(1)}{\beta_{1}r+o(r)}\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))$
$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}.$
Thus, this and $\widetilde{A}_{1}+\widetilde{A}_{2}=1$ concludes (4.35).
We next consider the case that $b_{1}=0$. In this case, $\rho^{(r)}_{1}=1-o(r)$, and therefore, from (A.6),
$\pi^{(r)}_{0}$ $=\frac{\beta_{2}}{1+\beta_{2}\ell_{0}}(r+o(r))=\frac{b_{2}}{c_{2}+b_{2}\ell_{0}}r+o(r),$ (A.13)
which is compatible with (4.62). Hence, by $\beta_{2}=b_{2}/c_{2}$,
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\ell^{(r)}_{0}\pi^{(r)}_{0}=\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}}.$
It remains to verify (3.4) and (3.5). This can be done through computing the limit of the distribution function $H(x)$ below as $r\downarrow 0$.
$H(x)\equiv\mathbb{P}[rL^{(r)}\leq x]=\sum_{n=0}^{[x]}\pi^{(r)}_{n},\qquad x\in\mathbb{R}_{+}.$
However, we omit the verification of (3.4) and (3.5) because they are part (i) of Theorem 3.1.
A.3 Proofs of Lemma 4.4
We first prove (4.22) and (4.23). Apply $f({\boldsymbol{x}})=x_{2}^{k+1}1(x_{2}\in B)$ to (4.1), then
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})1(T^{(r)}_{e,1}\in B)]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})1(T^{(r)}_{e,2}\in B)]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
where (4.1) is still applicable since the right-hand side of this equation is finite. Hence, (4.22) is obtained. (4.23) is similarly proved. We next prove (4.24).
$|(x\wedge 1/r)^{k}-x^{k}|$ $\leq x^{k}1(x>1/r)<rx^{k+1}1(x>1/r)$
$<r^{2}x^{k+2}1(x>1/r),\qquad x\geq 0,k\geq 1.$ (A.14)
Hence, the last inequality in (A.3) for $k=1$ and the uniform integrability condition yield
$|\mathbb{E}[\widehat{T}^{(r)}_{u}]-\mathbb{E}[T^{(r)}_{u}]|\leq r^{2}\mathbb{E}[(T^{(r)}_{u})^{3}1(T^{(r)}_{u}>1/r)]=o(r^{2}).$ (A.15)
On the other hand, the second inequality in (A.3) yields
$|\mathbb{E}[(\widehat{T}^{(r)}_{u})^{2}]-\mathbb{E}[(T^{(r)}_{u})^{2}]|\leq r\mathbb{E}[(T^{(r)}_{u})^{2}1(T^{(r)}_{u}>1/r)]=o(r).$ (A.16)
Thus, (A.15) and (A.16) prove (4.24) for $k=1,2$. Similarly, it follows from (4.22) with $B=(1/r,\infty)$ that, for $k=1,2$,
$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|$ $\leq\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}>1/r)]$
$=\frac{1}{k+1}\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}>1/r)]=o(r^{2-k}),$
To prove (4.26), we apply $f({\boldsymbol{x}})=x_{2}^{k}1(x_{2}>1/r,x_{3}>0)$ to (4.1), then
$k\mathbb{E}[(R^{(r)}_{e})^{k-1}1(R^{(r)}_{e}>1/r)]$ $=\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)]$
$\quad+\sum_{i=1,2}\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,i})^{k}1(T^{(r)}_{e,i}>1/r)]\mathbb{P}^{(r)}_{e}[w_{i}(L^{(r)}(0-))]$
$\geq\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)].$
Hence, applying (4.22) with $B=(1/r,\infty)$ and $k-1$ instead of $k$ to the above formula, we have
$|\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e})^{k}]|$ $\leq\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)]$
$\leq(\alpha^{(r)}_{e})^{-1}k\mathbb{E}[(R^{(r)}_{e}(0))^{k-1}1(R^{(r)}_{e}>1/r)]$
$\leq k\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k}1(T^{(r)}_{e,i}>1/r)],$
where the last term is $o(r^{2})$ for $k=1$ and $o(r)$ for $k=2$ by (A.15) and (A.16). Thus, we have (4.26). (4.27) is similarly obtained.
A.4 Proofs of Lemmas 4.7 and 4.8
Proof of Lemma 4.7.
From Taylor expansion of twice differentiable function $f$:
$|f(x)-(f(0)+f^{\prime}(0)x)|\leq\frac{1}{2}|x|^{2}\max_{|u|\leq|x|}|f^{\prime\prime}(u)|,\qquad x\in\mathbb{R},$ (A.17)
we have $|e^{-x}-1+x|\leq\frac{1}{2}|x|^{2}e^{|x|}$. Hence, for $i=1,2$,
$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}(\ell_{0}+o(r))}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}o(r)),$
$e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}(\ell_{0}+r\theta_{i}+o(r))}=e^{r\theta_{i}\ell^{(r)}_{0}}(1+\theta_{i}r+\theta_{i}^{2}o(r)).$
Thus, (4.35) is obtained. Using these asymptotic expansions, we have
$e^{r\theta_{2}(\ell^{(r)}_{0}+1)}-e^{r\theta_{1}(\ell^{(r)}_{0}+1)}$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+e^{r\theta^{(r)}_{2}\ell^{(r)}_{0}}(\theta_{2}r+\theta_{2}^{2}o(r))-e^{r\theta^{(r)}_{1}\ell^{(r)}_{0}}(\theta_{1}r+\theta_{1}^{2}o(r))$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+r(e^{r\theta^{(r)}_{2}\ell^{(r)}_{0}}\theta_{2}-e^{r\theta^{(r)}_{1}\ell^{(r)}_{0}}\theta_{1})+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}-e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r)$
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+r(e^{\theta_{2}\ell_{0}}(1+\theta_{2}o(r))\theta_{2}-e^{\theta_{1}\ell_{0}}(1+\theta_{1}o(r))\theta_{1})$
$\qquad+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}-e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r)$
$\quad=e^{r\theta_{2}\ell_{0}}-e^{r\theta_{1}\ell_{0}}-r(e^{\theta_{2}\ell_{0}}\theta_{2}-e^{\theta_{1}\ell_{0}}\theta_{1})+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}+e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r).$
This proves (4.36), and (4.37) is similarly proved. ∎
Proof of Lemma 4.8.
It follows from Lemma 4.2 that
$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}=1-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}+\frac{1}{2}|\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}|^{2}e^{|\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}|}$
$\quad=1+(\mu r\theta_{2}+O(r^{2}))\widehat{R}^{(r)}_{s}+O(r^{2})(\widehat{R}^{(r)}_{s})^{2}$
$\quad=1+\mu r\theta_{2}\widehat{R}^{(r)}_{s}+O(r^{2})(\widehat{R}^{(r)}_{s})^{2}.$
This proves (4.39), and (4.8) and (4.8) are similarly proved. ∎

### v3

@@
 Keywords: single server queue, level dependent arrival and service, stationary distribution, heavy traffic, diffusion approximation, Palm distribution, process limit, stochastic integral equation.
+Note: This version corrects Section 5.1 in addition to fixing typos.
 11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan & School of Data Science, Chinese University of Hong Kong, Shenzhen, China.
@@
 $\lim_{r\downarrow 0}(1-\rho^{(r)})\mathbb{E}[L^{(r)}]=\lim_{r\downarrow 0}\mathbb{P}(L^{(r)}=0)\mathbb{E}[L^{(r)}]$
-=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{2}\not=0.\end{cases}$ (3.13)
+=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.13)
 Let consider Theorem 3.1 when the inter-arrival times and service workload are all exponentially distributed. We call this case exponential.
@@
 5 Concluding remarks
-In this section, we discuss two subjects, a process limit for the 2-level $GI/G/1$ queue in heavy traffic and a verification of (3.11) in Remark 3.1.
+In this section, we intuitively answer two questions, what is a process limit for the 2-level $GI/G/1$ queue in heavy traffic and how to verify (3.11) in Remark 3.1.
 5.1 Process limit
@@
 From the assumptions (3.a)–(3.e), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
-Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation:
+Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
 $\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})$
@@
 which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
-We like to identify this stationary distribution $\widetilde{\nu}$ by that of the reflecting diffusion process.
-Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$, and define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
+For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$.
+So, we only consider the case that $b_{1}=0$.
+We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflecting diffusion process.
+Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$.
+We aim to characterize this $Z(\cdot)$ by a stochastic integral equation.
+To this end, define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
 $\sigma(x)=\sqrt{c_{1}}\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\sigma_{2}1(x>\ell_{0}),$
@@
 If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [7] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$.
-However, $Y(\cdot)$ may not be absolutely continuous.
-So far, assume that there is a weak solution of (5.2).
+However, $Y(\cdot)$ is nondecreasing, but may not be absolutely continuous.
+So far, we just assume that there is a weak solution of (5.2).
 For notational convenience, we continue to use the same notations for this weak solution.
@@
 […]
-then $M(\cdot)$ is a martingale. Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$. Then, it follows from (5.2) that
+then $M(\cdot)$ is a martingale.
+Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$.
+Then, $Z(t)$ and $Y(t)$ are continuous in $t$, so it follows from (5.2) that
 […]
-Hence, by Ito integration rule, we have, for $f\in C^{2}_{b}(\mathbb{R}_{+})$, where $C^{2}_{b}(\mathbb{R}_{+})$ is the set of all bounded and twice differentiable function from $\mathbb{R}_{+}$ to $\mathbb{R}$ which have the derivatives from the right at the origin, we have
+We next define local time $L_{x}(t)$ for $x\geq 0$ as
 […]
-where $\Gamma$ is operator on the space $C^{2}_{b}(\mathbb{R}_{+})$ defined as
-$\Gamma f(x)=-b(x)f^{\prime}(x)+\frac{1}{2}\sigma^{2}(x)f^{\prime\prime}(x),\qquad x\geq 0.$
-In particular, let $f$ be
-$f(x)=e^{\theta_{1}x}1(0\leq x\leq\ell_{0})+e^{\theta_{2}x}1(x>\ell_{0}),\qquad\theta_{i}\leq 0,i=1,2.$ (5.6)
-If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.5) for this $f$ yields
-$\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})+\theta_{1}\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$
-which implies that $\widetilde{\varphi}_{1}(\theta_{1})$ is the moment generating function of the exponential distribution and $\widetilde{\varphi}_{2}(\theta_{2})=0$ for all $\theta_{2}\leq 0$.
-Obviously, this is absurd.
-Thus, there is no solution $Z(\cdot)$ for (5.2) if $Z(\cdot)$ is a stationary process.
-What is a right stochastic integral equation for $Z(\cdot)$ Although $Z(\cdot)$ is not reflected at $\ell_{0}$, we may consider the discontinuous change of $\sigma(Z(u))$ at $Z(u)=\ell_{0}$ as something like reflection.
-This motivates us to define local times $L_{\ell_{0}}^{-}(t)$ and $L_{\ell_{0}}^{+}(t)$ at $\ell_{0}$ (e.g.
-see Theorem 22.5 of [6] for the definition of a local time) by
+where ${\rm sgn}(x-)=1(x>0)-1(x\leq 0)$ (e.g., see (4) of Chapter 22 of [6]). Note that, for any measurable function $f\geq 0$,
+[…]
+by Theorem 22.5 of [6].
+Note that $L_{x}(t)$ is $a.s.$ continuous in $t\geq 0$, but may not be so in $x$ if $Z(\cdot)$ is not a martingale.
+So, let $L^{+}_{x}(t)$ and $L^{-}_{x}(t)$ be the right and left continuous versions on $x$, respectively, of $L_{x}(t)$ for each fixed $t>0$.
+Namely, $L^{+}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x+\varepsilon}(t)$ and $L^{-}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x-\varepsilon}(t)$.
+As we will see, their difference is a key for our analysis.
+To derive a BAR for (5.2), we use the extended Ito formula for a convex test function $f$ for the local time $L^{-}_{x}(t)$, which is given as
+[…]
+where $f^{\prime}_{+}(x)=f^{\prime}(x+)$, that is, left derivative of $f$ at $x$, and $\mu_{f^{\prime}_{+}}$ is a measure on $(\mathbb{R}^{+},{\mathcal{B}}(\mathbb{R}_{+}))$ defined by
+$\mu_{f^{\prime}_{+}}((x,y])=f^{\prime}_{+}(x)-f^{\prime}_{+}(y),\qquad x<y\mbox{ in }\mathbb{R}_{+}.$
+Note that (5.7) is the left-continuous local time counterpart of Theorem 22.5 of [6], which takes the right-continuous local time $L^{+}_{x}(t)$ and $\mu_{f^{\prime}_{-}}$ instead of $L^{-}_{x}(t)$ and $\mu_{f^{\prime}_{+}}$.
+For (5.7), we take the test function $f$ with parameter $\theta\geq 0$, which is defined as $f(x)=e^{\theta h^{-}_{\ell_{0}}(x)}$, where $h^{-}_{\ell_{0}}(x)=(\ell_{0}-x)1(0\leq x\leq\ell_{0})$.
+Then, $f$ is a convex function for $x\geq 0$, and $f^{(n)}_{+}(x)=\theta^{n}e^{\theta x}1(0\leq x<\ell_{0})$ for $n\geq 1$.
+Since $\mu_{f^{\prime}_{+}}(\{\ell_{0}\})=-f^{\prime}(\ell_{0}-)$ and $\mu_{f^{\prime}_{+}}$ has the density $f^{\prime\prime}_{+}(x)=\theta^{2}e^{\theta x}1(0\leq x<\ell_{0})$, it follows from (5.4) and (5.6) that
+[…]
+$=\frac{1}{2}\theta^{2}\int_{0}^{t}e^{\theta Z(u)}1(0\leq Z(u)<\ell_{0})d[Z]_{u}-\frac{1}{2}\theta e^{\theta\ell_{0}}L^{-}_{\ell_{0}}(t).$
+Hence, the extended Ito formula (5.7) becomes
 […]
 […]
-and replace (5.2) by
+If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.1) for $t=1$ yields
+$\Big{(}-b_{1}\theta+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{1}(\theta)-\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{-}_{\ell_{0}}(1)]+\theta\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$ (5.9)
+where $\widetilde{\varphi}_{1}(\theta)$ is the moment generating function of $Z(1)$ on $0\leq Z(1)\leq\ell_{0}$.
+Similarly, let $\widetilde{\varphi}_{2}(\theta)$ be the moment generating function of $Z(1)$ on $Z(1)>\ell_{0}$, then, taking the right-continuous version $L^{+}_{x}(t)$ of the local time $L_{x}(t)$, and letting $f(x)=-e^{\theta h^{+}_{\ell_{0}}}$ for $\theta\leq 0$, where $h^{+}_{\ell_{0}}=(x-\ell_{0})1(x>\ell_{0})$, we have
+$\Big{(}-b_{2}\theta+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{2}(\theta)+\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{+}_{\ell_{0}}(1)]=0.$ (5.10)
+We finally note that
 […]
-Then, applying Ito integration rule to this stochastic integral equation for $f$ of (5.6) and taking for $Z(\cdot)$ to be stationary yield (5.1). This suggests the following conjectures.
+[…]
+Then, combining (5.9) and (5.10), we arrive at (5.1). Thus, the BAR of $Z(\cdot)$ is identical with (5.1). Based on this fact, we conjecture that
   • (a)
-    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to $Z(\cdot)$ satisfying (5.9) as $r\downarrow 0$.
+    there is a unique weak solution $Z(\cdot)$ for the stochastic integral equation (5.2),
   • (b)
-    Stochastic integral equation (5.9) has a unique weak solution.
-Both are hard to verify, so we leave them for future work.
-Since $Z(\cdot)$ of (5.9) is specified by $b_{i},c_{i},\sigma_{i}^{2}$ for $i=1,2$ and $\ell_{0}$, its stationary distribution must be $\widetilde{\nu}$ of (3.7).
-This supports the conjecture in Remark 3.1 because of (ii) of Theorem 3.1.
+    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to the $Z(\cdot)$ of (a),
+  • (c)
+    The $\widetilde{\nu}$ of (3.7) in Theorem 3.1 is the unique solution of the BAR (5.1) and the stationary distribution of $Z(t)$.
+These conjectures are challenging, but we expect that they are verifiable.
 5.2 Intuitive verification of (3.11)

### v4

@@
 This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed.
-To get its performance in tractable form, we study in the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation.
+To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation.
 It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters.
@@
 Keywords: single server queue, level dependent arrival and service, stationary distribution, heavy traffic, diffusion approximation, Palm distribution, process limit, stochastic integral equation.
-Note: This version corrects Section 5.1 in addition to fixing typos.
-11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan & School of Data Science, Chinese University of Hong Kong, Shenzhen, China.
+11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan.
 1 Introduction
@@
 However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential.
-Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of this 2-level $GI/G/1$ queues under a certain scaling in heavy traffic.
+Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution.
 This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold.
@@
 We take the stochastic framework introduced in [11] to study the 2-level $GI/G/1$ queue.
-Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining service time, respectively, at time $t$.
+Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining workload of a customer being served, respectively, at time $t$.
+Note that the remaining service time cannot be used for $R_{s}(t)$ here because the service rate may change in future.
 Define
@@
 Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with respect to the time-shift operator $\Psi$.
-We assume the following assumptions.
+Let $b_{1}\in\mathbb{R}$ and $b_{2},\ell_{0}>0$ be constants, and assume the following assumptions.
   • (3.a)
@@
   • (3.d)
-    $r\ell^{(r)}_{0}=\ell_{0}+o(r)$, and $\ell^{(r)}_{0}\geq 2$.
+    $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
   • (3.e)
@@
     $\{(T^{(r)}_{e,i})^{2};r\in(0,1]\}$ for $i=1,2$ and $\{(T^{(r)}_{s})^{2};r\in(0,1]\}$ are uniformly integrable.
-Furthermore, (3.b) and (3.c) imply that $1-\rho^{(r)}_{i}=rb_{i}/c_{i}+o(r)$, $\rho_{i}\equiv\frac{\lambda_{i}}{c_{i}\mu}=1$ and $\lambda_{i}=c_{i}\mu$.
+Furthermore, define $\rho_{i}=\lambda_{i}/(c_{i}\mu)$, then (3.b) and (3.c) imply that $1-\rho^{(r)}_{i}=rb_{i}/c_{i}+o(r)$, $\rho_{i}=1$ and $\lambda_{i}=c_{i}\mu$.
 Hence, $\lambda_{1}=\lambda_{2}$ if and only if $c_{1}=c_{2}$.
-It is easy to see that $X^{(r)}(t)$ has the stationary distribution by (3.e).
-This distribution is denoted by $\nu^{(r)}$.
 We take countably many $r$’s for a limiting operation.
@@
 In what follows, $\mathbb{P}$ is always this probability measure.
-By the stability condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution.
+By the condition (3.e), the Markov process $X^{(r)}(\cdot)$ has the stationary distribution.
 Let $X^{(r)}\equiv(L^{(r)},R^{(r)}_{e},R^{(r)}_{s})$ be a random vector subject to the stationary distribution of $X^{(r)}(t)$.
 For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$.
-For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
+For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
 Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$, and let $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper.
@@
     For each $i=1,2$, if $\nu^{(r)}_{i}(\mathbb{R}_{+})$ is bounded away from 00 as $r\downarrow 0$, then $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$, namely,
-    $\lim_{r\downarrow 0}\nu_{i|cd}^{(r)}(B)=\widetilde{\nu}_{icd}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.6)
+    $\lim_{r\downarrow 0}\nu_{i|cd}^{(r)}(B)=\widetilde{\nu}_{i|cd}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.6)
   • (ii)
@@
 Let $S^{(r)}(n)$ be the sojourn time of the $n$-th arriving customer in service, then the traffic intensity $\rho^{(r)}$ can be defined under the stationary setting as
-$\rho^{(r)}=\alpha^{(r)}_{e}\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}S^{(r)}_{k}.$ (3.12)
+$\rho^{(r)}=\alpha^{(r)}_{e}\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}S^{(r)}(k).$ (3.12)
 Hence, by Little’s law, we have $\rho^{(r)}=\mathbb{P}(L^{(r)}>0)=1-\mathbb{P}(L^{(r)}=0)$.
@@
 $\lim_{r\downarrow 0}(1-\rho^{(r)})\mathbb{E}[L^{(r)}]=\lim_{r\downarrow 0}\mathbb{P}(L^{(r)}=0)\mathbb{E}[L^{(r)}]$
-=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.13)
+=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.13)
 Let consider Theorem 3.1 when the inter-arrival times and service workload are all exponentially distributed. We call this case exponential.
 Proposition 3.1.
-Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{M}$ which is defined as $\widetilde{\nu}_{M}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{1}(M)+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{2}(M)$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
-$\widetilde{A}_{1}(M)=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
-and $\widetilde{A}_{2}(M)=1-\widetilde{A}_{1}(M)$.
+Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{\exp}$ which is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
+$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
+and $\widetilde{A}_{\exp,2}=1-\widetilde{A}_{\exp,1}$.
 Remark 3.2.
-Since $\sigma_{i}^{2}=2$ and $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$ in the exponential case, $\widetilde{\nu}_{M}$ and $\widetilde{A}_{i}(M)$ agree with $\widetilde{\nu}$ and $\widetilde{A}_{i}$ in Theorem 3.1, respectively, for $i=1,2$.
+Since $\sigma_{i}^{2}=2$ and $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$ in the exponential case, $\widetilde{\nu}_{\exp}$ and $\widetilde{A}_{\exp,i}$ agree with $\widetilde{\nu}$ and $\widetilde{A}_{\exp,i}$ in Theorem 3.1, respectively, for $i=1,2$.
 Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed.
@@
 In this section we prove Theorem 3.1.
-Compared with the exponential case, a major difficulty comes from the fact that $Z^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$.
-Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions.
+Compared with the exponential case, a major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$.
+Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below.
 This method is called a BAR approach, and has been used in [2, 3, 10].
 Recently, this approach is further studied in [11].
-The proof of Theorem 3.1 is composed of three parts. We first introduce Palm distributions and the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
+The proof of Theorem 3.1 is composed of three parts. We first introduce the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
 4.1 Palm distribution and BAR
@@
 $\mathbb{P}^{(r)}_{v}(A)=(\alpha^{(r)}_{v})^{-1}\mathbb{E}\left[\int_{(0,1]}1_{\Theta_{u}^{-1}A}N^{(r)}_{v}(du)\right],\qquad A\in{\mathcal{F}},v=e,s,$ (4.1)
-where $1_{A}$ is the indicator function of the event $A\in{\mathcal{F}}$.
+where $1_{A}$ is the indicator function of the event $A\in{\mathcal{F}}$. For $v=e,s$, $\mathbb{P}^{(r)}_{v}$ is called a Palm distribution concerning the counting process $N^{(r)}_{v}$.
+The Palm distributions play a major role for the derivation of the stationary equations in our analysis (e.g., see (4.1)).
+Intuitively, $\mathbb{P}^{(r)}_{v}$ is considered as the conditional probability distribution given the event that $N^{(r)}_{v}$ counts arrivals for $v=e$ (service completion for $v=s$) at time 00.
+This event occurs with probability 00 under the stationary framework, so such an event is often handled as a boundary condition in the stationary equation (e.g, see [4]).
+On the other hand, Palm distributions are defined through time averages under the stationary framework.
+This enables us to derive the stationary equation without any boundary condition, where the boundary conditions are handled by Palm distributions in the stationary equation.
+This is quite convenient because the boundary conditions may be only asymptotically satisfied.
+This is exactly our case.
 We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [11].
 For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$.
-Let $S$ be its state space, that is, $S\equiv\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right.
+Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right.
 Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
 ${\mathcal{H}}f({\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f({\boldsymbol{x}})-(c_{1}1(1\leq x_{1}\leq\ell_{0})+c_{2}1(x_{1}>\ell_{0}))\frac{\partial^{+}}{\partial x_{3}}f({\boldsymbol{x}}),\qquad{\boldsymbol{x}}\in S,$
-where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$, and define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
+where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$.
+In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit.
+Define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
 $X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
@@
 Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N_{s}(t)))$ when $\Delta N_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N_{s}(t)=0$.
-We further define difference operators $\Delta_{e}$ and $\Delta_{s}$ as
+For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
 $\Delta_{e}f(X^{(r)}(t))=f(X^{(r)}_{1}(t))-f(X^{(r)}(t-)),$
-$\Delta_{s}f(X^{(r)}(t))=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)),$
+$\Delta_{s}f(X^{(r)}(t))=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)).$
 Then, from the elementary observation of the time evolution of $X^{(r)}(t)$, we have
@@
 $\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
-To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{s}$.
+To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$.
 We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
 4.2 Preliminary lemmas
-We will prove Theorem 3.1 using the moment generating functions $\varphi_{i}(\theta_{2})$ and $\varphi_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$.
+We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{2})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$.
 Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
-$\varphi_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
-$\varphi_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
+$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
+$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
 To this end, we introduce test functions for the BAR (4.1). We first define functions $w^{(r)}_{i}$ and $g^{(r)}_{i,\theta}$ for $i=1,2$ and $\theta\in\mathbb{R}$ as
@@
 $e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1.$ (4.8)
-We note the following lemma, which is obtained from Lemma 5.8 of [2].
+Note that $\eta^{(r)}_{i}(\theta)$ is the value of the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ at $e^{-\theta}$, so it is infinitely many differentiable function of $\theta$.
+Obviously, $\zeta^{(r)}(\theta)$ has the same property.
+Hence, twice differentiating the equations in (4.8), we can compute the Taylor expansions of them around the origin, which yields the following lemma.
+It is formally obtained in [3, Lemma 5.8].
 Lemma 4.2.
@@
 $\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$
-Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L(^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L(^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
+Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
 $\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
@@
 For $k\geq 1$, we first prove (4.5). We apply $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
-$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,L^{(r)}(t)\leq\ell^{(r)}_{0}).$
+$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,1\leq L^{(r)}(t)\leq\ell^{(r)}_{0}).$
 Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ and $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, respectively.
@@
 Recall that
-$\Delta^{(r)}_{i}=\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
+$\Delta^{(r)}_{i}=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
 $\hskip 43.05542pt-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}],\qquad i=1,2,$
@@
 Hence, (4.66) holds. Applying it to (4.61) and (4.64) for $i=1$, we obtain
-$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b\not=0,\end{cases}$ (4.68)
+$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (4.68)
 and, similarly, for $i=2$, (4.55) yields
@@
 From the assumptions (3.a)–(3.e), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
-Recall that one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
+Recall that, if one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
 $\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})$
@@
 For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$.
-So, we only consider the case that $b_{1}=0$.
+So, we only consider the case that $b_{1}\not=0$.
 We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflecting diffusion process.
@@
 If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [6] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$.
-However, $Y(\cdot)$ is nondecreasing, but may not be absolutely continuous.
-So far, we just assume that there is a weak solution of (5.2).
-For notational convenience, we continue to use the same notations for this weak solution.
-To apply Ito integration rule to (5.2), let us compute the quadratic variation $\big{[}Z\big{]}_{t}$ of $Z(\cdot)$.
-Define $M(\cdot)\equiv\{M(t);t\geq 0\}$ as
-[…]
-then $M(\cdot)$ is a martingale.
-Denote the quadratic variations of $Z(\cdot)$ and $M(\cdot)$, respectively, by $\big{[}Z\big{]}_{t}$ and $\big{[}M\big{]}_{t}$.
-Then, $Z(t)$ and $Y(t)$ are continuous in $t$, so it follows from (5.2) that
-[…]
-We next define local time $L_{x}(t)$ for $x\geq 0$ as
-[…]
-where ${\rm sgn}(x-)=1(x>0)-1(x\leq 0)$ (e.g., see (4) of Chapter 22 of [6]). Note that, for any measurable function $f\geq 0$,
-[…]
-by Theorem 22.5 of [6].
-Note that $L_{x}(t)$ is $a.s.$ continuous in $t\geq 0$, but may not be so in $x$ if $Z(\cdot)$ is not a martingale.
-So, let $L^{+}_{x}(t)$ and $L^{-}_{x}(t)$ be the right and left continuous versions on $x$, respectively, of $L_{x}(t)$ for each fixed $t>0$.
-Namely, $L^{+}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x+\varepsilon}(t)$ and $L^{-}_{x}(t)=\lim_{\varepsilon\downarrow 0}L_{x-\varepsilon}(t)$.
-As we will see, their difference is a key for our analysis.
-To derive a BAR for (5.2), we use the extended Ito formula for a convex test function $f$ for the local time $L^{-}_{x}(t)$, which is given as
-[…]
-where $f^{\prime}_{+}(x)=f^{\prime}(x+)$, that is, left derivative of $f$ at $x$, and $\mu_{f^{\prime}_{+}}$ is a measure on $(\mathbb{R}^{+},{\mathcal{B}}(\mathbb{R}_{+}))$ defined by
-$\mu_{f^{\prime}_{+}}((x,y])=f^{\prime}_{+}(x)-f^{\prime}_{+}(y),\qquad x<y\mbox{ in }\mathbb{R}_{+}.$
-Note that (5.7) is the left-continuous local time counterpart of Theorem 22.5 of [6], which takes the right-continuous local time $L^{+}_{x}(t)$ and $\mu_{f^{\prime}_{-}}$ instead of $L^{-}_{x}(t)$ and $\mu_{f^{\prime}_{+}}$.
-For (5.7), we take the test function $f$ with parameter $\theta\geq 0$, which is defined as $f(x)=e^{\theta h^{-}_{\ell_{0}}(x)}$, where $h^{-}_{\ell_{0}}(x)=(\ell_{0}-x)1(0\leq x\leq\ell_{0})$.
-Then, $f$ is a convex function for $x\geq 0$, and $f^{(n)}_{+}(x)=\theta^{n}e^{\theta x}1(0\leq x<\ell_{0})$ for $n\geq 1$.
-Since $\mu_{f^{\prime}_{+}}(\{\ell_{0}\})=-f^{\prime}(\ell_{0}-)$ and $\mu_{f^{\prime}_{+}}$ has the density $f^{\prime\prime}_{+}(x)=\theta^{2}e^{\theta x}1(0\leq x<\ell_{0})$, it follows from (5.4) and (5.6) that
-[…]
-$=\frac{1}{2}\theta^{2}\int_{0}^{t}e^{\theta Z(u)}1(0\leq Z(u)<\ell_{0})d[Z]_{u}-\frac{1}{2}\theta e^{\theta\ell_{0}}L^{-}_{\ell_{0}}(t).$
-Hence, the extended Ito formula (5.7) becomes
-[…]
-[…]
-If $Z(\cdot)$ is a stationary process, then taking the expectation of (5.1) for $t=1$ yields
-$\Big{(}-b_{1}\theta+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{1}(\theta)-\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{-}_{\ell_{0}}(1)]+\theta\widetilde{\varphi}_{1}(0)\mathbb{E}[Y(1)]=0,$ (5.9)
-where $\widetilde{\varphi}_{1}(\theta)$ is the moment generating function of $Z(1)$ on $0\leq Z(1)\leq\ell_{0}$.
-Similarly, let $\widetilde{\varphi}_{2}(\theta)$ be the moment generating function of $Z(1)$ on $Z(1)>\ell_{0}$, then, taking the right-continuous version $L^{+}_{x}(t)$ of the local time $L_{x}(t)$, and letting $f(x)=-e^{\theta h^{+}_{\ell_{0}}}$ for $\theta\leq 0$, where $h^{+}_{\ell_{0}}=(x-\ell_{0})1(x>\ell_{0})$, we have
-$\Big{(}-b_{2}\theta+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta^{2}\Big{)}\widetilde{\varphi}_{2}(\theta)+\frac{1}{2}\theta e^{\theta\ell_{0}}\mathbb{E}[L^{+}_{\ell_{0}}(1)]=0.$ (5.10)
-We finally note that
-[…]
-[…]
-Then, combining (5.9) and (5.10), we arrive at (5.1). Thus, the BAR of $Z(\cdot)$ is identical with (5.1). Based on this fact, we conjecture that
+However, $Y(\cdot)$ may not be absolutely continuous.
+So far, in the previous versions of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function.
+We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1.
+So, it is conjectured in the previous versions that
   • (a)
@@
   • (c)
-    The $\widetilde{\nu}$ of (3.7) in Theorem 3.1 is the unique solution of the BAR (5.1) and the stationary distribution of $Z(t)$.
-These conjectures are challenging, but we expect that they are verifiable.
+    the stationary distribution of $Z(t)$ uniquely exists under the stability condition $b_{2}>0$, and agrees with $\widetilde{\nu}$ of (3.7) in Theorem 3.1 under the one of conditions in (ii) of it in addition to $b_{2}>0$.
+During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al.
+[1, Lemma 4.3] (see also [12] for a different proof), and (c) is affirmatively answered in [12].
+Furthermore, the stationary distribution of $Z(t)$ which is obtained in [12] does not need any of the extra conditions in Theorem 3.1.
+This is not surprising because limiting operation for $r\downarrow 0$ is not needed to derive the stationary distribution of $Z(t)$.
+Thus, we need not further discuss about the conjectures (a) and (c).
+As for the conjecture (b), it is still open (see [12]).
 5.2 Intuitive verification of (3.11)
@@
 Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, we have
-$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1).$ (5.13)
+$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (5.3)
+where $\sim$” stands for the same order as $r\downarrow 0$, that is, for non-zero valued functions $f_{1}$ and $f_{2}$, $f_{1}(r)\sim f_{2}(r)$ means that $\lim_{r\downarrow 0}f_{1}(r)/f_{2}(r)=1$.
 On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below.
@@
 $=\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))=\frac{b_{1}}{c_{1}}\frac{c_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}r+o(r),$ (A.12)
-which is compatible with (4.59). Hence, (A.9) yield
+which is compatible with (4.59). It follows from (A.9) and (A.2) that
 $\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})$ $=\lim_{r\downarrow 0}\frac{1-e^{-\beta_{1}\ell_{0}}+o(1)}{\beta_{1}r+o(r)}\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))$
-$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}.$
-Thus, this and $\widetilde{A}_{1}+\widetilde{A}_{2}=1$ concludes (4.35).
+$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)+c_{2}b_{1}},$ (A.13)
+where the last equality uses the fact that $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$.
 We next consider the case that $b_{1}=0$. In this case, $\rho^{(r)}_{1}=1-o(r)$, and therefore, from (A.6),
 $\pi^{(r)}_{0}$ $=\frac{\beta_{2}}{1+\beta_{2}\ell_{0}}(r+o(r))=\frac{b_{2}}{c_{2}+b_{2}\ell_{0}}r+o(r),$ (A.14)
-which is compatible with (4.62). Hence, by $\beta_{2}=b_{2}/c_{2}$,
-$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\ell^{(r)}_{0}\pi^{(r)}_{0}=\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}}.$
-It remains to verify (3.4) and (3.5). This can be done through computing the limit of the distribution function $H(x)$ below as $r\downarrow 0$.
-$H(x)\equiv\mathbb{P}[rL^{(r)}\leq x]=\sum_{n=0}^{[x]}\pi^{(r)}_{n},\qquad x\in\mathbb{R}_{+}.$
-However, we omit the verification of (3.4) and (3.5) because they are part (i) of Theorem 3.1.
+which is compatible with (4.62). Since $\beta_{2}=b_{2}/c_{2}$, (A.9) and (A.14) yield
+$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\ell^{(r)}_{0}\pi^{(r)}_{0}=\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}}.$ (A.15)
+Hence, from (3.14), (A.2) and (A.15) and using the fact that $\nu^{(r)}_{1}(\mathbb{R}_{+})+\nu^{(r)}_{2}(\mathbb{R}_{+})=1$, we have
+$\lim_{r\downarrow 0}\nu^{(r)}_{i}(\mathbb{R}_{+})$ $=\widetilde{A}_{\exp,i},\qquad i=1,2.$ (A.16)
+It remains to verify (3.4) and (3.5).
+This can be done through the Laplace transforms of $\{\pi^{(r)}_{n};0\leq n\leq\ell^{(r)}_{0}\}$ and $\{\pi^{(r)}_{n};n>\ell^{(r)}_{0}\}$ together with (A.16).
+We omit its detail because they are just computations.
 A.3 Proofs of Lemma 4.4
@@
 This proves (4.39), and (4.8) and (4.8) are similarly proved. ∎
+Acknowledgement
+I thank the editor and referees for their helpful comments and careful reading.
+This study is motivated by the BAR approach, coined by Jim Dai (e.g., see [3]).
+I am grateful to him for his grand design about the BAR approach, and to Rami Atar for providing his recent paper [1].
+I also benefit from Evsey Morozov for his comments on this paper.

### v5

@@
 This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed.
-To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under a certain scaling in heavy traffic, which corresponds to a diffusion approximation.
+To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under the scaling in heavy traffic which corresponds to a diffusion approximation.
 It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters.
@@
 Here, the system is an internet server, and its energy consumption depends on service rate.
-Then, the problem is to find an optimal level to minimize the operating energy of the server and customer’s waiting cost.
+Then, the problem is to find an optimal threshold to minimize the operating energy of the server and customer’s waiting cost.
 So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length.
-It may be considered to approximate using a quasi-birth-and-death process, and it is not hard to numerically compute its performance.
+It may be considered to approximate it by a quasi-birth-and-death process, then it is not hard to numerically compute its performance.
 However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential.
@@
 Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it.
-This conjecture is easily verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional.
+This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional.
 Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [2, 3, 10, 11].
-This verification, Theorem 3.1, requires a certain limitation on system parameters or the exponentially distributed workload brought by a customer.
+This verification, Theorem 3.1, requires either some limitation on system parameters or for the workload brought by a customer to be exponentially distributed.
 The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
@@
 We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1.
-A main result, Theorem 3.1, and the exponentially distributed case are presented in Section 3.2.
+A main result, Theorem 3.1, and the exponentially distributed case, Proposition 3.1, are presented in Section 3.2.
 In Section 4, we prove Theorem 3.1.
@@
 We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$.
-Thus, the inter-arrival times are independent, but they have may two different distributions depending on the queue length.
+Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length.
 Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are random variables which may have different distributions.
@@
 These dynamics is depicted in the sample path of $L^{(r)}(t)$ below.
+Figure 3.1: Sample path of the queue length process
 We will use the following notations.
@@
 Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov.
-Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with respect to the time-shift operator $\Psi$.
-Let $b_{1}\in\mathbb{R}$ and $b_{2},\ell_{0}>0$ be constants, and assume the following assumptions.
+Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with the time-shift operator $\Psi$.
+We assume the following assumptions.
   • (3.a)
@@
   • (3.c)
-    $c^{(r)}_{i}\mu^{(r)}-\lambda^{(r)}_{i}=r\mu^{(r)}b_{i}+o(r)$ as $r\downarrow 0$, and $b_{2}>0$.
+    There are constants $b_{1}\in\mathbb{R}$ and $b_{2}>0$ such that $c^{(r)}_{i}\mu^{(r)}-\lambda^{(r)}_{i}=r\mu^{(r)}b_{i}+o(r)$ as $r\downarrow 0$.
   • (3.d)
-    $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
+    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
   • (3.e)
     $\rho^{(r)}_{2}<1$.
+Here, we write $f(r)=o(g(r))$ as $r\downarrow 0$ for positive valued functions $f,g$ of $r$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$.
+Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$.
 Note that (3.a) implies that
@@
 Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$.
-Obviously, $\mathbb{E}[N_{e}(t)]\leq\mathbb{E}[N_{e,0}(t)]$ under the stationary framework.
-Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N_{e}(1)]\leq\mathbb{E}[N_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$.
-Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$.
+Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework.
+Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$.
+Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$.
 ∎
@@
 We are now ready to present a main result, which will be proved in Section 4.
-Recall that $L^{(r)}$ is a random variable subject to the stationary distribution $\nu^{(r)}$ of the $r$-th 2-level $GI/G/1$ queue.
-We decompose $\nu^{(r)}$ into marginal measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
+Recall that $L^{(r)}$ is a random variable subject to the stationary distribution of the $r$-th 2-level $GI/G/1$ queue.
+Define probability distribution $\nu^{(r)}$ under the assumptions (3.a)–(3.e) as
+$\nu^{(r)}(B)=\mathbb{P}(rL^{(r)}\in B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.1)
+We refer to this $\nu^{(r)}$ as a scaled stationary distribution.
+We decompose $\nu^{(r)}$ into measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
 $\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}>\ell^{(r)}_{0}),$
-$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}),$
+$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}).$
 Theorem 3.1.
@@
 $\sigma_{i}^{2}=\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2},\qquad\beta_{i}=\frac{2b_{i}}{c_{i}\sigma_{i}^{2}},\qquad i=1,2.$ (3.2)
-For each $r\in(0,1]$, define $\widetilde{\nu}_{i|cd}$ as the probability measures on ${\mathcal{B}}(\mathbb{R}_{+})$ which have the density functions $h_{i|cd}$ for $i=1,2$ which are
+For each $r\in(0,1]$, define $\widetilde{\nu}_{i|cd}$ as the probability measures on ${\mathcal{B}}(\mathbb{R}_{+})$ which have the density functions $h_{i|cd}$ for $i=1,2$ such that
 $h_{1|cd}(x)=1(0\leq x\leq\ell_{0})\times\left\{\begin{array}[]{ll}\displaystyle 1/\ell_{0},&b_{1}=0,\\ \displaystyle\frac{e^{\beta_{1}\ell_{0}}\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}e^{-\beta_{1}x},&b_{1}\not=0,\end{array}\right.$ (3.5)
@@
 Thus, (3.12) is equivalent to (3.11).
-We conjecture to have (3.10) without any condition in (ii).
+We conjecture that (3.11) holds without any condition in (ii).
 There are two supports for this conjecture.
@@
 =$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.14)
-Let consider Theorem 3.1 when the inter-arrival times and service workload are all exponentially distributed. We call this case exponential.
+Let consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
 Proposition 3.1.
-Under the settings of Theorem 3.1, if $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed, then, for any $\lambda_{1}$ and $\lambda_{2}$, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}_{\exp}$ which is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$, where
-$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.14)
+Under the settings of Theorem 3.1, the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
+$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{0}}{1+\beta_{2}\ell_{0}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.15)
 and $\widetilde{A}_{\exp,2}=1-\widetilde{A}_{\exp,1}$.
@@
 Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed.
-However, for a double check, we give its independent proof in Appendix A.2.
-This proof is straightforward because $L^{(r)}(\cdot)$ is a birth-and-death process, and helpful to intuitively convince (ii) of Theorem 3.1.
+However, for a double check, we give its direct proof in Appendix A.2, which is helpful to intuitively convince (ii) of Theorem 3.1.
+This proof is straightforward, but one must be careful because $L^{(r)}(\cdot)$ itself is not a continuous-time Markov chain even under the exponential assumptions (see Appendix A.2 for its reason).
 4 Proofs
 In this section we prove Theorem 3.1.
-Compared with the exponential case, a major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$.
+A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$.
 Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below.
@@
 Define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
-$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
-$L^{(r)}_{1}(t)=L(t-)+1(\Delta N_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
-where $n_{e,1}(t)=\sum_{n=1}^{N_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$.
+$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
+$L^{(r)}_{1}(t)=L(t-)+1(\Delta N^{(r)}_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
+where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$.
 We call $X^{(r)}_{1}(t)$ an intermediate state.
-Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N_{s}(t)))$ when $\Delta N_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N_{s}(t)=0$.
+Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$ when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
 For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
@@
   • (ii)
-    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N_{e}$ is counted at time 00 and just before when $N_{s}$ is counted at time 00.
+    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00.
 This means that we take a so called arrival first framework.
@@
 For each integer $k\geq 0$, let $f({\boldsymbol{x}})=x_{1}\wedge(k+1)$ for ${\boldsymbol{x}}\in S$, then obviously $f\in C_{b}^{p}(S)$.
-Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N_{e}(u)-N_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N_{s}(u)-N_{s}(u-)=1)$, we have
+Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
 $\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]$
@@
 $\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
-To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$.
+To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1.
 We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
 4.2 Preliminary lemmas
-We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{2})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$.
+We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{i})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$.
 Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
-$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}],$
+$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(L^{(r)}>\ell^{(r)}_{0})],$
 $\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
@@
 $w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{0}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{0}),\qquad z\in\mathbb{Z}_{+},$
-$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{1}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r)},\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
+$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{i}}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
 where $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are the unique solutions of the following equations for each $\theta,\theta\in\mathbb{R}$.
-$e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1.$ (4.8)
-Note that $\eta^{(r)}_{i}(\theta)$ is the value of the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ at $e^{-\theta}$, so it is infinitely many differentiable function of $\theta$.
-Obviously, $\zeta^{(r)}(\theta)$ has the same property.
-Hence, twice differentiating the equations in (4.8), we can compute the Taylor expansions of them around the origin, which yields the following lemma.
-It is formally obtained in [3, Lemma 5.8].
+$e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1,$ (4.8)
+because $\eta^{(r)}_{i}(\theta)$ is the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ taking the value $e^{-\theta}$ and $\zeta^{(r)}(\theta)$ is similarly determined.
+Furthermore, $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are infinitely many differentiable functions of $\theta$.
+Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin.
+It is formally proved in [3, Lemma 5.8].
 Lemma 4.2.
@@
 $\psi^{(r)}_{i}(\theta_{i})<\infty,\quad\mathbb{E}_{u}\left[f^{(r)}_{r{\boldsymbol{\theta}}}(L^{(r)},R^{(r)})\right]<\infty,\quad u=e,s,\quad\mbox{uniformly in $r\in(0,1]$},$ (4.15)
-because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$.
-These finiteness guarantees our computations bellow for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
-Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{{\boldsymbol{\theta}}}$. Then,
+because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
+Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
 $\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\sum_{i=1,2}\eta_{i}^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})$
@@
 $\qquad+e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.19)
-Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]\right)$.
+Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]\right)$.
 Then, by (4.2) and (4.2),
@@
 $(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
-and, for $k=1,2$,
-$\mathbb{E}[(\widehat{T}^{(r)}_{u})^{k}]=\mathbb{E}[(T^{(r)}_{u})^{k}]+o(r^{3-k}),\qquad u=(e,1),(e,2),s,$ (4.24)
-$\mathbb{E}[(\widehat{R}^{(r)}_{v})^{k}]=\mathbb{E}[(R^{(r)}_{v})^{k}]+o(r^{2-k}),\qquad v=e,s,$ (4.25)
+and, for $i=1,2$,
+$|\mathbb{E}[(\widehat{T}^{(r)}_{u})^{k}]-\mathbb{E}[(T^{(r)}_{u})^{k}]|=\mathbb{E}[(T^{(r)}_{u})^{k}1(T^{(r)}_{u}>1/r)]=o(r^{3-k}),\quad u=(e,i),s,$ (4.24)
+$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|=\mathbb{E}[(R^{(r)}_{v})^{k}1(R^{(r)}_{v}>1/r)]=o(r^{2-k}),\qquad v=e,s,$ (4.25)
 $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]=\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}]+o(r^{3-k}),$ (4.26)
 $\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}]=\mathbb{E}^{(r)}_{e}[(R^{(r)}_{s}(0-))^{k}]+o(r^{3-k}).$ (4.27)
-In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable.
+In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable for $k=1,2$.
 This lemma is easily proved using the BAR (4.1), but its proof is lengthy, so we defer it to Appendix A.3.
@@
 Proof.
-For $k=1$, multiply both sides of (4.5) and (4.5) by $\mu^{(r)}$ and $\lambda^{(r)}_{1}$, respectively, and take their difference.
-Since $c^{(r)}_{1}\mu^{(r)}-\lambda^{(r)}_{1}=r\mu^{(r)}b_{1}+o(r)$ by the assumption (3.c), $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.22) and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.23), we have
+For $k=1$, multiply both sides of (4.5) and (4.5) by $\mu^{(r)}$ and $\lambda^{(r)}_{1}$, respectively, and take their difference, then, noting that $c^{(r)}_{1}\mu^{(r)}-\lambda^{(r)}_{1}=r\mu^{(r)}b_{1}+o(r)$, we have
 $(r\mu^{(r)}b_{1}+o(r))\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]-c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]$
@@
 $\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
-because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1.
+because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.25).
 This proves (4.33).
@@
 Proof.
-(4.45) is immediate from Lemma 4.6 by applying (4.43) and (4.44). As for (4.46), we only prove its first equation because its second one is similarly proved. By the Schwartz’s inequality,
+The first equation of (4.45) is immediate from (4.2) and Lemma 4.6 together with (4.43) and (4.44), and the second equation follows from the first one by Lemma 4.1.
+As for (4.46), we only prove its first equation because its second one is similarly proved.
+By the Schwartz’s inequality,
 $\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}=\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3/2}(\widehat{R}^{(r)}_{s})^{1/2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}$
@@
 From (4.5) for $k=3$ and Lemma 4.4, we have
-$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad\leq\mathbb{E}[(T^{(r)}_{s})^{3}]+3c_{1}\mathbb{E}[(R^{(r)}_{s})^{2}1(R^{(r)}_{s}\geq 1/r,L^{(r)}\geq 1)]/\alpha^{(r)}$
-$\quad=\mathbb{E}[(T^{(r)}_{s})^{3}]+\frac{c_{1}}{\min_{i=1,2}c_{i}}\mathbb{E}[(T^{(r)}_{s})^{3}1(T^{(r)}_{s}\geq 1/r)],$
-which is uniformly bounded for $r\in(0.1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
+$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\leq\mathbb{E}[(T^{(r)}_{s})^{3}],$
+which is uniformly bounded for $r\in(0,1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
 (2nd Step)   In this step, we first expand the exponential terms in $D^{(r)}(\theta_{1},\theta_{2})$ of (4.2), using Lemmas 4.7 and 4.8 in the step 1, then
@@
 $\leq e^{|\theta_{1}|r\ell^{(r)}_{0}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
-Hence, (4.66) holds. Applying it to (4.61) and (4.64) for $i=1$, we obtain
+Hence, (4.66) holds. Applying it to (4.61) and (4.64), we obtain
 $\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (4.68)
-and, similarly, for $i=2$, (4.55) yields
+and, similarly, (4.55) yields
 $\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{0}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (4.69)
@@
 First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$.
-In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{e}$ under $\mathbb{P}^{(r)}_{s}$, so, by Lemma 4.4 for $k=1$,
-$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)$ $=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
+In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
+$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
 Applying this fact to the definitions of $\Delta^{(r)}_{i}$ yields
@@
 $\lambda_{2}\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
-Since $\mu c_{i}=\lambda_{i}$, this can be written as
+Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$, this can be written as
 $\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$
@@
 $\Delta^{(r)}_{1}-\Delta^{(r)}_{2}=(\lambda^{(r)}_{1}-\lambda^{(r)}_{2})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=o(r),$ (4.73)
-because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). On the other hand, if $b_{1}=0$, then (4.54) and (4.60) yield
+because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). If $b_{1}=0$, then (4.54) and (4.60) yield
 $\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-\Delta^{(r)}_{2})=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r),$
-Hence,
+Hence, by (4.3) and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$,
 $\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\mu b_{2}(1-\varphi^{(r)}_{1}(0))+o(1)$
-Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54) and (4.63) yield
+Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54), (4.63) and (4.3) yield
 $\frac{\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
-Again by $c_{1}=c_{2}$, this proves (4.72).
+Again by $c_{1}=c_{2}$ and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$, this proves (4.72).
 Let
@@
 5 Concluding remarks
-In this section, we intuitively answer two questions, what is a process limit for the 2-level $GI/G/1$ queue in heavy traffic and how to verify (3.11) in Remark 3.1.
+In this section, we intuitively discuss two questions, what is a process limit for the 2-level $GI/G/1$ queue in heavy traffic and how to verify (3.12) in Remark 3.1.
 5.1 Process limit
@@
 However, $Y(\cdot)$ may not be absolutely continuous.
-So far, in the previous versions of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function.
+So far, in the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function.
 We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1.
-So, it is conjectured in the previous versions that
+So, it was conjectured that
   • (a)
@@
 During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al.
-[1, Lemma 4.3] (see also [12] for a different proof), and (c) is affirmatively answered in [12].
-Furthermore, the stationary distribution of $Z(t)$ which is obtained in [12] does not need any of the extra conditions in Theorem 3.1.
-This is not surprising because limiting operation for $r\downarrow 0$ is not needed to derive the stationary distribution of $Z(t)$.
-Thus, we need not further discuss about the conjectures (a) and (c).
+[1, Lemma 4.3] (see also [12] for a different proof), while the author affirmatively answers (c) in [12].
+Those results are obtained for more general reflecting diffusion processes than $Z(\cdot)$, which are hard to include in this paper.
+Thus, we refer to [12] for them, and will not further discuss about the conjectures (a) and (c).
 As for the conjecture (b), it is still open (see [12]).
@@
 A.2 Proof of Proposition 3.1
-Because of the exponentially distributed assumption, for each fixed $r$, $L^{(r)}(\cdot)$ is a continuous-time Markov chain with transition rates $q^{(r)}_{i,j}$ for $i,j\in\mathbb{Z}_{+}$ given by
-$q^{(r)}_{0,0}=-c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{0,1}=\lambda^{(r)}_{1},$
-$q^{(r)}_{i,i}=-(c^{(r)}_{1}\mu^{(r)}+\lambda^{(r)}_{1}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{1},\quad 1\leq i\leq\ell^{(r)}_{0}-1,$
-$q^{(r)}_{i,i}=-(c^{(r)}_{1}\mu^{(r)}+\lambda^{(r)}_{2}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{2},\quad i=\ell^{(r)}_{0},$
-$q^{(r)}_{i,i}=-(c^{(r)}_{2}\mu^{(r)}+\lambda^{(r)}_{2}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{2}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{2},\quad i\geq\ell^{(r)}_{0}+1,$
-where all the other $q_{i,j}$’s vanish.
-Thus, $L^{(r)}(\cdot)$ is a birth-and-death process.
-By the stability condition (3.e), it has the stationary distribution.
-Denote it by $\pi^{(r)}_{n}=\mathbb{P}(L^{(r)}=n)$ for $n\geq 0$, then it is easy to see that
-$\pi^{(r)}_{n}=\left\{\begin{array}[]{ll}\pi^{(r)}_{0}(\rho^{(r)}_{1})^{n},&0\leq n\leq\ell^{(r)}_{0},\\ \pi^{(r)}_{0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}(\rho^{(r)}_{2})^{n-\ell^{(r)}_{0}},&n>\ell^{(r)}_{0},\end{array}\right.$ (A.3)
-where
-$\pi^{(r)}_{0}=\left\{\begin{array}[]{ll}\displaystyle\frac{(1-\rho^{(r)}_{1})(1-\rho^{(r)}_{2})}{1-\rho^{(r)}_{2}+(\rho^{(r)}_{2}-\rho^{(r)}_{1})(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}},&\rho^{(r)}_{1}\not=1,\\ \displaystyle\frac{1-\rho^{(r)}_{2}}{(\ell^{(r)}_{0}+1)(1-\rho^{(r)}_{2})+\rho^{(r)}_{2}},&\rho^{(r)}_{1}=1.\end{array}\right.$ (A.6)
-Hence, we can directly compute $\widetilde{\nu}_{i}(\mathbb{R}_{+})$ for $i=1,2$ through $\pi^{(r)}_{0}$ of (A.6), which yields
-$\nu^{(r)}_{1}(\mathbb{R}_{+})=\sum_{n=0}^{\ell^{(r)}_{0}}\pi^{(r)}_{n}=\left\{\begin{array}[]{ll}\displaystyle\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}+1}}{1-\rho^{(r)}_{1}}\pi^{(r)}_{0},&\rho^{(r)}_{1}\not=1,\\ \displaystyle(\ell^{(r)}_{0}+1)\pi^{(r)}_{0},&\rho^{(r)}_{1}=1,\end{array}\right.$ (A.9)
-$\nu^{(r)}_{2}(\mathbb{R}_{+})=\sum_{n=\ell^{(r)}_{0}+1}^{\infty}\pi^{(r)}_{n}=\frac{\rho^{(r)}_{2}}{1-\rho^{(r)}_{2}}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\pi^{(r)}_{0}.$ (A.10)
-We now consider the asymptotic behaviors of $\pi^{(r)}_{0}$ and $\widetilde{\nu}_{i}(\mathbb{R}_{+})$ for $i=1,2$. Form the definition of $\rho^{(r)}_{i}$ and the assumptions (3.b) and (3.c),
-$\rho^{(r)}_{i}=\frac{\lambda^{(r)}_{i}}{c^{(r)}_{i}\mu^{(r)}}=1-\frac{r\mu b_{i}+o(r)}{c^{(r)}_{i}\mu^{(r)}}=1-\frac{b_{i}}{c_{i}}r+o(r),\qquad r\downarrow 0.$
-We first consider the case that $b_{1}\not=0$. Since $\beta_{i}=b_{i}/c_{i}$ by the exponential assumption,
-$(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\big{(}1-\beta_{1}r+o(r)\big{)}^{r^{-1}\ell_{0}+o(1)}=e^{-\beta_{1}\ell_{0}}+o(1)),$ (A.11)
-and therefore (A.6) yields
-$\pi^{(r)}_{0}$ $=\frac{\beta_{1}\beta_{2}r^{2}+o(r^{2})}{\beta_{2}r+o(r)+(e^{-\beta_{1}\ell_{0}}+o(1))((\beta_{1}-\beta_{2})r+o(r))}$
-$=\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))=\frac{b_{1}}{c_{1}}\frac{c_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}r+o(r),$ (A.12)
-which is compatible with (4.59). It follows from (A.9) and (A.2) that
-$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})$ $=\lim_{r\downarrow 0}\frac{1-e^{-\beta_{1}\ell_{0}}+o(1)}{\beta_{1}r+o(r)}\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))$
-$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)+c_{2}b_{1}},$ (A.13)
-where the last equality uses the fact that $\beta_{i}=b_{i}/c_{i}$ for $i=1,2$.
-We next consider the case that $b_{1}=0$. In this case, $\rho^{(r)}_{1}=1-o(r)$, and therefore, from (A.6),
-$\pi^{(r)}_{0}$ $=\frac{\beta_{2}}{1+\beta_{2}\ell_{0}}(r+o(r))=\frac{b_{2}}{c_{2}+b_{2}\ell_{0}}r+o(r),$ (A.14)
-which is compatible with (4.62). Since $\beta_{2}=b_{2}/c_{2}$, (A.9) and (A.14) yield
-$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\ell^{(r)}_{0}\pi^{(r)}_{0}=\frac{b_{2}\ell_{0}}{c_{2}+b_{2}\ell_{0}}.$ (A.15)
-Hence, from (3.14), (A.2) and (A.15) and using the fact that $\nu^{(r)}_{1}(\mathbb{R}_{+})+\nu^{(r)}_{2}(\mathbb{R}_{+})=1$, we have
-$\lim_{r\downarrow 0}\nu^{(r)}_{i}(\mathbb{R}_{+})$ $=\widetilde{A}_{\exp,i},\qquad i=1,2.$ (A.16)
-It remains to verify (3.4) and (3.5).
-This can be done through the Laplace transforms of $\{\pi^{(r)}_{n};0\leq n\leq\ell^{(r)}_{0}\}$ and $\{\pi^{(r)}_{n};n>\ell^{(r)}_{0}\}$ together with (A.16).
-We omit its detail because they are just computations.
+We prove Proposition 3.1 only for $b_{1}\not=0$ because the case for $b_{1}=0$ can be similarly proved.
+We first consider this 2-level $M/M/1$ queue without index $r$ for notational convenience.
+Because of the exponential assumptions, we can describe this model by the continuous time Markov chain with state space
+$S=\{(m,i);m=1,2,i=0,1,\ldots,\ell_{0}\}\cup\{(3,i);i\geq\ell_{0}+1\},$
+and its transition diagram is depicted in Figure A.1.
+Figure A.1: Transition diagram
+Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$.
+Note that arrivals may find the queue length less than $\ell_{0}$ when their last arrivals found the queue length not less than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
+Because of this, the arrival rate at any time may depend on the queue length at the last arrival instant, and therefore $L(\cdot)$ cannot be a continuous-time Markov chain.
+We partition the state space $S$ of $X_{\exp}(\cdot)$ to the following three disjoint subsets.
+$S_{m}=\{(m,i)\in S;i\in[0,\ell_{0}]\},\quad m=1,2,\qquad S_{3}=\{(3,i):i\geq\ell_{0}+1\}.$
+By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{m,i}$ as
+$p_{m,i}=\mathbb{P}[J=m,L=i],\qquad m=1,2,3,(m,i)\in S_{m}.$
+Obviously, the stationary distribution, $\{p_{m,i};(m.i)\in S\}$, is uniquely determined by the following stationary equations. For state in $S_{1}$,
+$\lambda_{1}p_{1,0}=c_{1}\mu p_{1,1},$ (A.1)
+$(\lambda_{1}+c_{1}\mu)p_{1,i}=\lambda_{1}p_{1,i-1}+c_{1}\mu p_{1,i+1}+\lambda_{2}p_{2,i-1},\quad i\in[1,\ell_{0}-1],$ (A.2)
+$(\lambda_{1}+c_{1}\mu)p_{1,\ell_{0}}=\lambda_{1}p_{1,\ell_{0}-1}+\lambda_{2}p_{2,\ell_{0}-1},$ (A.3)
+and, for states in $S_{2}$,
+$\lambda_{2}p_{2,0}=c_{1}\mu p_{2,1},$ (A.4)
+$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,i+1},\qquad i\in[1,\ell_{0}-1],$ (A.5)
+$(\lambda_{2}+c_{1}\mu)p_{2,\ell_{0}}=c_{2}\mu p_{3,\ell_{0}+1},$ (A.6)
+For state in $S_{3}$, it follows from the equation for $\ell_{0}+1$ and the local balance that
+$(\lambda_{2}+c_{2}\mu)p_{3,\ell_{0}+1}=\lambda_{1}p_{1,\ell_{0}}+\lambda_{2}p_{2,\ell_{0}}+c_{2}\mu p_{\ell_{0}+2},$ (A.7)
+$\lambda_{2}p_{3,i}=c_{2}\mu p_{3,i+1},\qquad i\geq\ell_{0}+1.$ (A.8)
+In what follows, we use the following notations.
+$\rho_{ij}=\frac{\lambda_{i}}{c_{j}\mu},\qquad\gamma_{ij}=\frac{c_{i}\mu}{c_{1}\mu+\lambda_{j}},\qquad i,j=1,2.$
+We first note that, from (A.8),
+$p_{3,i}=\rho_{22}^{i-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad i\geq\ell_{0}+1.$ (A.9)
+Thus, $p_{3,i}$ for $i\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
+We next consider $p_{2,i}$ for $i\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
+$p_{2,0}=\frac{c_{1}\mu}{\lambda_{2}}p_{2,1}=\rho_{21}^{-1}p_{2,1},$ (A.10)
+$p_{2,i}=\gamma_{12}p_{2,i+1}=\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}},\quad i\in[1,\ell_{0}-1].$ (A.11)
+Thus, $p_{2,i}$ for $i\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written as
+$p_{2,i}=\gamma_{12}^{-1}p_{2,i-1}=\gamma_{12}^{-i+1}p_{2,1}=\gamma_{12}^{-i+1}\rho_{21}p_{2,0},\quad i\in[1,\ell_{0}-1].$ (A.12)
+Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{m}$ for $m=1,2,3$.
+We call them boundary probabilities.
+There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{0}}$ for large $\ell_{0}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}-1}=\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$ by (A.12) and (A.11), so we do not consider it.
+We first note that, from (A.6),
+$p_{2,\ell_{0}}=\gamma_{22}p_{3,\ell_{0}+1}.$ (A.13)
+Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $i=\ell_{0}+1$, (A.7) can be written as
+$p_{3,\ell_{0}+1}=\rho_{12}p_{1,\ell_{0}}+\rho_{22}p_{2,\ell_{0}},$ (A.14)
+and, substituting $p_{3,\ell_{0}+1}=\gamma_{22}^{-1}p_{2,\ell_{0}}$ of (A.13) to this equation, we have
+$(\gamma_{22}^{-1}-\rho_{22})p_{2,\ell_{0}}=\rho_{12}p_{1,\ell_{0}}.$
+Since $\gamma_{22}^{-1}-\rho_{22}=\frac{c_{1}\mu+\lambda_{2}}{c_{2}\mu}-\frac{\lambda_{2}}{c_{2}\mu}=\frac{c_{1}}{c_{2}}=\rho_{12}\rho_{11}^{-1}$, this equation yields
+$p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}.$ (A.15)
+Thus, the boundary probabilities $p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ are explicitly related by (A.13) and (A.15).
+We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{0}}$, but this is not easy.
+To resolve this problem, we consider $p_{1,i}$ for $i\in[0,\ell_{0}-1]$. From (A.2),
+$\lambda_{1}p_{1,i}-c_{1}\mu p_{1,i+1}=\lambda_{1}p_{1,i-1}-c_{1}\mu p_{1,i}+\lambda_{2}p_{2,i-1},\qquad i\in[1,\ell_{0}-1],$
+then this equation can be written as
+$\rho_{11}p_{1,i}-p_{1,i+1}=\rho_{11}p_{1,i-1}-p_{1,i}+\rho_{21}p_{2,i-1},\qquad i\in[1,\ell_{0}-1].$
+Summing up both sides of this equation for $i\in[1,j]$ for $j\in[1,\ell_{0}-1]$, we have
+$\rho_{11}p_{1,j}-p_{1,j+1}=\rho_{21}\sum_{i=1}^{j}p_{2,i-1},\qquad j\in[0,\ell_{0}-1],$ (A.16)
+because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $j=0$ by the convention that the empty sum vanishes, which we always take.
+Hence, applying (A.10) and (A.11) to (A.16), we have
+$\rho_{11}p_{1,j}-p_{1,j+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{i=1}^{j-1}\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}}\right)$
+$=\left(1+\rho_{21}\sum_{i=1}^{j-1}\gamma_{12}^{-(i-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
+$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[1,\ell_{0}-1].$
+Multiplying both sides by $\rho_{11}^{-(j+1)}$,
+$\rho_{11}^{-j}p_{1,j}-\rho_{11}^{-(j+1)}p_{1,j+1}$
+$\quad=\rho_{11}^{-(j+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(j\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[0,\ell_{0}-1].$
+Summing up both sides of this equation from $j=0$ to $j=k-1\leq\ell_{0}-1$, we have
+$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{j=0}^{k-1}\rho_{11}^{-(j+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{j=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(j-1)}-\rho_{11}^{-(j-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
+$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}}$
+$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}\rho_{11}^{-2}-\frac{\rho_{11}^{-k}-\rho_{11}^{-1}}{1-\rho_{11}}\right)\frac{\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}},\quad k\in[1,\ell_{0}].$
+Hence, multiplying both sides of this equation by $\rho_{11}^{k}$, we have
+$p_{1,k}$ $=\rho_{11}^{k}p_{1,0}-\frac{1-\rho_{11}^{k}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\frac{1-\rho_{11}^{k-1}}{1-\rho_{11}}\frac{\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
+$\quad-\frac{\gamma_{12}^{-(k-1)}-\rho_{11}^{k-1}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}},\qquad k\in[1,\ell_{0}].$ (A.17)
+In particular, for $k=\ell_{0}$, substituting $p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}$ of (A.15), we have
+$\rho_{11}^{\ell_{0}}\frac{p_{1,0}}{p_{1,\ell_{0}}}$ $=1+\frac{1-\rho_{11}^{\ell_{0}}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}\rho_{11}-\frac{1-\rho_{11}^{\ell_{0}-1}}{1-\rho_{11}}\frac{\rho_{11}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}$
+$\quad+\frac{\rho_{11}\gamma_{12}-\rho_{11}^{\ell_{0}}\gamma_{12}^{\ell_{0}}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}.$ (A.18)
+Thus, $p_{1,0}$ is explicitly related to $p_{1,\ell_{0}}$.
+We now apply these computations to the $r$-th 2-level $M/M/1$ queue, and compute the limit of its scaled stationary distribution through moment generating functions.
+For the $r$-th system, we put the superscript (r) as we have used for the $r$-th 2-level $GI/G/1$ queue.
+For example, $p^{(r)}_{m,i}$ is the stationary probability of $\{(J,L)=(m,i)\}$ for $(m,i)\in S$.
+Define the moment generating functions $g^{(r)}_{m}(\theta)$ for $m=1,2,3$ as
+$g^{(r)}_{m}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{m,k},\quad\theta\in\mathbb{R},m=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
+and define $g^{(r)}(\theta)=\sum_{m=1}^{3}g^{(r)}_{m}(\theta)$.
+Obviously, $g^{(r)}(\theta)$ is the moment generating function of $rL^{(r)}$.
+Thus, our task is to compute the limit of $g^{(r)}(\theta)$ as $r\downarrow 0$.
+To this end, we compute $g^{(r)}_{m}(\theta)$ for $m=1,2,3$, using the normalization condition that $g^{(r)}(0)=1$.
+However, it should be noted that this condition is not sufficient to prove the weak convergence of the distribution of $rL^{(r)}$, for which it is required that $\lim_{\theta\uparrow 0}\lim_{r\downarrow 0}g^{(r)}(\theta)=1$.
+This weak convergence will be verified through explicitly computing $g^{(r)}_{m}(\theta)$ for $m=1,2,3$.
+We first compute $g^{(r)}_{m}(\theta)$ for $m=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
+$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{i=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-i)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-i}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{0}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}},$ (A.19)
+$g^{(r)}_{3}(\theta)$ $=\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}.$ (A.20)
+On the other hand, computing $g^{(r)}_{1}(\theta)$ is a bit complicated. To ease it, we introduce the following functions.
+$f^{(r)}_{1}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}e^{r\theta k},\qquad f^{(r)}_{\rho_{1}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}\rho^{(r)}_{1})^{k},\qquad f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k},$
+then
+$f^{(r)}_{1}(\theta)=\frac{1-(e^{r\theta})^{\ell^{(r)}_{0}}}{1-e^{r\theta}}1(\theta\not=0)+\ell^{(r)}_{0}1(\theta=0),\qquad f^{(r)}_{\rho_{1}}(\theta)=\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-e^{r\theta}\rho^{(r)}_{1}},$
+$f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\frac{e^{r\theta\ell^{(r)}_{0}}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{-\ell^{(r)}_{0}+1}.$
+Using these function, we compute $g^{(r)}_{1}(\theta)$ from (A.2) as given below.
+$g^{(r)}_{1}(\theta)$ $=\sum_{k=0}^{\ell^{(r)}_{0}}(e^{r\theta}\rho^{(r)}_{1})^{k}p^{(r)}_{1,0}-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta k}-(e^{r\theta}\rho^{(r)}_{1})^{k}}{1-\rho^{(r)}_{1}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad+\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta(k-1)}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{1-\rho^{(r)}_{1}}\frac{e^{r\theta}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k-1}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell_{0}}$
+$=\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(\theta)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad+\left(f^{(r)}_{1}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$ (A.21)
+We now compute the limits of $g^{(r)}_{m}(0)$ for $m=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
+$(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=O(\gamma_{12}^{\ell_{0}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.22)
+Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}$ decays exponentially fast as the function of $1/r$. On the other hand,
+$\lim_{r\downarrow 0}(1-\rho^{(r)}_{i})/r=\beta_{i},\qquad i=1,2,$ (A.23)
+$\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{0}/r+o(r)}=e^{-\beta_{1}\ell_{0}},$ (A.24)
+where recall that $\beta_{i}=b_{i}/c_{i}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.
+Lemma A.1.
+For the sequence of the 2-level $M/M/1$ queue satisfying (3.b)–(3.e),
+$\lim_{r\downarrow 0}\frac{p^{(r)}_{2,\ell_{0}}}{p^{(r)}_{1,\ell_{0}}}=1,\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{3,\ell_{0}+1}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}},\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}.$ (A.25)
+Proof.
+The first two limits in (A.25) aer immediate from (A.15) and (A.13) because
+$\lim_{r\downarrow 0}\gamma^{(r)}_{22}=\lim_{r\downarrow 0}\frac{c^{(r)}_{2}\mu^{(r)}}{c^{(r)}_{1}\mu+\lambda^{(r)}_{2}}=\frac{c_{2}\mu}{c_{1}\mu+\lambda_{2}}=\frac{\lambda_{2}}{\lambda_{1}+\lambda_{2}}.$
+For the third limit, we have, from (A.2) for the $r$-th system,
+$\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}$ $=\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{-\ell^{(r)}_{0}}\left(1+\frac{\rho^{(r)}_{1}\gamma^{(r)}_{12}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}\right).$
+Hence, applying (A.24), we have the third limit of (A.25). ∎
+We now tentatively assume that there exists finite $\widetilde{p}_{1,\ell_{0}}$ such that
+$\widetilde{p}_{1,\ell_{0}}=\lim_{r\downarrow 0}\frac{p^{(r)}_{1,\ell^{(r)}_{0}}}{r}.$ (A.26)
+The existence of this $\widetilde{p}_{1,\ell_{0}}$ will be verified at the final step (see (A.31)). Applying these asymptotic properties, it follows from (A.2) and (A.20) with help of (A.26) that
+$\lim_{r\downarrow 0}g_{2}^{(r)}(0)$ $=\lim_{r\downarrow 0}\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}}$
+$=\frac{c_{1}\mu+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{1,\ell^{(r)}_{0}}=0,$ (A.27)
+$\lim_{r\downarrow 0}g^{(r)}_{3}(0)$ $=\lim_{r\downarrow 0}\frac{1}{1-\rho^{(r)}_{2}}p^{(r)}_{3,\ell^{(r)}_{0}+1}=\frac{c_{2}}{b_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
+$=\beta_{2}^{-1}\gamma_{22}^{-1}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{2}^{-1}\widetilde{p}_{1,\ell_{0}}.$ (A.28)
+To compute the limit of $g^{(r)}_{1}(0)$, we first compute it from (A.2) as
+$g^{(r)}_{1}(0)$ $=\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad+\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}.$ (A.29)
+To proceed to further compute, we note that
+$\lim_{r\downarrow 0}rf^{(r)}_{1}(0)=\lim_{r\downarrow 0}r\ell^{(r)}_{0}=\ell_{0},$
+$\lim_{r\downarrow 0}rf^{(r)}_{\rho_{1}}(0)=\lim_{r\downarrow 0}r\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-\rho^{(r)}_{1}}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}}),$
+$\lim_{r\downarrow 0}f^{(r)}_{\gamma_{12}^{-1}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\gamma^{(r)}_{12}=\frac{c_{1}\mu}{\lambda_{2}}=\rho_{21}^{-1}.$
+Then, we have, using Lemma A.1,
+$\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}})\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,0}$
+$\quad=\beta_{1}^{-1}\left(e^{\beta_{1}\ell_{0}}-1\right)\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\widetilde{p}_{1,\ell_{0}},$
+$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}=0,$
+$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}=0,$
+and
+$\lim_{r\downarrow 0}\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
+$\quad=\lim_{r\downarrow 0}f^{(r)}_{\gamma^{-1}_{12}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}p^{(r)}_{2,\ell^{(r)}_{0}}=0.$
+Hence, from (A.2)
+$\lim_{r\downarrow 0}g^{(r)}_{1}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{1}^{-1}(e^{\beta_{1}\ell_{0}}-1)\widetilde{p}_{1,\ell_{0}}.$ (A.30)
+Hence, by (A.2) and (A.2), we have
+$1=\lim_{r\downarrow 0}g^{(r)}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\left(\beta_{2}^{-1}+\beta_{1}^{-1}(e^{\beta_{1}\ell_{0}}-1)\right)\widetilde{p}_{1,\ell_{0}}.$ (A.31)
+This proves that $\widetilde{p}_{1,\ell_{0}}$ indeed exists and is finite, and given by
+$\widetilde{p}_{1,\ell_{0}}=\frac{\lambda_{2}\beta_{1}\beta_{2}}{(\lambda_{1}+\lambda_{2})(\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1))}.$ (A.32)
+We are now in a final step to find $\widetilde{g}(\theta)\equiv\lim_{r\downarrow 0}g^{(r)}(\theta)$ for $\theta<0$.
+This can be obtained from (A.2), (A.2) and (A.20) in the following way.
+From (A.20) and Lemma A.1,
+$\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$ $=\lim_{r\downarrow 0}\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
+$=\lim_{r\downarrow 0}\frac{r}{1-(1+r\theta+o(r\theta))(1-r\beta_{2}+o(r))}e^{\ell^{(r)}_{0}\theta}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
+$=\frac{e^{\ell_{0}\theta}}{\beta_{2}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,\ell^{(r)}_{0}}$
+$=\frac{e^{\ell_{0}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},$
+while, from (A.2) and Lemma A.1,
+$\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)$ $=\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\lim_{r\downarrow 0}\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}+1}}{1-e^{r\theta}\rho^{(r)}_{1}}p^{(r)}_{1,0}$
+$=\frac{1-e^{\ell_{0}(\theta-\beta_{1})}}{\beta_{1}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}\widetilde{p}_{1,\ell_{0}}$
+$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\ell_{0}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}$
+Thus, we have, for $\theta\leq 0$ satisfying $\theta\not=\beta_{1}$,
+$\lim_{r\downarrow 0}g^{(r)}(\theta)$ $=\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)+\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$
+$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\ell_{0}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}+\frac{e^{\ell^{(r)}_{0}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}.$
+This completes the proof of Proposition 3.1.
 A.3 Proofs of Lemma 4.4
@@
 where (4.1) is still applicable since the right-hand side of this equation is finite. Hence, (4.22) is obtained. (4.23) is similarly proved. We next prove (4.24).
-$|(x\wedge 1/r)^{k}-x^{k}|$ $\leq x^{k}1(x>1/r)<rx^{k+1}1(x>1/r)$
+$|(x\wedge 1/r)^{k}-x^{k}|$ $\;=\;x^{k}1(x>1/r)<rx^{k+1}1(x>1/r)$
 $<r^{2}x^{k+2}1(x>1/r),\qquad x\geq 0,k\geq 1.$ (A.33)

### v6

@@
 This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed.
-To get its performance in tractable form, we study the limit of the stationary distribution of this two-level queue under the scaling in heavy traffic which corresponds to a diffusion approximation.
-It is shown that this limiting distribution is truncated exponential below the threshold level and exponential above it under suitably chosen system parameters.
+To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic.
+This limit corresponds to its diffusion approximation.
+It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters.
 This result is proved using the so called BAR approach studied in [2, 3, 10, 12].
-We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distributions.
-Keywords: single server queue, level dependent arrival and service, stationary distribution, heavy traffic, diffusion approximation, Palm distribution, process limit, stochastic integral equation.
-11footnotetext: Department of Information Sciences, Tokyo University of Science, Noda, Chiba, Japan.
+We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
+Keywords: single server queue; level dependent arrival and service; stationary distribution; heavy traffic; diffusion approximation; Palm distribution; stochastic integral equation; multi-level SRBM.
 1 Introduction
@@
 So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length.
-It may be considered to approximate it by a quasi-birth-and-death process, then it is not hard to numerically compute its performance.
-However, it is hard to derive it in an analytically tractable form except for special cases such as all the distributions to be exponential.
-Because of this, a simple approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution.
+However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured.
+Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution.
 This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold.
@@
 Thus, theoretical study for its diffusion approximation is challenging.
-As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue (e.g., see Kingman [7]).
+As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see Kingman [7]).
 Furthermore, Miyazawa [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room.
@@
 In Section 5.2, we intuitively verify the conjecture in Remark 3.1.
-2 Framework for the 2-level GI/G/1 queues
+2 Framework for the 2-level GI/G/1 queue
 We take the stochastic framework introduced in [12] to study the 2-level $GI/G/1$ queue.
@@
 To this end, we introduce a sequence of this queueing models indexed by $r\in(0,1]$.
-Since we consider their countable sequence, one may consider $r=1/m$ for positive integer $m$, but we keep $r$ just for convenience.
+Since this sequence can be countable, one may take $r=1/\sqrt{n}$ for positive integer $n$ so that the scaling is identical with the standard diffusion approximation, but we keep $r$ just for convenience.
 3.1 Assumptions
-For the $r$-th model, let $\ell^{(r)}_{0}$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$.
+For the $r$-th model, let $\ell^{(r)}_{0}\geq 2$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$.
 We refer to this number as the queue length.
@@
 Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length.
-Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are random variables which may have different distributions.
+Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are independent random variables which may have different distributions.
 As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers.
@@
   • (3.d)
-    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$, and $\ell^{(r)}_{0}\geq 2$.
+    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$.
   • (3.e)
     $\rho^{(r)}_{2}<1$.
-Here, we write $f(r)=o(g(r))$ as $r\downarrow 0$ for positive valued functions $f,g$ of $r$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$.
+Here, for positive valued functions $f,g$ of $r$, we write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$.
 Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$.
@@
 For convenience, let $R^{(r)}=(R^{(r)}_{e},R^{(r)}_{s})$.
-For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting distribution agrees with that of $T^{(r)}_{e,i}$.
+For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting times have the same distribution as that of $T^{(r)}_{e,i}$.
 Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$, and let $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper.
@@
 Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework.
-Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}(n)\wedge T_{e,2}(n)]<\infty$.
-Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)]<\infty$.
+Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$.
+Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}]<\infty$.
 ∎
@@
 We refer to this $\nu^{(r)}$ as a scaled stationary distribution.
-We decompose $\nu^{(r)}$ into measures $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
+We decompose $\nu^{(r)}$ using finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$, which are defined as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
 $\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}>\ell^{(r)}_{0}),$
@@
 We conjecture that (3.11) holds without any condition in (ii).
-There are two supports for this conjecture.
-One is from a process limit for the 2-level $GI/G/1$ queue in heavy traffic.
-Another is from the intuitive verification of (3.12).
-They are detailed in Sections 5.1 and 5.2, repectively.
+This is supported by the intuitive verification of (3.12), which is detailed in Section 5.2.
 Let $S^{(r)}(n)$ be the sojourn time of the $n$-th arriving customer in service, then the traffic intensity $\rho^{(r)}$ can be defined under the stationary setting as
@@
 =$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.14)
-Let consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
+Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
 Proposition 3.1.
@@
 Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed.
-However, for a double check, we give its direct proof in Appendix A.2, which is helpful to intuitively convince (ii) of Theorem 3.1.
-This proof is straightforward, but one must be careful because $L^{(r)}(\cdot)$ itself is not a continuous-time Markov chain even under the exponential assumptions (see Appendix A.2 for its reason).
+However, for a double check, we give its direct proof in Appendix A.2, which is helpful to convince (ii) of Theorem 3.1.
+This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of more than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
+This Markov chain will be detailed in Appendix A.2.
 4 Proofs
@@
 $\mathbb{P}(\Theta_{t}^{-1}A)=\mathbb{P}(A),\qquad A\in{\mathcal{F}}.$
-[…]
+Then, for each $r\in(0,1]$, $X^{(r)}(\cdot)$ and the increments of $N^{(r)}_{e}$ and $N^{(r)}_{s}$ are jointly stationary.
+Since $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite by Lemma 3.1, we can define probability distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ on $(\Omega,{\mathcal{F}})$ by
 $\mathbb{P}^{(r)}_{v}(A)=(\alpha^{(r)}_{v})^{-1}\mathbb{E}\left[\int_{(0,1]}1_{\Theta_{u}^{-1}A}N^{(r)}_{v}(du)\right],\qquad A\in{\mathcal{F}},v=e,s,$ (4.1)
@@
 For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$.
-Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded functions from $S$ to $\mathbb{R}_{+}$ which continuously and partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right.
+Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded and continuous functions from $S$ to $\mathbb{R}_{+}$ which are partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right.
 Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
@@
 In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit.
-Define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
+For $t\geq 0$, define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
 $X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
-$L^{(r)}_{1}(t)=L(t-)+1(\Delta N^{(r)}_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
+$L^{(r)}_{1}(t)=L^{(r)}(t-)+1(\Delta N^{(r)}_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
 where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$.
@@
 Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$ when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
+In these definitions, $X^{(r)}(0-)$ is required, but formally not defined.
+However, this does not cause any problem.
+Please see (i) of Remark 4.1 for its details.
 For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
@@
 where the summation is well defined because $\Delta_{e}f(X^{(r)}(u))$ and $\Delta_{s}f(X^{(r)}(u))$ vanish except for finitely many $u\in(0,t]$.
-Taking the expectation of the above equation by $\mathbb{P}$, we have a stationary equation:
+Take the expectation of the above equation for $t=1$ by $\mathbb{P}$, then the stationarity of $X(\cdot)$ and the definition of the Palm distribution yield
+[…]
+$\mathbb{E}\left[\sum_{u\in(0,1]}\Delta_{v}f(X^{(r)}(u))]\right]=\alpha^{(r)}_{v}\mathbb{E}^{(r)}_{v}\left[\Delta_{v}f(X^{(r)}(0))\right],\qquad v=e,s.$
+Hence, we have a stationary equation:
 $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$
@@
   • (i)
-    $X^{(r)}(0-)$ is not defined for $X^{(r)}(\cdot)=\{X^{(r)}(t);t\geq 0\}$, but well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$.
-Thus, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined.
+    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distribution $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$.
+So, $X^{(r)}(0-)$ can be considered as a random variable generated by $\mathbb{P}^{(r)}_{e}$.
+However, one may be uncomfortable for this interpretation.
+In this case, please consider the process $X(t)$ to start at time $-1$ subject to the stationary distribution, then the distribution of $X(\cdot)=\{X(t);t\geq 0\}$ is unchanged because $X(\cdot)$ is a stationary process.
+In either way, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions.
   • (ii)
@@
 We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$.
-This prevents to explicitly refer to the intermediate state $X^{(r)}_{1}(0)$.
+This prevents to directly refer to the intermediate state $X^{(r)}_{1}(0)$ in our arguments.
 Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
 $\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
-Apply this to (4.1) yields
+Applying this to (4.1) yields
 $\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
@@
 $w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{0}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{0}),\qquad z\in\mathbb{Z}_{+},$
-$g^{(r)}_{i,\theta}(x_{1},x_{2})=e^{-\eta^{(r)}_{i}}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
-where $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are the unique solutions of the following equations for each $\theta,\theta\in\mathbb{R}$.
+$g^{(r)}_{i,\theta}(x_{1},x_{2})=\exp\left(-\eta^{(r)}_{i}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r)\right),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
+where $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are the unique solutions of the following equations for each $\theta\in\mathbb{R}$.
 $e^{\theta}\mathbb{E}\left[e^{-\eta^{(r)}_{i}(\theta)(T^{(r)}_{e,i}\wedge 1/r)}\right]=1,\quad i=1,2,\quad e^{-\theta}\mathbb{E}\left[e^{-\zeta^{(r)}(\theta)(T^{(r)}_{s}\wedge 1/r)}\right]=1,$ (4.8)
 because $\eta^{(r)}_{i}(\theta)$ is the inverse function of the Laplace transform of the finite positive random variable $T^{(r)}_{e,i}\wedge 1/r$ taking the value $e^{-\theta}$ and $\zeta^{(r)}(\theta)$ is similarly determined.
-Furthermore, $\eta^{(r)}_{i}(\theta),\zeta^{(r)}(\theta)$ are infinitely many differentiable functions of $\theta$.
+Then, for each ${\boldsymbol{\theta}}\equiv(\theta_{1},\theta_{2})\in\mathbb{R}^{2}$, define test function $f^{(r)}_{{\boldsymbol{\theta}}}$ for the BAR as
+$f^{(r)}_{{\boldsymbol{\theta}}}(z,x_{1},x_{2})=\sum_{i=1,2}e^{\theta_{i}z}g^{(r)}_{i,\theta_{i}}(x_{1},x_{2})w^{(r)}_{i}(z),\qquad(z,x_{1},x_{2})\in S.$ (4.9)
+Functions $\eta^{(r)}_{i}(\theta)$ and $\zeta^{(r)}(\theta)$ play key roles in our analysis.
+Consider their properties.
+First, it is easy to see that they are infinitely many differentiable functions of $\theta$.
 Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin.
@@
 $\zeta^{(r)}(r\theta)=-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}r^{2}\theta^{2}+o(r^{2}).$ (4.11)
-Furthermore, by , for $i=1,2$, there are constant $d_{e,i},d_{s},a>0$ such that, for $r\in(0,1]$ and $\theta\in\mathbb{R}$ satisfying $r|\theta|<a$,
+Furthermore, for $i=1,2$, there are constant $d_{e,i},d_{s},a>0$ such that, for $r\in(0,1]$ and $\theta\in\mathbb{R}$ satisfying $r|\theta|<a$,
 $\left|\eta^{(r)}_{i}(r\theta)(u_{1}\wedge 1/r)+\zeta^{(r)}(r\theta)(u_{2}\wedge 1/r)\right|\leq|\theta|\left(d_{e,i}(ru_{1}\wedge 1)+d_{s}(ru_{2}\wedge 1)\right).$ (4.12)
@@
 $\lim_{r\downarrow 0}\mathbb{E}\left[g^{(r)}_{i,r\theta_{i}}(R^{(r)})\right]=1,\qquad\lim_{r\downarrow 0}\mathbb{E}_{u}\left[g^{(r)}_{i,r\theta_{i}}(R^{(r)})\right]=1,\quad u=e,s.$ (4.13)
-We define test function $f^{(r)}_{{\boldsymbol{\theta}}}$ for the BAR as
-$f^{(r)}_{{\boldsymbol{\theta}}}(z,x_{1},x_{2})=\sum_{i=1,2}e^{\theta_{i}z}g^{(r)}_{i,\theta_{i}}(x_{1},x_{2})w^{(r)}_{i}(z),$ (4.13)
-and auxiliary functions $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ as
+We next introduce auxiliary functions $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ as
 $\psi^{(r)}_{i}(\theta_{i})=\mathbb{E}\left[e^{r\theta_{i}L^{(r)}}g^{(r)}_{i,r\theta_{i}}(R^{(r)})w^{(r)}_{i}(L^{(r)})\right],\qquad\theta_{1}\in\mathbb{R},\theta_{2}\leq 0.$ (4.14)
@@
 because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
-Based on these results, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
-$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\sum_{i=1,2}\eta_{i}^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})$
+Based on these facts, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
+$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\sum_{i=1,2}\eta_{i}^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})$
 $\qquad+\sum_{i=1,2}c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})-c^{(r)}_{1}\zeta^{(r)}(r\theta_{1})\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)}))\right].$ (4.16)
@@
 Hence, from (4.2), we have
-$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\left(-\beta_{i}+\theta_{i}\right)\mu\theta_{i}r^{2}\psi^{(r)}_{i}(\theta_{i})$
-$\hskip 64.58313pt+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}),$ (4.17)
-We next compute the Palm expectation terms in (4.1). In what follows, we use the following simplified notations.
+$\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\left(-\beta_{i}+\theta_{i}\right)\mu\theta_{i}r^{2}\psi^{(r)}_{i}(\theta_{i})$
+$\qquad+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}).$ (4.17)
+We next compute the Palm expectation terms in (4.1). In what follows, we use the following abbreviated notations.
 $\widehat{T}^{(r)}_{e,i}\equiv T^{(r)}_{e,i}(0)\wedge 1/r,\quad i=1,2,\qquad\widehat{T}^{(r)}_{s}\equiv T^{(r)}_{s}(0-)\wedge 1/r,$
 $\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$
-Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ and $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
-$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
-$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{{\boldsymbol{\theta}},r}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
+Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
+$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
+$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
 Hence, from the definitions of $X^{(r)}_{1}(0)$ and $\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))$, we have
@@
 $(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}\in B)],$ (4.22)
-$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
+$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]$
+$\qquad\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
 and, for $i=1,2$,
@@
 Hence, dividing both sides of (4.3) by $r$ and letting $r\downarrow 0$, we have
-$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
-Since $e^{r\theta\ell_{0}}$ can be arbitrarily small, this implies that
+$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
+Since $e^{r\theta\ell^{(r)}_{0}}=e^{(\ell_{0}+o(1))\theta}$ can be arbitrarily small by choosing small $\theta<0$, we have
 $\lim_{r\downarrow 0}\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=0.$
@@
 where the 3rd equation is obtained by (4.3), and the last one is by Corollary 4.3 and Lemma 4.8. Since $D^{(r)}(0,\theta)=O(r^{2})$ from (4.3), we have
-$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\qquad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
+$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\quad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
 because $\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are nonnegative.
@@
 Recall that, if one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
-$\Big{(}\!-b_{1}\theta_{1}+\frac{1}{2}c_{1}\sigma_{1}^{2}\theta_{1}^{2}\Big{)}\widetilde{\varphi}_{1}(\theta_{1})+\Big{(}\!-b_{2}\theta_{2}+\frac{1}{2}c_{2}\sigma_{2}^{2}\theta_{2}^{2}\Big{)}\widetilde{\varphi}_{2}(\theta_{2})$
-$\;\;+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)+\beta_{2}e^{\theta_{2}\ell_{0}}\theta\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\quad\theta\in\mathbb{R}\times\mathbb{R}_{-},$ (5.1)
+$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
+$\qquad+\beta_{2}e^{\theta_{2}\ell_{0}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (5.1)
 which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
@@
 So, we only consider the case that $b_{1}\not=0$.
-We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflecting diffusion process.
+We like to identify this stationary distribution $\widetilde{\nu}$ by that of a reflected diffusion process.
 Assume that the diffusion scaled process $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to a nonnegative valued stochastic process $Z(\cdot)\equiv\{Z(t);t\geq 0\}$ as $r\downarrow 0$.
@@
 To this end, define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
-$\sigma(x)=\sqrt{c_{1}}\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\sigma_{2}1(x>\ell_{0}),$
-$b(x)=b_{1}1(x\leq\ell_{0})+b_{2}1(x>\ell_{0}),$
+$b(x)=b_{1}\mu 1(x\leq\ell_{0})+b_{2}\mu 1(x>\ell_{0}),$
+$\sigma(x)=\sqrt{c_{1}}\mu\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\mu\sigma_{2}1(x>\ell_{0}),$
 then one may expect that $Z(\cdot)$ satisfies the following stochastic integral equation.
@@
 However, $Y(\cdot)$ may not be absolutely continuous.
-So far, in the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered the stationary distribution using the generalized Ito formula for a convex function.
-We got a stationary equation similar to (5.1) for the $Z(\cdot)$, but could not completely obtained the stationary distribution although we had partial information which is compatible with Theorem 3.1.
+In the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered its stationary distribution using the Ito formula, which yields
+$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\in\mathbb{R}.$ (5.3)
+where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$.
+However, we had not solved this stationary equation.
 So, it was conjectured that
@@
     the stationary distribution of $Z(t)$ uniquely exists under the stability condition $b_{2}>0$, and agrees with $\widetilde{\nu}$ of (3.8) in Theorem 3.1 under the one of conditions in (ii) of it in addition to $b_{2}>0$.
-During the reviewing process of this paper, the author have known that (a) is resolved by Atar et al.
-[1, Lemma 4.3] (see also [12] for a different proof), while the author affirmatively answers (c) in [12].
-Those results are obtained for more general reflecting diffusion processes than $Z(\cdot)$, which are hard to include in this paper.
-Thus, we refer to [12] for them, and will not further discuss about the conjectures (a) and (c).
-As for the conjecture (b), it is still open (see [12]).
+During the review process of this paper, the author have known that (a) is resolved by Atar et al.
+[1, Lemma 4.3] (see also Miyazawa [11] for a different proof), while (c) can be well answered because the stationary distribution of the $Z(\cdot)$ is fully obtained in [11], in which it is claimed that the limiting distribution $\widetilde{\nu}$ of this paper is identical with the stationary distribution of the $Z(\cdot)$, but this claim is only true if and only if $c_{1}=c_{2}$.
+This fact also can be seen from the comparison of (5.3) with (5.1), which does not require the $\widetilde{\nu}$ itself.
+However, the $\widetilde{\nu}$ has not yet been completely specified.
+As for (a), the stationary distributions are obtained for more general reflected diffusion processes than $Z(\cdot)$.
+We refer to [11, 13] for them, and will not further discuss about the conjectures (a) and (c).
+As for the conjecture (b), it is still open.
 5.2 Intuitive verification of (3.12)
@@
 Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{0}$ from below or above.
-Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ after time $B^{(r)}_{2}$.
-Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, we have
-$\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (5.3)
+Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ at time $\tau^{(r)}_{2}$ after time $B^{(r)}_{2}$.
+Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $f^{(r)}_{\rm up}\equiv\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, the mean sojourn time for $L^{(r)}(t)$ to be above $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
+$f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (5.4)
 where $\sim$” stands for the same order as $r\downarrow 0$, that is, for non-zero valued functions $f_{1}$ and $f_{2}$, $f_{1}(r)\sim f_{2}(r)$ means that $\lim_{r\downarrow 0}f_{1}(r)/f_{2}(r)=1$.
 On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below.
-Then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [11] that
-$\mathbb{E}[B^{(r)}_{1}]\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (5.4)
-Since $B^{(r)}_{2}+B^{(r)}_{1}$ is the cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
-$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{\mathbb{E}[B^{(r)}_{1}]}{\mathbb{E}[B^{(r)}_{1}]+\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
-Thus, for $b_{1}\not=0$, (3.12) is intuitively obtained for $i=1$, and it is similarly obtained for $i=2$. The case for $b_{1}=0$ is similarly derived. Thus, we have another support of (3.12).
+Let $f^{(r)}_{\rm down}=1-f^{(r)}_{\rm up}$, then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that the mean sojourn time for $L^{(r)}(t)$ to be below $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
+$f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]$ $\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}$
+$=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (5.5)
+Since $f^{(r)}_{\rm up}B^{(r)}_{2}+f^{(r)}_{\rm down}B^{(r)}_{1}$ is the mean cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
+$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]}{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]+f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
+Thus, for $b_{1}\not=0$, (3.12) is intuitively obtained for $i=1$, and it is similarly obtained for $i=2$. The case for $b_{1}=0$ is similarly derived. Thus, we have support of (3.12).
 Appendix
@@
 We first consider this 2-level $M/M/1$ queue without index $r$ for notational convenience.
-Because of the exponential assumptions, we can describe this model by the continuous time Markov chain with state space
-$S=\{(m,i);m=1,2,i=0,1,\ldots,\ell_{0}\}\cup\{(3,i);i\geq\ell_{0}+1\},$
-and its transition diagram is depicted in Figure A.1.
+Because of the exponential assumption, this model can be described by a continuous-time Markov chain.
+However, $L(\cdot)$ itself cannot be a Markov chain because the arrival rate may depend on the queue length at the last arrival instant as discussed at the end of Section 3.2.
+Taking this into account, we describe this model by the continuous time Markov chain with state space
+$S=\{(j,\ell);j=1,2,\ell=0,1,\ldots,\ell_{0}\}\cup\{(3,\ell);\ell\geq\ell_{0}+1\},$
+and its transition diagram is depicted in Figure A.1. Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$.
 Figure A.1: Transition diagram
-Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$.
-Note that arrivals may find the queue length less than $\ell_{0}$ when their last arrivals found the queue length not less than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
-Because of this, the arrival rate at any time may depend on the queue length at the last arrival instant, and therefore $L(\cdot)$ cannot be a continuous-time Markov chain.
-We partition the state space $S$ of $X_{\exp}(\cdot)$ to the following three disjoint subsets.
-$S_{m}=\{(m,i)\in S;i\in[0,\ell_{0}]\},\quad m=1,2,\qquad S_{3}=\{(3,i):i\geq\ell_{0}+1\}.$
-By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{m,i}$ as
-$p_{m,i}=\mathbb{P}[J=m,L=i],\qquad m=1,2,3,(m,i)\in S_{m}.$
-Obviously, the stationary distribution, $\{p_{m,i};(m.i)\in S\}$, is uniquely determined by the following stationary equations. For state in $S_{1}$,
+We partition the state space $S$ of $X_{\exp}(\cdot)$ into the following three disjoint subsets.
+$S_{j}=\{(j,\ell)\in S;\ell\in[0,\ell_{0}]\},\quad j=1,2,\qquad S_{3}=\{(3,\ell):\ell\geq\ell_{0}+1\}.$
+By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{j,\ell}$ as
+$p_{j,\ell}=\mathbb{P}[J=j,L=\ell],\qquad j=1,2,3,(j,\ell)\in S_{j}.$
+Obviously, the stationary distribution, $\{p_{j,\ell};(j,\ell)\in S\}$, is uniquely determined by the following stationary equations. For state in $S_{1}$,
 $\lambda_{1}p_{1,0}=c_{1}\mu p_{1,1},$ (A.1)
-$(\lambda_{1}+c_{1}\mu)p_{1,i}=\lambda_{1}p_{1,i-1}+c_{1}\mu p_{1,i+1}+\lambda_{2}p_{2,i-1},\quad i\in[1,\ell_{0}-1],$ (A.2)
+$(\lambda_{1}+c_{1}\mu)p_{1,\ell}=\lambda_{1}p_{1,\ell-1}+c_{1}\mu p_{1,\ell+1}+\lambda_{2}p_{2,\ell-1},\quad\ell\in[1,\ell_{0}-1],$ (A.2)
 $(\lambda_{1}+c_{1}\mu)p_{1,\ell_{0}}=\lambda_{1}p_{1,\ell_{0}-1}+\lambda_{2}p_{2,\ell_{0}-1},$ (A.3)
@@
 $\lambda_{2}p_{2,0}=c_{1}\mu p_{2,1},$ (A.4)
-$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,i+1},\qquad i\in[1,\ell_{0}-1],$ (A.5)
+$(\lambda_{2}+c_{1}\mu)p_{2,\ell}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{0}-1],$ (A.5)
 $(\lambda_{2}+c_{1}\mu)p_{2,\ell_{0}}=c_{2}\mu p_{3,\ell_{0}+1},$ (A.6)
-For state in $S_{3}$, it follows from the equation for $\ell_{0}+1$ and the local balance that
+For states in $S_{3}$, it follows from the equation for $(3,\ell_{0}+1)$ and the local balance that
 $(\lambda_{2}+c_{2}\mu)p_{3,\ell_{0}+1}=\lambda_{1}p_{1,\ell_{0}}+\lambda_{2}p_{2,\ell_{0}}+c_{2}\mu p_{\ell_{0}+2},$ (A.7)
-$\lambda_{2}p_{3,i}=c_{2}\mu p_{3,i+1},\qquad i\geq\ell_{0}+1.$ (A.8)
+$\lambda_{2}p_{3,\ell}=c_{2}\mu p_{3,\ell+1},\qquad\ell\geq\ell_{0}+1.$ (A.8)
 In what follows, we use the following notations.
-$\rho_{ij}=\frac{\lambda_{i}}{c_{j}\mu},\qquad\gamma_{ij}=\frac{c_{i}\mu}{c_{1}\mu+\lambda_{j}},\qquad i,j=1,2.$
+$\rho_{j,k}=\frac{\lambda_{j}}{c_{k}\mu},\qquad\gamma_{j,k}=\frac{c_{j}\mu}{c_{1}\mu+\lambda_{k}},\qquad j,k=1,2.$
 We first note that, from (A.8),
-$p_{3,i}=\rho_{22}^{i-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad i\geq\ell_{0}+1.$ (A.9)
-Thus, $p_{3,i}$ for $i\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
-We next consider $p_{2,i}$ for $i\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
+$p_{3,\ell}=\rho_{22}^{\ell-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad\ell\geq\ell_{0}+1.$ (A.9)
+Thus, $p_{3,\ell}$ for $\ell\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
+We next consider $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
 $p_{2,0}=\frac{c_{1}\mu}{\lambda_{2}}p_{2,1}=\rho_{21}^{-1}p_{2,1},$ (A.10)
-$p_{2,i}=\gamma_{12}p_{2,i+1}=\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}},\quad i\in[1,\ell_{0}-1].$ (A.11)
-Thus, $p_{2,i}$ for $i\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written as
-$p_{2,i}=\gamma_{12}^{-1}p_{2,i-1}=\gamma_{12}^{-i+1}p_{2,1}=\gamma_{12}^{-i+1}\rho_{21}p_{2,0},\quad i\in[1,\ell_{0}-1].$ (A.12)
-Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{m}$ for $m=1,2,3$.
+$p_{2,\ell}=\gamma_{12}p_{2,\ell+1}=\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}},\quad\ell\in[1,\ell_{0}-1].$ (A.11)
+Thus, $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written by (A.4) as
+$p_{2,\ell}=\gamma_{12}^{-1}p_{2,\ell-1}=\gamma_{12}^{-\ell+1}p_{2,1}=\gamma_{12}^{-\ell+1}\rho_{21}p_{2,0},\quad\ell\in[1,\ell_{0}-1].$ (A.12)
+Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{j}$ for $j=1,2,3$.
 We call them boundary probabilities.
@@
 $p_{2,\ell_{0}}=\gamma_{22}p_{3,\ell_{0}+1}.$ (A.13)
-Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $i=\ell_{0}+1$, (A.7) can be written as
+Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $\ell=\ell_{0}+1$, (A.7) can be written as
 $p_{3,\ell_{0}+1}=\rho_{12}p_{1,\ell_{0}}+\rho_{22}p_{2,\ell_{0}},$ (A.14)
@@
 We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{0}}$, but this is not easy.
-To resolve this problem, we consider $p_{1,i}$ for $i\in[0,\ell_{0}-1]$. From (A.2),
-$\lambda_{1}p_{1,i}-c_{1}\mu p_{1,i+1}=\lambda_{1}p_{1,i-1}-c_{1}\mu p_{1,i}+\lambda_{2}p_{2,i-1},\qquad i\in[1,\ell_{0}-1],$
+To resolve this problem, we consider $p_{1,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.2),
+$\lambda_{1}p_{1,\ell}-c_{1}\mu p_{1,\ell+1}=\lambda_{1}p_{1,\ell-1}-c_{1}\mu p_{1,\ell}+\lambda_{2}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1],$
 then this equation can be written as
-$\rho_{11}p_{1,i}-p_{1,i+1}=\rho_{11}p_{1,i-1}-p_{1,i}+\rho_{21}p_{2,i-1},\qquad i\in[1,\ell_{0}-1].$
-Summing up both sides of this equation for $i\in[1,j]$ for $j\in[1,\ell_{0}-1]$, we have
-$\rho_{11}p_{1,j}-p_{1,j+1}=\rho_{21}\sum_{i=1}^{j}p_{2,i-1},\qquad j\in[0,\ell_{0}-1],$ (A.16)
-because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $j=0$ by the convention that the empty sum vanishes, which we always take.
+$\rho_{11}p_{1,\ell}-p_{1,\ell+1}=\rho_{11}p_{1,\ell-1}-p_{1,\ell}+\rho_{21}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1].$
+Summing up both sides of this equation for $\ell\in[1,m]$ for $m\in[1,\ell_{0}-1]$, we have
+$\rho_{11}p_{1,m}-p_{1,m+1}=\rho_{21}\sum_{\ell=1}^{m}p_{2,\ell-1},\qquad m\in[0,\ell_{0}-1],$ (A.16)
+because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $m=0$ by the convention that the empty sum vanishes, which we always take.
 Hence, applying (A.10) and (A.11) to (A.16), we have
-$\rho_{11}p_{1,j}-p_{1,j+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{i=1}^{j-1}\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}}\right)$
-$=\left(1+\rho_{21}\sum_{i=1}^{j-1}\gamma_{12}^{-(i-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
-$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[1,\ell_{0}-1].$
-Multiplying both sides by $\rho_{11}^{-(j+1)}$,
-$\rho_{11}^{-j}p_{1,j}-\rho_{11}^{-(j+1)}p_{1,j+1}$
-$\quad=\rho_{11}^{-(j+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(j\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[0,\ell_{0}-1].$
-Summing up both sides of this equation from $j=0$ to $j=k-1\leq\ell_{0}-1$, we have
-$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{j=0}^{k-1}\rho_{11}^{-(j+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{j=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(j-1)}-\rho_{11}^{-(j-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
+$\rho_{11}p_{1,m}-p_{1,m+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{\ell=1}^{m-1}\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}}\right)$
+$=\left(1+\rho_{21}\sum_{\ell=1}^{m-1}\gamma_{12}^{-(\ell-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
+$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[1,\ell_{0}-1].$
+Multiplying both sides by $\rho_{11}^{-(m+1)}$,
+$\rho_{11}^{-m}p_{1,m}-\rho_{11}^{-(m+1)}p_{1,m+1}$
+$\quad=\rho_{11}^{-(m+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(m\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[0,\ell_{0}-1].$
+Summing up both sides of this equation from $m=0$ to $m=k-1\leq\ell_{0}-1$, we have
+$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{m=0}^{k-1}\rho_{11}^{-(m+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{m=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(m-1)}-\rho_{11}^{-(m-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
 $\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}}$
@@
 For the $r$-th system, we put the superscript (r) as we have used for the $r$-th 2-level $GI/G/1$ queue.
-For example, $p^{(r)}_{m,i}$ is the stationary probability of $\{(J,L)=(m,i)\}$ for $(m,i)\in S$.
-Define the moment generating functions $g^{(r)}_{m}(\theta)$ for $m=1,2,3$ as
-$g^{(r)}_{m}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{m,k},\quad\theta\in\mathbb{R},m=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
-and define $g^{(r)}(\theta)=\sum_{m=1}^{3}g^{(r)}_{m}(\theta)$.
+For example, $p^{(r)}_{j,\ell}$ is the stationary probability of $\{(J,L)=(j,\ell)\}$ for $(j,\ell)\in S$.
+Define the moment generating functions $g^{(r)}_{j}(\theta)$ for $j=1,2,3$ as
+$g^{(r)}_{j}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{j,k},\quad\theta\in\mathbb{R},j=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
+and define $g^{(r)}(\theta)=\sum_{j=1}^{3}g^{(r)}_{j}(\theta)$.
 Obviously, $g^{(r)}(\theta)$ is the moment generating function of $rL^{(r)}$.
 Thus, our task is to compute the limit of $g^{(r)}(\theta)$ as $r\downarrow 0$.
-To this end, we compute $g^{(r)}_{m}(\theta)$ for $m=1,2,3$, using the normalization condition that $g^{(r)}(0)=1$.
+To this end, we compute $g^{(r)}_{j}(\theta)$ for $j=1,2,3$, using the normalization condition that $g^{(r)}(0)=1$.
 However, it should be noted that this condition is not sufficient to prove the weak convergence of the distribution of $rL^{(r)}$, for which it is required that $\lim_{\theta\uparrow 0}\lim_{r\downarrow 0}g^{(r)}(\theta)=1$.
-This weak convergence will be verified through explicitly computing $g^{(r)}_{m}(\theta)$ for $m=1,2,3$.
-We first compute $g^{(r)}_{m}(\theta)$ for $m=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
-$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{i=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-i)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-i}p^{(r)}_{2,\ell^{(r)}_{0}}$
+This convergence will be verified through explicitly computing $g^{(r)}_{j}(\theta)$ for $j=1,2,3$.
+We first compute $g^{(r)}_{j}(\theta)$ for $j=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
+$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{\ell=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-\ell)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-\ell}p^{(r)}_{2,\ell^{(r)}_{0}}$
 $=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{0}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}},$ (A.19)
@@
 $\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$ (A.21)
-We now compute the limits of $g^{(r)}_{m}(0)$ for $m=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
+We now compute the limits of $g^{(r)}_{j}(0)$ for $j=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
 $(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=O(\gamma_{12}^{\ell_{0}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.22)
 Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}$ decays exponentially fast as the function of $1/r$. On the other hand,
-$\lim_{r\downarrow 0}(1-\rho^{(r)}_{i})/r=\beta_{i},\qquad i=1,2,$ (A.23)
+$\lim_{r\downarrow 0}(1-\rho^{(r)}_{\ell})/r=\beta_{\ell},\qquad\ell=1,2,$ (A.23)
 $\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{0}/r+o(r)}=e^{-\beta_{1}\ell_{0}},$ (A.24)
-where recall that $\beta_{i}=b_{i}/c_{i}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.
+where recall that $\beta_{\ell}=b_{\ell}/c_{\ell}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.
 Lemma A.1.

### v7

@@
 This limit corresponds to its diffusion approximation.
-It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters.
-This result is proved using the so called BAR approach studied in [2, 3, 10, 12].
+It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters and generally distributed inter-arrival times and workloads brought by customers.
+This result is proved under a mild limitation on arrival parameters using the so called BAR approach studied in [3, 4, 10, 12].
 We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
-Keywords: single server queue; level dependent arrival and service; stationary distribution; heavy traffic; diffusion approximation; Palm distribution; stochastic integral equation; multi-level SRBM.
 1 Introduction
@@
 So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length.
-However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured.
+However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured in this exponential case.
 Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution.
-This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold.
+This limiting distribution corresponds to the stationary distribution of a one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold, which is recently studied in [11, 13].
 Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
@@
 In general, theoretical study of diffusion approximation is hard for state-dependent queues.
-For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see Mandelbaum and Pats [8], Yamada [14]), while the arrival process and service rate discontinuously change in the present model.
+For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see [8, 15]), while the arrival process and service rate discontinuously change in the present model.
 Thus, theoretical study for its diffusion approximation is challenging.
-As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see Kingman [7]).
-Furthermore, Miyazawa [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room.
-Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]).
+As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see [7]).
+Furthermore, [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room.
+Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see [6]).
 Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it.
@@
 Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [3, 4, 10, 12].
-This verification, Theorem 3.1, requires either some limitation on system parameters or for the workload brought by a customer to be exponentially distributed.
-The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
+This verification, Theorem 3.1, requires a mild limitation on the parameters of the arrival processes (see (3.f)).
+Its proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
+However, the proof has a complicated structure, so we split it into many components so that they are orthogonal as long as possible.
+This is the reason why we spend two sections for the proof.
 In this paper, we only study the limit of the scaled stationary distribution, but, as our results are related to the diffusion approximation, it is interesting to identify a reflected Brownian motion as a process limit for the 2-level $GI/G/1$ queue.
@@
 We intuitively discuss what can be expected about this process limit in Section 6.1.
-This paper is made up by five sections.
+This paper is made up by six sections.
 We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1.
 A main result, Theorem 3.1, and the exponentially distributed case, Proposition 3.1, are presented in Section 3.2.
-In Section 4, we prove Theorem 3.1.
+In Section 4, we introduce the so called BAR approach, and apply it to the 2-level $GI/G/1$ queue.
+Using these results, Theorem 3.1 is proved in Section 5.
 Finally, in Section 6, we intuitively discuss two topics.
@@
 We denote those point processes by $N_{e}$ and $N_{s}$ respectively.
-According to Miyazawa [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
+According to [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
 (2.a) Measurability: $A\in{\mathcal{F}}_{t}\;\mbox{ implies }\;\Theta_{s}^{-1}A\in{\mathcal{F}}_{s+t},\qquad s,t\geq 0,$
@@
 Hence, $\lambda_{1}=\lambda_{2}$ if and only if $c_{1}=c_{2}$.
+Most of our arguments are valid under the assumptions (3.a)–(3.e), but we require the following condition for their final step to determine the limiting distribution of the scaled queue length in heavy traffic.
+  • (3.f)
+    $\sigma_{e,1}^{2}=\sigma_{e,2}^{2}$ if $\lambda_{1}=\lambda_{2}$.
+Remark 3.1.
+The condition (3.f) comes out from technical requirements for the proof of Theorem 3.1.
+It is a mild limitation because no condition is needed for $\lambda_{1}\not=\lambda_{2}$, but we think that it is likely not needed.
+However, removing (3.f) seems to be beyond the framework of this paper.
+See Remark 5.1 for its details.
 We take countably many $r$’s for a limiting operation.
@@
 Lemma 3.1.
-Under the stationary assumption, $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
+Assume the assumptions (3.a)–(3.e) and the stationarity of $X^{(r)}(\cdot)$, then $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
 Proof.
@@
 Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$.
-Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}]<\infty$.
+Similarly, let $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1}))^{-1}T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1})^{-1})T_{s}]<\infty$.
 ∎
@@
 Define probability distribution $\nu^{(r)}$ under the assumptions (3.a)–(3.e) as
-$\nu^{(r)}(B)=\mathbb{P}(rL^{(r)}\in B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.1)
+$\nu^{(r)}(B)=\mathbb{P}(rL^{(r)}\in B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$
 We refer to this $\nu^{(r)}$ as a scaled stationary distribution.
 We decompose $\nu^{(r)}$ using finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$, which are defined as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
-$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}>\ell^{(r)}_{0}),$
-$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}).$
+$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;\ell^{(r)}_{0}<L^{(r)}),$
+$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|\ell^{(r)}_{0}<L^{(r)}).$
 Theorem 3.1.
-Assume the assumptions (3.a)–(3.e), and let
+Assume the assumptions (3.a)–(3.f). Let
 $\sigma_{i}^{2}=\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2},\qquad\beta_{i}=\frac{2b_{i}}{c_{i}\sigma_{i}^{2}},\qquad i=1,2.$ (3.1)
@@
   • (i)
-    For each $i=1,2$, if $\nu^{(r)}_{i}(\mathbb{R}_{+})$ is bounded away from 00 as $r\downarrow 0$, then $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$, namely,
+    For each $i=1,2$, if $\nu^{(r)}_{i}(\mathbb{R}_{+})$ is bounded away from 00 as $r\downarrow 0$, that is,
+    $\liminf_{r\downarrow 0}\varphi^{(r)}_{i}(0)>0,\qquad i=1,2,$ (3.6)
+    then $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$, namely,
     $\lim_{r\downarrow 0}\nu_{i|cd}^{(r)}(B)=\widetilde{\nu}_{i|cd}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.7)
@@
     $\widetilde{A}_{1}=\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)},\;&b_{1}\not=0,\end{cases}\qquad\widetilde{A}_{2}=1-\widetilde{A}_{1}.$ (3.10)
-    If either $\lambda_{1}=\lambda_{2}$ (equivalently, $c_{1}=c_{2}$) or $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,1]$, then $\alpha^{(r)}_{e}$ converges to $\widetilde{\alpha}_{e}$ and $\nu^{(r)}$ weakly converges to $\widetilde{\nu}$, namely
-    $\lim_{r\downarrow 0}\nu^{(r)}(B)=\widetilde{\nu}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.11)
+    If either $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$, equivalently, $c^{(r)}_{1}-c^{(r)}_{2}=O(r)$, or $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,1]$, then
+    $\lim_{r\downarrow 0}\nu^{(r)}(B)=\widetilde{\nu}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}),$ (3.11)
+    $\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\widetilde{\alpha}_{e}.$ (3.12)
   • (iii)
@@
 Corollary 3.1.
-Under either one of the assumptions in (ii) of Theorem 3.1,
+Under the assumptions (3.a)–(3.f) and either one of the assumptions in (ii) of Theorem 3.1,
 $\lim_{r\downarrow 0}(1-\rho^{(r)})\mathbb{E}[L^{(r)}]=\lim_{r\downarrow 0}\mathbb{P}(L^{(r)}=0)\mathbb{E}[L^{(r)}]$
 =$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.15)
-Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
+Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed.
+We call this model the $r$-th 2-level $M/M/1$ queue.
+Note that (3.f) is automatically satisfied in this exponential model.
 Proposition 3.1.
-Under the settings of Theorem 3.1, the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
+Under the assumptions (3.a)–(3.e), the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
 $\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{0}}{1+\beta_{2}\ell_{0}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.16)
@@
 However, for a double check, we give its direct proof in Appendix A.2, which is helpful to convince (ii) of Theorem 3.1.
-This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of more than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
+This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of greater than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
 This Markov chain will be detailed in Appendix A.2.
-4 Proofs
-In this section we prove Theorem 3.1.
+4 BAR’s for the 2-level $GI/G/1$ queue
+In this section, we prepare auxiliary notations and results for proving Theorem 3.1.
 A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$.
@@
 Recently, this approach is further studied in [12].
-The proof of Theorem 3.1 is composed of three parts. We first introduce the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
+Throughout this section, we assume (3.a)–(3.e), but (3.f) is not assumed. The proof of Theorem 3.1 will be done in Section 5.
 4.1 Palm distribution and BAR
@@
 Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
-${\mathcal{H}}f({\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f({\boldsymbol{x}})-(c_{1}1(1\leq x_{1}\leq\ell_{0})+c_{2}1(x_{1}>\ell_{0}))\frac{\partial^{+}}{\partial x_{3}}f({\boldsymbol{x}}),\qquad{\boldsymbol{x}}\in S,$
-where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$.
+${\mathcal{H}}f(z,{\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}})-(c_{1}1(1\leq z\leq\ell_{0})+c_{2}1(z>\ell_{0}))\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}}),\quad(z,{\boldsymbol{x}})\in S,$
+where ${\boldsymbol{x}}=(x_{1},x_{2})\in\mathbb{R}_{+}^{2}$.
 In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit.
@@
 $\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]=0,\qquad f\in C_{b}^{p},$ (4.2)
-The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short.
+The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short, then we call analysis based on this BAR as a BAR approach.
+A major difficulty in this BAR approach is to evaluate the expectations under the Palm distributions (the last two terms in (4.1)).
+This evaluation will be done for the 2-level $GI/G/1$ queue in Section 4.3.
 Remark 4.1.
@@
   • (i)
-    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distribution $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$.
-So, $X^{(r)}(0-)$ can be considered as a random variable generated by $\mathbb{P}^{(r)}_{e}$.
-However, one may be uncomfortable for this interpretation.
-In this case, please consider the process $X(t)$ to start at time $-1$ subject to the stationary distribution, then the distribution of $X(\cdot)=\{X(t);t\geq 0\}$ is unchanged because $X(\cdot)$ is a stationary process.
-In either way, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions.
+    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$.
+So, $X^{(r)}(0-)$ is a random variable well defined under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$.
+Hence, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions.
+If we define the stationary process $X^{(r)}(\cdot)$ on the time axis (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ are well defined even under $\mathbb{P}$.
   • (ii)
@@
 Proof.
-For each integer $k\geq 0$, let $f({\boldsymbol{x}})=x_{1}\wedge(k+1)$ for ${\boldsymbol{x}}\in S$, then obviously $f\in C_{b}^{p}(S)$.
+For each integer $k\geq 0$, let $f(z,{\boldsymbol{x}})=z\wedge(k+1)$ for $(z,{\boldsymbol{x}})\in S$, then obviously $f\in C_{b}^{p}(S)$.
 Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
 $\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]$
-$\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
+$\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{P}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
 Applying this to (4.1) yields
-$\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
-Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$. Because of this equality, the above formula also yields (4.3). We next let $f({\boldsymbol{x}})=x_{2}$, and apply (4.1), then
+$\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{P}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
+Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$.
+Because of this equality, the above formula also yields (4.3).
+We next let $f(z,{\boldsymbol{x}})=x_{1}$, and apply (4.1), then
 $-1+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[T^{(r)}_{e,1}1(L^{(r)}(0-)<\ell^{(r)}_{0})+T^{(r)}_{e,2}1(L^{(r)}(0-)\geq\ell^{(r)}_{0})\right]=0$
-because $(R^{(r)}_{e})^{\prime}(t)=-1$. This proves (4.4). Similarly, let $f({\boldsymbol{x}})=x_{3}$, then
+because $(R^{(r)}_{e})^{\prime}(t)=-1$. This proves (4.4). Similarly, let $f(z,{\boldsymbol{x}})=x_{2}$, then
 $(R^{(r)}_{s})^{\prime}(t)=-c_{1}^{(r)}1(1\leq L^{(r)}(0)\leq\ell^{(r)}_{0})-c_{2}^{(r)}1(L^{(r)}(0)>\ell^{(r)}_{0}),$ (4.6)
@@
 To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1.
-We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
-4.2 Preliminary lemmas
+We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 5.
+4.2 Asymptotic BAR for the pre-limit process $X^{(r)}(\cdot)$
 We will prove Theorem 3.1 using the moment generating functions $\varphi^{(r)}_{i}(\theta_{i})$ and $\varphi^{(r)}_{i|cd}(\theta_{i})$ of $\nu^{(r)}_{i}$ and $\nu^{(r)}_{i|cd}$, respectively, for $i=1,2$.
 Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
-$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(L^{(r)}>\ell^{(r)}_{0})],$
-$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
+$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(\ell^{(r)}_{0}<L^{(r)})],$
+$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|\ell^{(r)}_{0}<L^{(r)}].$
 To this end, we introduce test functions for the BAR (4.1). We first define functions $w^{(r)}_{i}$ and $g^{(r)}_{i,\theta}$ for $i=1,2$ and $\theta\in\mathbb{R}$ as
@@
 Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin.
-It is formally proved in [3, Lemma 5.8].
+It is formally proved in [4, Lemma 5.8].
 Lemma 4.2.
-Under the assumptions (3.a)–(3.d), we have the following expansions as $r\downarrow 0$.
-$\eta_{i}^{(r)}(r\theta)=\lambda^{(r)}_{i}r\theta+\frac{1}{2}(\lambda^{(r)}_{i})^{3}(\sigma^{(r)}_{e,i})^{2}r^{2}\theta^{2}+o(r^{2}),\qquad i=1,2,$ (4.10)
-$\zeta^{(r)}(r\theta)=-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}r^{2}\theta^{2}+o(r^{2}).$ (4.11)
+Under the assumptions (3.a)–(3.d), we have the following expansions as $r\downarrow 0$ for each fixed $\theta\in\mathbb{R}$.
+$\eta_{i}^{(r)}(r\theta)=\lambda^{(r)}_{i}r\theta+\frac{1}{2}(\lambda^{(r)}_{i})^{3}(\sigma^{(r)}_{e,i})^{2}r^{2}\theta^{2}+o((\theta r)^{2}),\qquad i=1,2,$ (4.10)
+$\zeta^{(r)}(r\theta)=-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}r^{2}\theta^{2}+o((\theta r)^{2}).$ (4.11)
 Furthermore, for $i=1,2$, there are constant $d_{e,i},d_{s},a>0$ such that, for $r\in(0,1]$ and $\theta\in\mathbb{R}$ satisfying $r|\theta|<a$,
-$\left|\eta^{(r)}_{i}(r\theta)(u_{1}\wedge 1/r)+\zeta^{(r)}(r\theta)(u_{2}\wedge 1/r)\right|\leq|\theta|\left(d_{e,i}(ru_{1}\wedge 1)+d_{s}(ru_{2}\wedge 1)\right).$ (4.12)
+$\left|\eta^{(r)}_{i}(r\theta)(u_{1}\wedge 1/r)\right|\leq|r\theta|d_{e,i}(u_{1}\wedge 1/r),\qquad i=1,2,$ (4.12)
+$\left|\zeta^{(r)}(r\theta)(u_{2}\wedge 1/r)\right|\leq|r\theta|d_{s}(u_{2}\wedge 1/r).$ (4.13)
+Remark 4.2.
+In (4.10) and (4.11), $o((\theta r)^{2})$ can be replaced by $o(r^{2})$ because $\theta$ is fixed, but $o((\theta r)^{2})$ is convenient for our computation because we need to divide both sides of them by $\theta$ in some cases.
+We will take this convention throughout the paper as long as it is possibe.
 By this lemma, $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ is bounded by $e^{|\theta_{i}|(d_{e,i}+d_{s})}$ for sufficiently small $r$ for each $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, and therefore, we have the following facts.
@@
 $\psi^{(r)}_{i}(\theta_{i})<\infty,\quad\mathbb{E}_{u}\left[f^{(r)}_{r{\boldsymbol{\theta}}}(L^{(r)},R^{(r)})\right]<\infty,\quad u=e,s,\quad\mbox{uniformly in $r\in(0,1]$},$ (4.16)
-because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
+because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$.
+These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
+We also note the following fact.
+Corollary 4.3.
+For $\theta_{1}\in\mathbb{R}_{+}$ and $\theta_{2}\leq 0$
+$\lim_{r\downarrow 0}\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|=0,\quad i=1,2.$ (4.17)
+Proof.
+By Corollary 4.2, we have
+$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{0})]\right|$
+$\leq e^{|\theta_{1}|r\ell^{(r)}_{0}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
+This proves (4.17). ∎
 Based on these facts, we compute $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$ in (4.1) for $f^{(r)}_{r{\boldsymbol{\theta}}}$. Then,
@@
 $\qquad+\sum_{i=1,2}c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})\psi^{(r)}_{i}(\theta_{i})-c^{(r)}_{1}\zeta^{(r)}(r\theta_{1})\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)}))\right].$ (4.18)
-Since $\lambda^{(r)}_{i}-c^{(r)}_{i}\mu^{(r)}=-b_{i}\mu^{(r)}r+o(r)$, it follows from Lemma 4.2, $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ and $\lambda_{i}=c_{i}\mu$ that
-$\eta^{(r)}_{i}(r\theta_{i})+c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})$ $=-b_{i}r^{2}\mu\theta_{i}+\frac{1}{2}c_{i}(\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2})r^{2}\mu\theta_{i}^{2}+o(r^{2})$
-$=-\frac{1}{2}c_{i}\sigma_{i}^{2}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o(r^{2}).$
+Since $\lambda^{(r)}_{i}-c^{(r)}_{i}\mu^{(r)}=-b_{i}\mu^{(r)}r+o(r)$, it follows from Lemma 4.2, $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ and $\lambda_{i}=c_{i}\mu$ that, each fixed $\theta_{i}\in\mathbb{R}$,
+$\eta^{(r)}_{i}(r\theta_{i})+c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})$ $=-b_{i}r^{2}\mu\theta_{i}+\frac{1}{2}c_{i}(\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2})r^{2}\mu\theta_{i}^{2}+o((\theta_{i}r)^{2})$
+$=-\frac{1}{2}c_{i}\sigma_{i}^{2}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta_{i}r)^{2}),\qquad r\downarrow 0.$
 Hence, from (4.2), we have
 $\mathbb{E}\left[{\mathcal{H}}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]=\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\left(-\beta_{i}+\theta_{i}\right)\mu\theta_{i}r^{2}\psi^{(r)}_{i}(\theta_{i})$
-$\qquad+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}).$ (4.17)
+$\quad+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o((\theta_{1}r)^{2})+o((\theta_{2}r)^{2}).$ (4.19)
 We next compute the Palm expectation terms in (4.1). In what follows, we use the following abbreviated notations.
-$\widehat{T}^{(r)}_{e,i}\equiv T^{(r)}_{e,i}(0)\wedge 1/r,\quad i=1,2,\qquad\widehat{T}^{(r)}_{s}\equiv T^{(r)}_{s}(0-)\wedge 1/r,$
-$\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$
+$\widehat{T}^{(r)}_{e,i}\equiv T^{(r)}_{e,i}\wedge 1/r,\quad i=1,2,\qquad\widehat{T}^{(r)}_{s}\equiv T^{(r)}_{s}\wedge 1/r,$ (4.20)
+$\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s-}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$ (4.21)
 Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
@@
 $\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})\Big{]}$
-$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
-$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.18)
+$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
+$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.22)
 where the third equality is obtained by applying (4.8). Similarly, we have
 $\mathbb{E}^{(r)}_{s}[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{s}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0)))1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0)=\ell^{(r)}_{0}+1)\Big{]}$
+$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0-)=\ell^{(r)}_{0}+1)\Big{]}$
 $\quad=e^{r\theta_{1}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
@@
 Then, by (4.2) and (4.2),
-$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
 $\quad+\alpha^{(r)}_{e}(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
@@
 For each fixed $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, we have, as $r\downarrow 0$,
-$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o(r^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
-$\quad-c_{1}\left(\mu r\theta_{1}+\mu^{3}\sigma_{s}^{2}r^{2}\theta_{1}^{2}/2+o(r^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})]=D^{(r)}(\theta_{1},\theta_{2}).$ (4.21)
+$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
+$\quad-c_{1}\left(\mu r\theta_{1}-\mu^{3}\sigma_{s}^{2}r^{2}\theta_{1}^{2}/2+o((\theta_{1}r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})]=D^{(r)}(\theta_{1},\theta_{2}).$ (4.25)
 This BAR is called an asymptotic BAR, and a starting point of our proof of Theorem 3.1.
-We will divide this BAR by $r^{2}$, and take its limit as $r\downarrow 0$, expanding the exponential functions in (4.3).
-To perform this expansion, we consider asymptotic behaviors on $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ as $r\downarrow 0$.
+4.3 Expansion of the asymptotic BAR
+We aim to derive the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{2}(\theta_{2})$ from the BAR (4.3).
+For this, we will divide it for either $\theta_{2}=0$ or $\theta_{1}=0$ by $r^{2}$, and take its limit as $r\downarrow 0$, expanding the exponential functions in (4.3).
+We perform this computations in three steps.
+Recall that (3.a)–(3.e) are assumed but (3.f) is not assumed throughout Section 4.
+(Step 1)   In this step, we show that the moments of $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ are well behaved as $r\downarrow 0$.
+This is also examined for $\widehat{R}^{(r)}_{s-}1(L^{(r)}=\ell^{(r)}_{0})$ and $\widehat{R}^{(r)}_{e}1(L^{(r)}=\ell^{(r)}_{0-})$ under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$, respectively.
+These are basis for the expansion of the BAR.
+We then introduce the set of key quantities, $\Delta^{(r)}_{i}$ for $i=1,2$.
 Lemma 4.4.
-Under the assumption (3.a), for $k=0,1,2$ as $r\downarrow 0$,
-$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}\in B)],$ (4.22)
-$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]$
-$\qquad\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
-and, for $i=1,2$,
+Under the assumption (3.a), for $k=0,1,2$ and $r\in(0,1]$,
+$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}],$ (4.26)
+$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}]\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}]+(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[T_{s}^{k}],$ (4.27)
+and, for $i=1,2$, as $r\downarrow 0$
 $|\mathbb{E}[(\widehat{T}^{(r)}_{u})^{k}]-\mathbb{E}[(T^{(r)}_{u})^{k}]|=\mathbb{E}[(T^{(r)}_{u})^{k}1(T^{(r)}_{u}>1/r)]=o(r^{3-k}),\quad u=(e,i),s,$ (4.28)
-$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|=\mathbb{E}[(R^{(r)}_{v})^{k}1(R^{(r)}_{v}>1/r)]=o(r^{2-k}),\qquad v=e,s,$ (4.25)
-$\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]=\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}]+o(r^{3-k}),$ (4.26)
-$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}]=\mathbb{E}^{(r)}_{e}[(R^{(r)}_{s}(0-))^{k}]+o(r^{3-k}).$ (4.27)
-In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable for $k=1,2$.
-This lemma is easily proved using the BAR (4.1), but its proof is lengthy, so we defer it to Appendix A.3.
-Because of the finiteness of the third moments in the assumption (3.a), the following corollary is immediate from Lemma 4.4.
-Corollary 4.3.
-For each $k=1,2$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
-Note that the weak limit of $\nu^{(r)}$ should be obtained from (4.3).
-However, we will see that it requires the relation between the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, which is not easy to get from (4.3).
-To relax this difficulty, we prepare two lemmas below.
+$|\mathbb{E}[(\widehat{R}^{(r)}_{v})^{k}]-\mathbb{E}[(R^{(r)}_{v})^{k}]|=\mathbb{E}[(R^{(r)}_{v})^{k}1(R^{(r)}_{v}>1/r)]=o(r^{2-k}),\qquad v=e,s.$ (4.29)
+In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly bounded for $k=1,2$ by (3.a).
+This lemma is easily proved using the BAR (4.1), but its proof is lengthy, so we defer it to Appendix A.3. We prepare one more lemma and its corollary, which are helpful for the expansion of the BAR.
 Lemma 4.5.
-Under the assumption (3.a), we have, for $k=1,2,3$ and as $r\downarrow 0$,
+Under the assumption (3.a), we have, for $k=1,2,3$ as $r\downarrow 0$,
 $c^{(r)}_{1}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(1\leq L^{(r)}\leq\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.28)
+$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.30)
 $c^{(r)}_{2}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.29)
+$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.31)
 $k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}\leq\ell^{(r)}_{0})]$
@@
 $k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)},$ (4.31)
+$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)}.$ (4.33)
 Proof.
-For $k\geq 1$, we first prove (4.5). We apply $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
+For $k\geq 1$, we first prove (4.5). We apply $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
 $\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,1\leq L^{(r)}(t)\leq\ell^{(r)}_{0}).$
-Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ and $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, respectively.
+Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ and $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, respectively.
 ∎
+The following corollary is immediate from (4.5), (4.5), (3.a) and Lemma 4.4.
+Corollary 4.4.
+For each $k=1,2,3$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
+Note that the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{2}(\theta_{2})$ should be obtained from (4.3).
+However, we will see that it requires the relation between the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, which is not easy to get from (4.3).
+To consider this problem, we introduce the following quantities.
 Define, for $i=1,2$,
-$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
 $\quad+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}].$ (4.34)
@@
 Lemma 4.6.
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-r\mu b_{1}\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.33)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.35)
 $\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]+o(r).$ (4.36)
@@
 $\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
-$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
-because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.25).
+$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
+because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r)=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r)=o(r^{2})$ by (4.29).
 This proves (4.35).
@@
 We will use Lemma 4.6 to determine the limits of $\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]$ and $\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]$, equivalently, $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, as $r\downarrow 0$ under some conditions.
-4.3 Proof of Theorem 3.1
-We first prove (i). We are now ready to take the limit of the BAR of (4.3) as $r\downarrow 0$. For this, we will use the following elementary expansions.
+(Step 2)   In this step, we expand $D^{(r)}(\theta,0)$ and $D^{(r)}(0,\theta)$ concerning $r$ as $r\downarrow 0$.
+We will see that $\Delta^{(r)}_{i}$’s describe the expansion terms of $D^{(r)}(\theta,0)$ of order $r$.
+We here will introduce ${\mathcal{E}}^{(r)}_{i}$’s for describing their expansion terms of order $r^{2}$.
+We start with the following elementary expansions. Recall Remark 4.2 for our convention of the order terms.
 Lemma 4.7.
 For each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$,
-$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}+o(r),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r)+o(r),\quad i=1,2,$ (4.35)
-$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+o(r),$ (4.36)
-$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r).$ (4.37)
+$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}(1+o(\theta_{i}r)),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r+o(\theta_{i}r)),\quad i=1,2,$ (4.37)
+$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)),$ (4.38)
+$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}$
+$\quad=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)).$ (4.39)
 Lemma 4.8.
-For each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$,
-$e^{-\eta^{(r)}_{i}(r\theta_{i})\widehat{R}^{(r)}_{e}}=1-r\lambda_{i}\theta_{i}\widehat{R}^{(r)}_{e}-2^{-1}r^{2}\lambda_{i}^{3}\sigma_{e,i}^{2}r\theta_{i}^{2}(\widehat{R}^{(r)}_{e})^{2}+o(r^{2}),$ (4.38)
-$e^{-\zeta^{(r)}(r\theta_{i})\widehat{R}^{(r)}_{s}}=1+\mu r\theta_{i}\widehat{R}^{(r)}_{s}-2^{-1}r^{i}\mu^{3}\sigma_{s}^{2}\theta_{i}^{2}(\widehat{R}^{(r)}_{s})^{2}+o(r^{2}),\qquad i=1,2,$ (4.39)
+Let $S_{n}(x)=\sum_{k=1}^{n}x^{k}$ for $x\in\mathbb{R}$ and integer $n\geq 1$, then, for each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$, for $i=1,2$,
+$e^{-\eta^{(r)}_{i}(r\theta_{i})\widehat{R}^{(r)}_{e}}=1-r\theta_{i}(\lambda^{(r)}_{i}\widehat{R}^{(r)}_{e})-2^{-1}(r\theta_{i})^{2}[(\lambda^{(r)}_{i}\sigma^{(r)}_{e,i})^{2}(\lambda^{(r)}_{i}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{i}\widehat{R}^{(r)}_{e})^{2}]$
+$\qquad+S_{3}(\widehat{R}^{(r)}_{e})o((r\theta)^{2}),$ (4.40)
+$e^{-\zeta^{(r)}(r\theta_{i})\widehat{R}^{(r)}_{s-}}=1+r\theta_{i}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-2^{-1}(r\theta_{i})^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]$
+$\qquad+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}),$ (4.41)
 $e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}$
-$\hskip 34.44434pt=r(\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}+2^{-1}r^{2}(\lambda_{2}^{3}\sigma_{e,2}^{2}r\theta_{2}^{2}-\lambda_{1}^{3}\sigma_{e,1}^{2}r\theta_{1}^{2})(\widehat{R}^{(r)}_{e})^{2}+o(r^{2}).$ (4.40)
-$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}}$
-$\hskip 34.44434pt=\mu r(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}-2^{-1}r^{2}\mu^{3}\sigma_{s}^{2}(\theta_{2}^{2}-\theta_{1}^{2})(\widehat{R}^{(r)}_{s})^{2})+o(r^{2}),$ (4.41)
-These lemmas are proved in Appendix A.4. From BAR (4.3) in Lemma 4.3, we aim to determine the limiting measures $\widetilde{\nu}_{1}$ and $\widetilde{\nu}_{2}$. This will be done step by step.
-(1st Step)   We expand (4.3) using Lemmas 4.7 and 4.8.
-For this, we need the asymptotic behaviors concerning $\widehat{R}^{(r)}_{e}$ and $\widehat{R}^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{s}$ and $\mathbb{P}^{(r)}_{e}$, respectively.
-To see them, we first put $\theta_{2}=\theta_{1}=\theta$ in (4.3), then
-$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o(r^{2})\big{)}\psi^{(r)}_{i}(\theta)$
-$\quad=c_{1}\left(\mu r\theta+\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o(r^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta),$ (4.42)
-where applying (4.35), (4.8) and Corollary 4.3 to (4.2) yields
-$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+\theta r+o(r))\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r(\lambda_{2}-\lambda_{1})\theta(1+\theta r)\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r^{2})$
-$\quad=\alpha^{(r)}_{e}e^{\theta\ell_{0}}r(\lambda_{2}-\lambda_{1})\theta\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r^{2}).$
-Hence, dividing both sides of (4.3) by $r$ and letting $r\downarrow 0$, we have
-$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
-Since $e^{r\theta\ell^{(r)}_{0}}=e^{(\ell_{0}+o(1))\theta}$ can be arbitrarily small by choosing small $\theta<0$, we have
-$\lim_{r\downarrow 0}\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=0.$
-and therefore, by Corollary 4.2, we have
-$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}=0]=0.$ (4.43)
-We next put $\theta_{1}=0$ and $\theta_{2}=\theta$ in (4.3). In this case,
-$\frac{1}{2}c_{2}\sigma_{2}^{2}\big{(}(\beta_{2}-\theta)\mu\theta r^{2}+o(r^{2})\big{)}\psi^{(r)}_{2}(\theta)+D^{(r)}(0,\theta)=0,$
-and, from (4.2),
+$\quad=r(\theta_{2}\lambda^{(r)}_{2}-\theta_{1}\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}+2^{-1}r^{2}[\theta_{2}^{2}(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}\lambda^{(r)}_{2}-\theta_{1}^{2}(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}\lambda^{(r)}_{1}]\widehat{R}^{(r)}_{e}$
+$\qquad-2^{-1}r^{2}[\theta_{2}^{2}(\lambda^{(r)}_{2})^{2}-\theta_{1}^{2}(\lambda^{(r)}_{1})^{2}](\widehat{R}^{(r)}_{e})^{2}+S_{3}(\widehat{R}^{(r)}_{e})o((r\theta)^{2}),$ (4.42)
+$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}$
+$\quad=r(\theta_{2}-\theta_{1})(\mu^{(r)}\widehat{R}^{(r)}_{s-})-2^{-1}r^{2}(\theta_{2}^{2}-\theta_{1}^{2})[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]$
+$\qquad+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}).$ (4.43)
+These lemmas are proved in Appendix A.4.
+We first compute $D^{(r)}(\theta,0)$.
+$D^{(r)}(\theta,0)$ $=\alpha^{(r)}_{e}(1-e^{r\theta\ell^{(r)}_{0}})\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
+$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(1-e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\qquad+\alpha^{(r)}_{e}(e^{r\theta(\ell^{(r)}_{0}+1)}-1)\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
+$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-1)1(L^{(r)}(0)=\ell^{(r)}_{0})]$
+$=-\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.44)
+Define ${\mathcal{E}}^{(r)}_{1}$ is defined as
+${\mathcal{E}}^{(r)}_{1}$ $=2^{-1}\mathbb{E}^{(r)}_{e}[\{(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\quad-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.45)
+Applying (4.8) and
+$e^{r\theta}e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}=(1+r\theta+o(r\theta))\{1-r\theta(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})$
+$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}]+o((r\theta)^{2})\}$
+$\quad=1-r\theta(-1+\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-2^{-1}(r\theta)^{2}[((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}]$
+$\qquad+S_{3}(\widehat{R}^{(r)}_{e})o((r\theta)^{2}),$
+to (4.3) and recalling the definition of $\Delta^{(r)}_{1}$, we have
+$D^{(r)}(\theta,0)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(-r\theta\Delta^{(r)}_{1}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{1})+o((r\theta)^{2}),$ (4.46)
+because $\mathbb{E}^{(r)}_{s}[S_{3}(\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
+Similarly, define ${\mathcal{E}}^{(r)}_{2}$ as
+${\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{e}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.47)
+Then, $D^{(r)}(0,\theta)$ is computed from (4.2) as
 $D^{(r)}(0,\theta)=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}-1)\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
+$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
 $\quad+\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(1-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])$
-$\quad+\alpha^{(r)}_{e}(\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]-e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})])$
-$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})])$
-$=\alpha^{(r)}_{e}r\theta e^{r\theta\ell^{(r)}_{0}}\big{(}\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda_{2}(1+r\theta)\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{)}+O(r^{2}),$
-where the 3rd equation is obtained by (4.3), and the last one is by Corollary 4.3 and Lemma 4.8. Since $D^{(r)}(0,\theta)=O(r^{2})$ from (4.3), we have
-$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\quad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
-because $\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are nonnegative.
-These results lead to the following asymptotic results, which strengthens Corollary 4.3.
+$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}$
+$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[1+r\theta\mu^{(r)}\widehat{R}^{(r)}_{s-}-2^{-1}(r\theta)^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$\quad-(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}[1-r\theta(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})$
+$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}]1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}+o((r\theta)^{2})$
+$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r\theta\Delta^{(r)}_{2}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{2})+o((r\theta)^{2}).$ (4.48)
+(Step 3)   In this step, we derive asymptotic expansions for $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$. First, note that (4.35) and (4.36) can be written as
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\varphi^{(r)}_{1}(0)+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.49)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (4.50)
+Then, substituting (4.46) into the asymptotic BAR (4.3) with $\theta_{2}=0$ and applying (4.49), we have
+$\frac{1}{2}rc_{1}\mu\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-r\theta_{1}{\mathcal{E}}^{(r)}_{1})+\theta_{1}o(r),$
+$\quad=c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0]+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+\theta_{1}o(r),$ (4.51)
+Similarly, substituting (4.3) into (4.3) with $\theta_{1}=0$ and applying (4.50), we have
+$\frac{1}{2}rc_{2}\mu\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\big{(}\Delta^{(r)}_{2}+r\theta_{2}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r)$
+$\quad=re^{\theta_{2}\ell_{0}}\big{(}\mu b_{2}\varphi^{(r)}_{2}(0)+\theta_{2}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r).$ (4.52)
+We next compute $\mathbb{P}[L^{(r)}=0]$ for (4.3).
 Lemma 4.9.
-$\lim_{r\downarrow 0}\mathbb{P}^{(r)}_{e}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,\quad\lim_{r\downarrow 0}\mathbb{P}^{(r)}_{s}[L^{(r)}(0-)=\ell^{(r)}_{0}]=0,$ (4.45)
-$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,\quad\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=0.$ (4.46)
+$\mathbb{P}[L^{(r)}=0]=\begin{cases}\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b=0,\\ \frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),&b\not=0.\end{cases}$ (4.53)
 Proof.
-The first equation of (4.45) is immediate from (4.2) and Lemma 4.6 together with (4.43) and (4.44), and the second equation follows from the first one by Lemma 4.1.
-As for (4.46), we only prove its first equation because its second one is similarly proved.
-By the Schwartz’s inequality,
-$\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}=\big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3/2}(\widehat{R}^{(r)}_{s})^{1/2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\big{)}^{2}$
-$\leq\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})].$ (4.47)
-From (4.5) for $k=3$ and Lemma 4.4, we have
-$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{3}1(L^{(r)}(0-)=\ell^{(r)}_{0})]\leq\mathbb{E}[(T^{(r)}_{s})^{3}],$
-which is uniformly bounded for $r\in(0,1]$, and therefore (4.3) and (4.44) prove (4.46). ∎
-(2nd Step)   In this step, we first expand the exponential terms in $D^{(r)}(\theta_{1},\theta_{2})$ of (4.2), using Lemmas 4.7 and 4.8 in the step 1, then
-$r^{-1}D^{(r)}(\theta_{1},\theta_{2})$
-$=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}-e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(r^{-1}+\mu\theta_{2}\widehat{R}^{(r)}_{s}-2^{-1}r\mu^{3}\sigma_{s}^{2}\theta_{2}^{2}(\widehat{R}^{(r)}_{s})^{2})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(\mu(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}-2^{-1}r\mu^{3}\sigma_{s}^{2}(\theta_{2}^{2}-\theta_{1}^{2})(\widehat{R}^{(r)}_{s})^{2})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r))$
-×$\hskip 51.6665pt\times\mathbb{E}^{(r)}_{s}[(r^{-1}-\lambda_{1}\theta_{1}\widehat{R}^{(r)}_{e}-2^{-1}r\lambda_{1}^{3}\sigma_{e,1}^{2}\theta_{1}^{2}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}(1+\theta_{2}r)+o(r))\mathbb{E}^{(r)}_{s}[((\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}$
-$\hskip 51.6665pt+2^{-1}r(\lambda_{2}^{3}\sigma_{e,2}^{2}r\theta_{2}^{2}-\lambda_{1}^{3}\sigma_{e,1}^{2}r\theta_{1}^{2})(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})]+o(r).$ (4.48)
-We next apply Lemma 4.9 to this formula, then
-$r^{-1}D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}-e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[(r^{-1}+\mu\theta_{2}\widehat{R}^{(r)}_{s})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}+o(r))\mathbb{E}^{(r)}_{e}[\mu(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r))\mathbb{E}^{(r)}_{s}[(r^{-1}-\lambda_{1}\theta_{1}\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}(1+\theta_{2}r)+o(r))\mathbb{E}^{(r)}_{s}[(\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r)$
-$=\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}\theta_{2}-e^{\theta_{1}\ell_{0}}\theta_{1})(\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}])$
-$\quad+\alpha^{(r)}_{e}(e^{\theta_{2}\ell_{0}}\lambda_{2}\theta_{2}-e^{\theta_{1}\ell_{0}}\lambda_{1}\theta_{1})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+o(r).$ (4.49)
-Recall that
-$\Delta^{(r)}_{i}=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\hskip 43.05542pt-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}],\qquad i=1,2,$
-then (4.3) can be written as
-$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}r\big{(}e^{\theta_{2}\ell_{0}}\theta_{2}\Delta^{(r)}_{2}-e^{\theta_{1}\ell_{0}}\theta_{1}\Delta^{(r)}_{1}\big{)}+o(r^{2}).$ (4.50)
-Substituting this into BAR (4.3) and dividing by $r$, we have
-$\frac{1}{2}\sum_{i=1,2}r\mu c_{i}\sigma_{i}^{2}\big{(}\beta_{i}-\theta_{i}\big{)}\theta_{i}\psi^{(r)}_{i}(\theta_{i})$
-$\quad=c_{1}\mu\theta_{1}\mathbb{P}[L^{(r)}=0]+\alpha^{(r)}_{e}\big{(}e^{\theta_{2}\ell_{0}}\theta_{2}\Delta^{(r)}_{2}-e^{\theta_{1}\ell_{0}}\theta_{1}\Delta^{(r)}_{1}\big{)}+o(r).$ (4.51)
-Substituting $\theta_{2}=0$ or $\theta_{1}=0$, this asymptotic BAR is split into two equations:
-$\frac{1}{2}r\mu c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-\alpha^{(r)}_{e}e^{\theta_{1}\ell_{0}}\Delta^{(r)}_{1}+o(r),$ (4.52)
-$\frac{1}{2}r\mu c_{2}\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\Delta^{(r)}_{2}+o(r).$ (4.53)
-(3rd Step)   We compute $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$ from (4.52) and (4.53) separately. First consider (4.53). Note that (4.34) yields
-$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (4.54)
-Hence, from (4.53), we have
-$\frac{1}{2}r\mu c_{2}\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})$ $=\frac{1}{2}e^{\theta_{2}\ell_{0}}r\mu c_{2}\sigma_{2}^{2}\beta_{2}\varphi^{(r)}_{2}(0)+o(r),$
-which is simply written as
-$\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})$ $=e^{\theta_{2}\ell_{0}}\beta_{2}\varphi^{(r)}_{2}(0)+o(1).$ (4.55)
-We next consider (4.52). From (4.33) of Lemma 4.6,
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-r\mu b_{1}\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r).$ (4.56)
-Hence, from (4.52), we have
-$r\mu c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})$
-$\qquad=2c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0)]+2r\mu b_{1}e^{\theta_{1}\ell_{0}}\varphi^{(r)}_{1}(0)+o(r).$ (4.57)
-This yields
-$\mathbb{P}[L^{(r)}=0)]=r\frac{c_{1}\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})-2b_{1}e^{\theta_{1}\ell_{0}}\varphi^{(r)}_{1}(0)}{2c_{1}(1-e^{\theta_{1}\ell_{0}})}+o(r).$ (4.58)
-We separately consider (4.58) for $b_{1}=0$ and $b_{1}\not=0$.
-(Case of $b_{1}=0$)   In this case, $\beta_{1}=b_{1}=0$, and therefore (4.58) yields
-$\mathbb{P}[L^{(r)}=0)=r\lim_{\theta_{1}\to 0}\frac{r\sigma_{1}^{2}\theta_{1}\psi^{(r)}_{1}(\theta_{1})}{2(e^{\theta_{1}\ell_{0}}-1)}+o(r)=\frac{r\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)+o(r).$ (4.59)
-Hence, for $b_{1}=0$, it follows from (4.56) and (4.59) that
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)+o(r),$ (4.60)
-$\psi^{(r)}_{1}(\theta_{1})=\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\varphi^{(r)}_{1}(0)+o(r).$ (4.61)
-(Case of $b_{1}\not=0$)   By substituting $\theta_{1}=\beta_{1}$ into (4.58), we have
-$\mathbb{P}[L^{(r)}=0)]=\frac{rb_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\varphi^{(r)}_{1}(0)+o(r)$ (4.62)
-and therefore, it follows from (4.56), (4.3) and (4.62) that
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\frac{r\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)+o(r),$ (4.63)
-$\psi^{(r)}_{1}(\theta_{1})=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\varphi^{(r)}_{1}(0)+o(r).$ (4.64)
-(4th Step)   We now complete the proof of Theorem 3.1. We first prove (i). Note that the bounded away assumption can be written as
-$\liminf_{r\downarrow 0}\varphi^{(r)}_{i}(0)>0,\qquad i=1,2.$ (4.65)
-Under this condition, for $\theta_{1}\in\mathbb{R}_{+}$ and $\theta_{2}\leq 0$,
+From (4.3), we have
+$\mathbb{P}[L^{(r)}=0]$ $=r\frac{c_{1}\mu\sigma_{1}^{2}\big{(}\theta_{1}-\beta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}(b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r).$ (4.54)
+If $b_{1}=0$, then $\beta_{1}=b_{1}=0$, and therefore (4.54) yields
+$\mathbb{P}[L^{(r)}=0]$ $=r\lim_{\theta_{1}\to 0}\frac{\theta_{1}(c_{1}\mu\sigma_{1}^{2}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r)$
+$=\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),$
+which proves (4.53) for $b_{1}=0$. If $b_{1}\not=0$, then substituting $\theta_{1}=\beta_{1}$ into (4.54) yields
+$\mathbb{P}[L^{(r)}=0]=\frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),$
+which is (4.53) for $b_{1}\not=0$. ∎
+We also compute $\alpha^{(r)}_{e}\Delta^{(r)}_{1}$ to be used later.
+Lemma 4.10.
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}\Big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\frac{1}{c_{1}\mu}\beta_{1}e^{\beta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\Big{)}+o(r),&b\not=0.\end{cases}$ (4.55)
+This lemma is immediate from substituting $\mathbb{P}^{(r)}_{e}[L^{(r)}=0]$ of Lemma 4.9 into (4.49).
+We are now ready to compute $\psi^{(r)}_{i}(\theta_{i})$ for $i~{}1,2$, which will be a key to prove Theorem 3.1.
+Lemma 4.11.
+$\psi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}\not=0,\end{cases}$ (4.56)
+$\psi^{(r)}_{2}(\theta_{2})$ $=\frac{\beta_{2}e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\varphi^{(r)}_{2}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}e^{\theta_{2}\ell_{0}}+o(1).$ (4.57)
+Proof.
+We first compute $\psi^{(r)}_{1}(\theta_{1})$. Assume that $b_{1}=0$. Substituting $\mathbb{P}^{(r)}_{e}[L^{(r)}=0]$ of Lemma 4.9 into (4.3) for $\beta_{1}=b_{1}=0$, we have
+$-rc_{1}\mu\sigma_{1}^{2}\theta_{1}\psi^{(r)}(\theta_{1})$
+$\quad=(1-e^{\theta_{1}\ell_{0}})\frac{r}{\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+r2e^{\theta_{1}\ell_{0}}\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r),$
+which proves (4.56) for $b_{1}=0$. Next assume that $b_{1}\not=0$. Similarly to the case of $b_{1}=0$, it follows from Lemma 4.9 and (4.3) for $b_{1}\not=0$ that
+$\frac{r}{2}c_{1}\mu\sigma_{1}^{2}(\beta_{1}-\theta_{1})\psi^{(r)}(\theta_{1})$
+$\quad=\frac{re^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})}{(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r)$
+$\quad=\frac{rb_{1}\mu\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\big{(}e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+r\frac{\beta_{1}e^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})+\theta_{1}e^{\theta_{1}\ell_{0}}(e^{\beta_{1}\ell_{0}}-1)}{(e^{\beta_{1}\ell_{0}}-1)}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r).$
+Hence, we have
+$\psi^{(r)}_{1}(\theta_{1})$ $=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}-\theta_{1}e^{\theta_{1}\ell_{0}}+(\theta_{1}-\beta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
+$=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}$
+$\quad+\frac{\beta_{1}(e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}-(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
+$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2e^{\theta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
+Thus, we have (4.56) for $b_{1}\not=0$. We next compute $\psi^{(r)}_{2}(\theta_{2})$. From (4.3), we have
+$\psi^{(r)}_{2}(\theta_{2})=\frac{e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\beta_{2}\varphi^{(r)}_{2}(0)+(\beta_{2}-(\beta_{2}-\theta_{2}))\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}+\theta_{2}o(r),$
+which proves (4.57). ∎
+5 Proof of Theorem 3.1
+From Lemma 4.11, we can see that, if
+$\lim_{r\downarrow 0}{\mathcal{E}}^{(r)}_{1}=\lim_{r\downarrow 0}{\mathcal{E}}^{(r)}_{2}=0,$ (5.1)
+then the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{1}(\theta_{2})$ can be obtained as long as the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{1}(0)$ exist.
+This suggests that (5.1) must be a key condition for Theorem 3.1.
+Taking this into account, we separate the proof of Theorem 3.1 in two steps by two propositions.
+Proposition 5.1.
+In Theorem 3.1, the assumption (3.f) can be replaced by (5.1).
+Proposition 5.2.
+Under the assumptions (3.a)–(3.e), (5.1) holds if
+$\lim_{r\downarrow 0}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})=0.$ (5.2)
+Furthermore, (5.2) holds if (3.f) holds. Hence, (3.f) implies (5.1).
+In the next two subsections, we prove these propositions. For their proofs, we will use the following lemmas, which are proved in Appendix A.5, respectively.
+Lemma 5.1.
+Under the assumptions (3.a)–(3.e), if
+$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,$ (5.3)
+then
+$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.4)
+Lemma 5.2.
+Under the assumptions (3.a)–(3.e), if (5.4) holds, then
+$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.5)
+Remark 5.1.
+From the proof of Proposition 5.2, we can see that (5.4) and (5.5) imply (5.2), and therefore (5.3) is sufficient for (5.2) and (5.1).
+We think that (5.3) is likely true because $L^{(r)}(\cdot)$ would weakly converge to a reflected diffusion and the stationary distribution of the reflected diffusion has a density.
+However, the verification of (5.3) would need a process limit for $L^{(r)}(\cdot)$, so this verification is beyond the present framework.
+5.1 Proof of Proposition 5.1
+Throughout this proof, the part (i), (ii) and (iii) of Theorem 3.1 under the condition (5.1) instead of (3.f) are denoted by the part (i)’, (ii)’ and (iii)’, respectively.
+We first prove the part (i)’. Since the condition in the part (i)’ is given by (3.6), Corollary 4.3 implies that
 $\lim_{r\downarrow 0}|\psi^{(r)}_{i}(\theta_{i})/\varphi^{(r)}_{i}(0)-\varphi^{(r)}_{i|cd}(\theta_{i})|=0.$ (5.6)
-is equivalent to
-$\lim_{r\downarrow 0}|(\psi^{(r)}_{i}(\theta_{1})-\varphi^{(r)}_{i}(\theta_{1}))|=0,$ (4.67)
-since $\varphi^{(r)}_{i|cd}(\theta_{i})=\varphi^{(r)}_{i}(\theta_{i})/\varphi^{(r)}_{i}(0)$. Here, (4.67) holds because (4.13) implies that
-$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{0})]\right|$
-$\leq e^{|\theta_{1}|r\ell^{(r)}_{0}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
-Hence, (4.66) holds. Applying it to (4.61) and (4.64), we obtain
+By this formula, we can immediately see that Lemma 4.11 and (5.1) yield
 $\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (5.7)
-and, similarly, (4.55) yields
 $\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{0}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (5.8)
 Since $\widetilde{\varphi}_{i|cd}$ is the moment generating function of $\widetilde{\nu}_{i|cd}$, (5.7) and (5.8) are equivalent to that $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$ for $i=1,2$, respectively.
-Thus, (i) is proved.
-We next prove (ii).
-For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumption in (ii).
+Thus, the part (i)’ is proved.
+We next prove the part (ii)’.
+For this, we compute the limits of $\varphi^{(r)}_{i}(0)$ for $i=1,2$ under the assumptions in the part (ii)’.
 First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$.
 In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
-$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
+$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
 Applying this fact to the definitions of $\Delta^{(r)}_{i}$ yields
 $\lambda^{(r)}_{2}\Delta^{(r)}_{1}-\lambda^{(r)}_{1}\Delta^{(r)}_{2}$
-$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (4.70)
-Thus, if $b_{1}=0$, then it follows from (4.54) and (4.60) that
-$\lambda_{2}\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
-Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$, this can be written as
-$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$
+$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (5.9)
+Compute $\alpha^{(r)}_{e}\Delta^{(r)}_{1}$ by Lemma 4.10 and (5.1), and recall (4.50), then
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}b_{1}\mu\varphi^{(r)}_{1}(0)+o(r),&b\not=0,\\ \end{cases}$ (5.10)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (5.11)
+Then, for $b_{1}=0$, (5.1) implies that
+$\lambda_{2}\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
+Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$ for $i=1,2$, this can be written as
+$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$ (5.12)
 Hence,
 $\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}\sigma_{1}^{2}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.13)
-On the other hand, if $b_{1}\not=0$, then, from (4.54) and (4.63),
-$\frac{\lambda_{2}\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
-Hence,
+Similarly, for $b_{1}\not=0$, from (5.10) and (5.11), we have
+$\frac{\lambda_{2}b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$ (5.14)
+Hence, $\mu c_{i}=\lambda_{i}$ for $i=1,2$ yields
 $\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.15)
-Thus, (4.65) is satisfied for $i=1,2$.
-We next consider the case that $\lambda_{1}=\lambda_{2}$, equivalently, $c_{1}=c_{2}$. Then, from the definitions of $\Delta^{(r)}_{i}$, we have
-$\Delta^{(r)}_{1}-\Delta^{(r)}_{2}=(\lambda^{(r)}_{1}-\lambda^{(r)}_{2})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=o(r),$ (4.73)
-because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). If $b_{1}=0$, then (4.54) and (4.60) yield
-$\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-\Delta^{(r)}_{2})=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r),$
-Hence, by (4.3) and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$,
-$\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\mu b_{2}(1-\varphi^{(r)}_{1}(0))+o(1)$
-Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54), (4.63) and (4.3) yield
-$\frac{\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
-Again by $c_{1}=c_{2}$ and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$, this proves (4.72).
-Let
-$\widetilde{\varphi}_{i}(\theta)=\widetilde{\varphi}_{i|cd}(\theta)\widetilde{\varphi}_{i}(0),\qquad i=1,2,$
-then $\widetilde{\varphi}_{i}$ is the moment generating function of $\widetilde{\nu}_{i}$, so (4.68) and (4.69) together with (4.71) and (4.71) prove that $\nu^{(r)}_{i}$ weakly converges to $\widetilde{\nu}_{i}$ for $i=1,2$.
-Thus, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}$ of (3.8).
-Similarly, by the assumption (3.c), (4.5), (4.59) and (4.62), we have
-$\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\lim_{r\downarrow 0}(c^{(r)}_{1}\mu^{(r)}(\varphi^{(r)}_{1}(0)-\mathbb{P}[L^{(r)}=0])+c^{(r)}_{2}\mu^{(r)}\varphi^{(r)}_{2}(0))=\lambda_{1}\widetilde{\varphi}_{1}(0)+\lambda_{2}\widetilde{\varphi}_{2}(0).$
-It remains to prove (iii). Assume that $\lambda_{1}>\lambda_{2}$, then $\Delta^{(r)}_{1}>\Delta^{(r)}_{2}$ by (4.73). For $b=0$, this yields
-$\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
+We next consider the case that $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$. Define $\overline{\varphi}_{1}(0)$ and $\underline{\varphi}_{1}(0)$ as
+$\overline{\varphi}_{1}(0)=\limsup_{r\downarrow 0}\varphi^{(r)}_{1}(0),\qquad\underline{\varphi}_{1}(0)=\liminf_{r\downarrow 0}\varphi^{(r)}_{1}(0),$
+and let $\{\overline{r}_{n}\in(0,1];n\geq 1\}$ and $\{\underline{r}_{n}\in(0,1];n\geq 1\}$ be the sequences such that $\varphi^{(\overline{r}_{n})}_{1}(0)$ and $\varphi^{(\underline{r}_{n})}_{1}(0)$ attain $\overline{\varphi}_{1}(0)$ and $\underline{\varphi}_{1}(0)$ as $n\to\infty$, respectively.
+We aim to show $\overline{\varphi}_{1}(0)=\underline{\varphi}_{1}(0)$ using $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$.
+To see this, we first consider $\overline{\varphi}_{1}(0)$. Since $\varphi^{(\overline{r}_{n})}_{1}(0)$ converges to $\overline{\varphi}_{1}(0)$, applying (5.1) and (4.17) to Lemma 4.11, we have
+$\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(\theta_{1})=\widetilde{\varphi}_{1}(\theta_{1})\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(\theta_{1})\overline{\varphi}_{1}(0).$ (5.16)
+Since the distribution $\widetilde{\nu}_{1}$ has the bounded density $h_{1|cd}$ on $[0,\ell_{0}]$, this implies that
+$\lim_{n\to\infty}\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]\leq\lim_{n\to\infty}\mathbb{P}[\overline{r}_{n}L^{(\overline{r}_{n})}\in[\ell_{0}-\varepsilon,\overline{r}_{n}\ell^{(\overline{r}_{n})}_{0}]]$
+[…]
+because $0\leq\overline{\varphi}_{1}(0)\leq 1$. Hence, $\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]=o(1)$ as $n\to\infty$. By Lemma 5.1, this implies that
+$\lim_{n\to\infty}\mathbb{E}^{(\overline{r}_{n})}_{s}[\widehat{R}^{(\overline{r}_{n})}_{e}1(L^{(\overline{r}_{n})}(0)=\ell^{(\overline{r}_{n})}_{0})]=0.$ (5.18)
+Then, from the definitions of $\Delta^{(r)}_{i}$ and $\lambda^{(\overline{r}_{n})}_{1}-\lambda^{(\overline{r}_{n})}_{2}=O(\overline{r}_{n})$, we have
+$\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2}=(\lambda^{(\overline{r}_{n})}_{1}-\lambda^{(\overline{r}_{n})}_{2})\mathbb{E}^{(\overline{r}_{n})}_{s}[\widehat{R}^{(\overline{r}_{n})}_{e}1(L^{(\overline{r}_{n})}(0)=\ell^{(\overline{r}_{n})}_{0})]=o(\overline{r}_{n}),$ (5.19)
+If $b_{1}=0$, then (4.50) and Lemma 4.10 yield
+$\alpha^{(\overline{r}_{n})}_{e}(\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2})=\frac{rc_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)-r\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)+o(\overline{r}_{n}),$
+Hence, by (5.19) and $\varphi^{(\overline{r}_{n})}_{1}(0)+\varphi^{(\overline{r}_{n})}_{2}(0)=1$,
+$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)=\mu b_{2}(1-\varphi^{(\overline{r}_{n})}_{1}(0))+o(1).$
+This is identical with (5.12) for $r=\overline{r}_{n}$ because $c_{1}=c_{2}$ by $\lambda_{1}=\lambda_{2}$.
+Hence, we have (5.13) for $r=\overline{r}_{n}$ and $b_{1}=0$.
+Similarly, if $b_{1}\not=0$, then, again by (4.50), (5.11) and (5.19) for $r=\overline{r}_{n}$, we have
+$\frac{b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(\overline{r}_{n})}_{1}(0)-\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)=o(1).$
+This is identical with (5.14) for $r=\overline{r}_{n}$ by $\lambda_{1}=\lambda_{2}$. Hence, we have (5.15) for $r=\overline{r}_{n}$ and $b_{1}\not=0$. Thus, we have
+$\overline{\varphi}_{1}(0)=\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(0)>0.$
+We can repeat the exactly same arguments for $\{\underline{r}_{n};n\geq 0\}$, and therefore we also have
+$\underline{\varphi}_{1}(0)=\lim_{n\to\infty}\varphi^{(\underline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(0)>0.$
+Hence, $\overline{\varphi}_{1}(0)=\underline{\varphi}_{1}(0)$ as claimed, and therefore $\lim_{r\downarrow 0}\varphi^{(r)}_{1}(0)=\widetilde{\varphi}_{1}(0)>0$.
+Hence, the condition (3.6) in the part (i)’ is satisfied, so we can use the results in that part.
+Namely, we have (5.7) and (5.8), equivalently, (3.7).
+From this together with (5.13) and (5.15), we have (3.11).
+It remains to prove (3.12) in the part (ii)’. To see this, apply the assumption (3.c), (5.13), (5.15) and Lemma 4.9 to (4.5), then we have
+$\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\lim_{r\downarrow 0}(c^{(r)}_{1}\mu^{(r)}(\varphi^{(r)}_{1}(0)-\mathbb{P}[L^{(r)}=0])+c^{(r)}_{2}\mu^{(r)}\varphi^{(r)}_{2}(0))=\lambda_{1}\widetilde{\varphi}_{1}(0)+\lambda_{2}\widetilde{\varphi}_{2}(0),$
+which proves (3.12).
+Finally, we prove the part (iii)’. Assume that $\lambda_{1}>\lambda_{2}$, then $\Delta^{(r)}_{1}>\Delta^{(r)}_{2}$ by (5.19). For $b=0$, this yields
+$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
 and therefore $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$.
 Similarly, for $b\not=0$, we also have $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$.
-Hence, (4.65) holds for $i=1$, so $\nu^{(r)}_{1|cd}$ weakly converges to $\widetilde{\nu}_{1|cd}$ as $r\downarrow 0$ by (i).
+Hence, (3.6) holds for $i=1$, so $\nu^{(r)}_{1|cd}$ weakly converges to $\widetilde{\nu}_{1|cd}$ as $r\downarrow 0$ by the part (i)’.
 Similarly, if $\lambda_{1}<\lambda_{2}$, then (3.6) holds for $i=2$, so $\nu^{(r)}_{2|cd}$ weakly converges to $\widetilde{\nu}_{2|cd}$.
+Thus, the proof of Proposition 5.1 is completed.
+5.2 Proof of Proposition 5.2
+Using (5.2), we prove (5.1) by reductio ad absurdum.
+First, assume that ${\mathcal{E}}^{(r)}_{1}$ does not vanish as $r\downarrow 0$.
+Then, there are a sequence $\{r_{n}\in(0,1];n\geq 0\}$ and a constant $\delta_{1}\not=0$ such that $r_{n}\downarrow 0$ and ${\mathcal{E}}^{(r_{n})}_{1}\to\delta_{1}$ as $n\to\infty$, because ${\mathcal{E}}^{(r)}_{1}$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
+We next choose a subsequence $\{r^{\prime}_{n}\in(0,1];n\geq 0\}\subset\{r_{n}\in(0,1];n\geq 0\}$ such that $\alpha^{(r)}_{e}$ and $\varphi^{(r^{\prime}_{n})}_{1}(0)$ converge to constants $a_{e}>0$ and $d_{1}\in[0,1]$, respectively, as $n\to\infty$, which are possible because $\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and above by Corollary 4.1 and $\varphi^{(r^{\prime}_{n})}_{1}(0)\in[0,1]$.
+Then, by Lemma 4.11, we have
+$\lim_{n\to\infty}\varphi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}\not=0.\end{cases}$
+This shows that $\nu^{(r)}_{1}$ weakly converges to a finite measure on $[0,\ell_{0}]$, but this measure cannot have a negative mass at $\ell_{0}$, so we must have that $\delta_{1}<0$ because $\delta_{1}\not=0$.
+We next consider the sequence $\{{\mathcal{E}}^{(r^{\prime}_{n})}_{2};n\geq 1\}$.
+If this sequence does not vanish as $n\to\infty$, then there is a subsequence $\{r^{\prime\prime}_{n}\in(0,1];n\geq 0\}\subset\{r^{\prime}_{n}\in(0,1];n\geq 0\}$ such that ${\mathcal{E}}^{(r^{\prime\prime}_{n})}_{2}$ converges some $\delta_{2}\not=0$ as $n\to\infty$.
+Then, by the exactly same argument as used for $\{{\mathcal{E}}^{(r_{n})}_{1};n\geq 1\}$ using (4.57) instead of (4.56), we must have that $\delta_{2}<0$.
+Hence, we have
+$\lim_{r^{\prime\prime}_{n}\downarrow 0}({\mathcal{E}}^{(r^{\prime\prime}_{n})}_{1}+{\mathcal{E}}^{(r^{\prime\prime}_{n})}_{2})=\delta_{1}+\delta_{2}<0.$
+This contradicts (5.2), and therefore ${\mathcal{E}}^{(r)}_{1}$ must vanish as $r\downarrow 0$.
+By (5.2), this further implies that ${\mathcal{E}}^{(r)}_{2}$ vanishes as $r\downarrow 0$.
+Hence, we have (5.1).
+We now prove that (3.f) implies (5.2). By the definitions of ${\mathcal{E}}^{(r)}_{1}$ and ${\mathcal{E}}^{(r)}_{2}$, we have
+${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
+$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.20)
+We separately consider the two cases that $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\not=\lambda_{2}$.
+First, consider the case that $\lambda_{1}=\lambda_{2}$.
+In this case, $\sigma_{e,1}^{2}=\sigma_{e,2}^{2}$ by (3.f).
+Hence, (5.2) implies (5.2).
+Next consider the case that $\lambda_{1}\not=\lambda_{2}$. In this case, we consider the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$. From (4.2), we have
+$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.21)
+Applying (4.37), (4.8) and Corollary 4.4 to this equation, we have
+$D^{(r)}(\theta,\theta)$ $=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}\big{[}(r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}$
+$\quad+2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}\lambda^{(r)}_{2}-(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}\lambda^{(r)}_{1}\}\widehat{R}^{(r)}_{e}$
+$\quad-2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2})^{2}-(\lambda^{(r)}_{1})^{2}\}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+o((r\theta)^{2}),$
+$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+(r\theta)^{2}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})\big{\}}+o((r\theta)^{2}).$
+This implies that
+$D^{(r)}(\theta,\theta)$ $=r\theta\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+O((r\theta)^{2}).$ (5.22)
+Substitute this formula into the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$, which is given by
+$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta)$
+$\quad=c_{1}\left(\mu r\theta-\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o((\theta r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta).$ (5.23)
+Then, picking up the terms of order $r\theta$, we have
+$c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r\theta).$
+Since $\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=O(r)$ by Corollary 4.2 and Lemma 4.9, this yields
+$(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r).$ (5.24)
+Since $\lambda_{1}\not=\lambda_{2}$ by (f.1), (5.24) implies (5.4), which further implies (5.5). Hence,
+$\lim_{r\downarrow 0}\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,$ (5.25)
+and therefore we have (5.2) by (5.2) and Lemma 5.2. Thus, the proof of Proposition 5.2 is completed.
 6 Concluding remarks
@@
 Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic.
-From the assumptions (3.a)–(3.e), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
+From the assumptions (3.a)–(3.f), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
 Recall that, if one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
-$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
+$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
 $\qquad+\beta_{2}e^{\theta_{2}\ell_{0}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (6.1)
-which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
-For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$.
+which can be obtained from (4.3) and (4.2) similarly to $D^{(r)}(\theta,\theta)$ from (5.22) and (5.2), recall that $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
+The BAR for $b_{1}=0$ is obtained from (6.1) by letting $b_{1}\downarrow 0$.
 So, we only consider the case that $b_{1}\not=0$.
@@
 where $W(\cdot)$ is the standard Brownian motion, $Y(\cdot)\equiv\{Y(t);t\geq 0\}$ is a non-deceasing process satisfying that $\int_{0}^{t}1(Z(u)>0)Y(du)=0$ for $t\geq 0$.
-If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [6] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$.
-However, $Y(\cdot)$ may not be absolutely continuous.
-In the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered its stationary distribution using the Ito formula, which yields
-$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\in\mathbb{R}.$ (5.3)
+Recently, Miyazawa [11] (see also [1, 2, 13]) shows that (6.2) has a unique weak solution, and $Z(t)$ has a stationary distribution if only if $b_{2}>0$, which corresponds to $\rho^{(r)}_{2}<1$.
+Assume that $Z(\cdot)$ is stationary, then, applying the Ito formula to (6.2) for test function $f(x)=e^{\theta x}$, we have
+$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\leq 0.$ (6.3)
 where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$.
-However, we had not solved this stationary equation.
-So, it was conjectured that
+(6.3) is the stationary equation for $Z(\cdot)$.
+Then, we can see that (6.1) for $\theta_{1}=\theta_{2}=\theta$ agrees with (6.3) if and only if $c_{1}=c_{2}$, equivalently, $\lambda_{1}=\lambda_{2}$.
+Hence, if $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to $Z(\cdot)$ as $r\downarrow 0$, then we must have $\lambda_{1}=\lambda_{2}$.
+Of course, this does not exclude the possibility that $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to some process other than $Z(\cdot)$.
+We conjecture that this is indeed possible because the stationary distribution of $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to the $\widetilde{\nu}$ of (3.9) for $\lambda_{1}\not=\lambda_{2}$ at least when $T_{s}$ is exponentially distribution by (ii) of Theorem 3.1.
+Thus, we have two challenging problems.
   • (a)
-    there is a unique weak solution $Z(\cdot)$ for the stochastic integral equation (5.2),
+    Prove that $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to $Z(\cdot)$ when $\lambda_{1}=\lambda_{2}$.
   • (b)
-    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to the $Z(\cdot)$ of (a),
-  • (c)
-    the stationary distribution of $Z(t)$ uniquely exists under the stability condition $b_{2}>0$, and agrees with $\widetilde{\nu}$ of (3.8) in Theorem 3.1 under the one of conditions in (ii) of it in addition to $b_{2}>0$.
-During the review process of this paper, the author have known that (a) is resolved by Atar et al.
-[1, Lemma 4.3] (see also Miyazawa [11] for a different proof), while (c) can be well answered because the stationary distribution of the $Z(\cdot)$ is fully obtained in [11], in which it is claimed that the limiting distribution $\widetilde{\nu}$ of this paper is identical with the stationary distribution of the $Z(\cdot)$, but this claim is only true if and only if $c_{1}=c_{2}$.
-This fact also can be seen from the comparison of (5.3) with (5.1), which does not require the $\widetilde{\nu}$ itself.
-However, the $\widetilde{\nu}$ has not yet been completely specified.
-As for (a), the stationary distributions are obtained for more general reflected diffusion processes than $Z(\cdot)$.
-We refer to [11, 13] for them, and will not further discuss about the conjectures (a) and (c).
-As for the conjecture (b), it is still open.
+    Find a process limit of $L^{(r)}_{\rm dif}(\cdot)$ when $\lambda_{1}\not=\lambda_{2}$.
+Both problems require theoretical investigations based on process limits, which is beyond the present framework. So, we leave them for future study.
 6.2 Intuitive verification of (3.13)
@@
 A.1 Proofs of Corollary 3.1
-We only verify (3.1) for $b_{1}\not=0$ because it is similarly proved for $b_{1}=0$. First, from (4.62) and (3.10),
-$\frac{1}{r}\mathbb{P}[L^{(r)}=0)]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
+We only verify (3.1) for $b_{1}\not=0$ because it is similarly proved for $b_{1}=0$. Since ${\mathcal{E}}^{(r)}_{1}=o(1)$ by Proposition 5.2, it follows from (4.53) and (3.10) that
+$\frac{1}{r}\mathbb{P}[L^{(r)}=0]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
 $=\frac{b_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1).$
@@
 $\lambda_{2}p_{2,0}=c_{1}\mu p_{2,1},$ (A.4)
-$(\lambda_{2}+c_{1}\mu)p_{2,\ell}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{0}-1],$ (A.5)
+$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{0}-1],$ (A.5)
 $(\lambda_{2}+c_{1}\mu)p_{2,\ell_{0}}=c_{2}\mu p_{3,\ell_{0}+1},$ (A.6)
@@
 A.3 Proofs of Lemma 4.4
-We first prove (4.22) and (4.4). Apply $f({\boldsymbol{x}})=x_{2}^{k+1}1(x_{2}\in B)$ to (4.1), then
-$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})1(T^{(r)}_{e,1}\in B)]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
-$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})1(T^{(r)}_{e,2}\in B)]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
-where (4.1) is still applicable since the right-hand side of this equation is finite. Hence, (4.22) is obtained. (4.4) is similarly proved. We next prove (4.24).
+We first prove (4.26) and (4.27). Apply $f(z,{\boldsymbol{x}})=x_{1}^{k+1}$ to (4.1), then
+$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
+$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
+where (4.1) is applicable since the right-hand side of this equation is finite. Hence, (4.26) is obtained. Similarly, we have (4.27) applying $f(z,{\boldsymbol{x}})=x_{2}^{k+1}$ to (4.1) because
+$f^{\prime}(X^{(r)}(t))=-(k+1)(R^{(r)}_{s}(t))^{k}(1-1(L^{(r)}(t)=0)),$
+and $R^{(r)}(t)$ has the same distribution as $T^{(r)}_{s}$ when $L^{(r)}(t)=0$. We next prove (4.28). For this, we use the following inequality,
 $|(x\wedge 1/r)^{k}-x^{k}|$ $\;=\;x^{k}1(x>1/r)<rx^{k+1}1(x>1/r)$
 $<r^{2}x^{k+2}1(x>1/r),\qquad x\geq 0,k\geq 1.$ (A.33)
-Hence, the last inequality in (A.3) for $k=1$ and the uniform integrability condition yield
+Hence, the last inequality in (A.3) for $k=1$ and the uniform boundedness of $\mathbb{E}[(T^{(r)}_{u})^{3}$ yield
 $|\mathbb{E}[\widehat{T}^{(r)}_{u}]-\mathbb{E}[T^{(r)}_{u}]|\leq r^{2}\mathbb{E}[(T^{(r)}_{u})^{3}1(T^{(r)}_{u}>1/r)]=o(r^{2}).$ (A.34)
@@
 $|\mathbb{E}[(\widehat{T}^{(r)}_{u})^{2}]-\mathbb{E}[(T^{(r)}_{u})^{2}]|\leq r\mathbb{E}[(T^{(r)}_{u})^{2}1(T^{(r)}_{u}>1/r)]=o(r).$ (A.35)
-Thus, (A.34) and (A.35) prove (4.24) for $k=1,2$. Similarly, it follows from (4.22) with $B=(1/r,\infty)$ that, for $k=1,2$,
-$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|$ $\leq\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}>1/r)]$
-$=\frac{1}{k+1}\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}>1/r)]=o(r^{2-k}),$
-To prove (4.26), we apply $f({\boldsymbol{x}})=x_{2}^{k}1(x_{2}>1/r,x_{3}>0)$ to (4.1), then
-$k\mathbb{E}[(R^{(r)}_{e})^{k-1}1(R^{(r)}_{e}>1/r)]$ $=\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)]$
-$\quad+\sum_{i=1,2}\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,i})^{k}1(T^{(r)}_{e,i}>1/r)]\mathbb{P}^{(r)}_{e}[w_{i}(L^{(r)}(0-))]$
-$\geq\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)].$
-Hence, applying (4.22) with $B=(1/r,\infty)$ and $k-1$ instead of $k$ to the above formula, we have
-$|\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e})^{k}]|$ $\leq\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}1(R^{(r)}_{e}>1/r)]$
-$\leq(\alpha^{(r)}_{e})^{-1}k\mathbb{E}[(R^{(r)}_{e}(0))^{k-1}1(R^{(r)}_{e}>1/r)]$
-$\leq k\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k}1(T^{(r)}_{e,i}>1/r)],$
-where the last term is $o(r^{2})$ for $k=1$ and $o(r)$ for $k=2$ by (A.34) and (A.35). Thus, we have (4.26). (4.27) is similarly obtained.
+Thus, (4.28) is proved for $k=1,2$.
+We next prove (4.29) for $v=e$.
+Since $\mathbb{E}[(R^{(r)}_{e})^{k}]$ is uniformly bounded for $k=1,2$ by (4.27), we can prove (4.29) by the same argument as the proof of (4.28), but the order of $r$ is one less because of (4.27).
+We omit the proof for $v=s$ because it is similar to the case of $v=e$ except for (4.26) to be used instead of (4.27).
 A.4 Proofs of Lemmas 4.7 and 4.8
 Proof of Lemma 4.7.
-From Taylor expansion of twice differentiable function $f$:
-$|f(x)-(f(0)+f^{\prime}(0)x)|\leq\frac{1}{2}|x|^{2}\max_{|u|\leq|x|}|f^{\prime\prime}(u)|,\qquad x\in\mathbb{R},$ (A.36)
-we have $|e^{-x}-1+x|\leq\frac{1}{2}|x|^{2}e^{|x|}$. Hence, for $i=1,2$,
-$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}(\ell_{0}+o(r))}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}o(r)),$
-$e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}(\ell_{0}+r\theta_{i}+o(r))}=e^{r\theta_{i}\ell^{(r)}_{0}}(1+\theta_{i}r+\theta_{i}^{2}o(r)).$
-Thus, (4.35) is obtained. Using these asymptotic expansions, we have
-$e^{r\theta_{2}(\ell^{(r)}_{0}+1)}-e^{r\theta_{1}(\ell^{(r)}_{0}+1)}$
-$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+e^{r\theta^{(r)}_{2}\ell^{(r)}_{0}}(\theta_{2}r+\theta_{2}^{2}o(r))-e^{r\theta^{(r)}_{1}\ell^{(r)}_{0}}(\theta_{1}r+\theta_{1}^{2}o(r))$
-$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+r(e^{r\theta^{(r)}_{2}\ell^{(r)}_{0}}\theta_{2}-e^{r\theta^{(r)}_{1}\ell^{(r)}_{0}}\theta_{1})+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}-e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r)$
-$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}}+r(e^{\theta_{2}\ell_{0}}(1+\theta_{2}o(r))\theta_{2}-e^{\theta_{1}\ell_{0}}(1+\theta_{1}o(r))\theta_{1})$
-$\qquad+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}-e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r)$
-$\quad=e^{r\theta_{2}\ell_{0}}-e^{r\theta_{1}\ell_{0}}-r(e^{\theta_{2}\ell_{0}}\theta_{2}-e^{\theta_{1}\ell_{0}}\theta_{1})+(e^{r\theta_{2}\ell^{(r)}_{0}}\theta_{2}^{2}+e^{r\theta_{1}\ell^{(r)}_{0}}\theta_{1}^{2})o(r).$
-This proves (4.36), and (4.37) is similarly proved. ∎
+For a $(n+1)$-times continuously differentiable function $f$, we have, from Taylor expansion (e.g. see Theorem 19.9 of [14]),
+[…]
+$\leq\frac{1}{(n+1)!}|x|^{n+1}\max_{|u|\leq|x|}|f^{(n+1)}(u)|,\quad x\in\mathbb{R},n\geq 0.$ (A.36)
+Applying $f(x)=e^{-x}$ to this inequality, we have $|e^{-x}-1+x|\leq\frac{1}{2}|x|^{2}e^{|x|}$. Hence, for each fixed $\theta\in\mathbb{R}$, as $r\downarrow 0$
+$e^{r\theta\ell^{(r)}_{0}}=e^{\theta(\ell_{0}+o(r))}=e^{\theta\ell_{0}}(1+\theta o(r))+o(\theta o(r))=e^{\theta\ell_{0}}(1+o(\theta r)),$
+$e^{r\theta(\ell^{(r)}_{0}+1)}=e^{\theta r\ell^{(r)}_{0}}e^{r\theta}=e^{\theta\ell_{0}}(1+o(\theta r))(1+\theta r+o((\theta r)^{2}))=e^{\theta\ell_{0}}(1+\theta r+o(\theta r)).$
+Thus, (4.37) is obtained. (4.38) and (4.7) are immediate from these asymptotic expansions. ∎
 Proof of Lemma 4.8.
-It follows from Lemma 4.2 that
-$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}=1-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}+\frac{1}{2}|\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}|^{2}e^{|\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}|}$
-$\quad=1+(\mu r\theta_{2}+O(r^{2}))\widehat{R}^{(r)}_{s}+O(r^{2})(\widehat{R}^{(r)}_{s})^{2}$
-$\quad=1+\mu r\theta_{2}\widehat{R}^{(r)}_{s}+O(r^{2})(\widehat{R}^{(r)}_{s})^{2}.$
-This proves (4.39), and (4.8) and (4.8) are similarly proved. ∎
+Since $\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}$ and $\eta^{(r)}_{i}(r\theta)\widehat{R}^{(r)}_{e}$ are uniformly bounded in $r<a$ for a sufficiently small $a>0$ and vanish as $r\downarrow 0$, it follows from Lemma 4.2 and (A.4) that
+$e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}=1-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}+\frac{1}{2}(\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-})^{2}+O\big{(}|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
+$\quad=1-\left(-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}(r\theta)^{2}+o((r\theta)^{2})\right)\widehat{R}^{(r)}_{s-}$
+$\qquad+\frac{1}{2}\left(-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}(r\theta)^{2}+o((r\theta)^{2})\right)^{2}(\widehat{R}^{(r)}_{s-})^{2}$
+$\qquad+(\widehat{R}^{(r)}_{s-})^{3}O\big{(}(r\theta)^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
+$\quad=1+r\theta(\mu^{(r)}\widehat{R}^{(r)}_{s-})-\frac{1}{2}(r\theta)^{2}\left((\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\right)+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}),$
+where recall that $S_{3}(\widehat{R}^{(r)}_{s-})=\widehat{R}^{(r)}_{s-}+(\widehat{R}^{(r)}_{s-})^{2}+(\widehat{R}^{(r)}_{s-})^{3}$. This proves (4.8), and (4.8) and (4.8) are similarly proved. ∎
+A.5 Proofs of Lemmas 5.1 and 5.2
+We first prove Lemma 5.1, namely, that (5.3) implies (5.4). To prove this, we use the BAR (4.1) for $f(z,x_{1},x_{2})=(x_{1}\wedge 1/r)(x_{2}\wedge 1/r)1(z=\ell^{(r)}_{0})$, then we have
+$\mathbb{E}[R^{(r)}_{e}1(R^{(r)}_{e}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]+\mathbb{E}[R^{(r)}_{s}1(R^{(r)}_{s}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]$
+$\quad=\alpha^{(r)}_{e}\mathbb{E}[\widehat{T}^{(r)}_{e,2}]\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\alpha^{(r)}_{e}\mathbb{E}[\widehat{T}^{(r)}_{s}]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (A.37)
+Then, we can see that the left-hand side of this equation vanishes as $r\downarrow 0$ by (5.3) because, by Schwartz’s inequality,
+$\big{(}\mathbb{E}[R^{(r)}_{v}1(R^{(r)}_{v}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]\big{)}^{2}\leq\mathbb{E}[(R^{(r)}_{v})^{2}]\mathbb{P}[L^{(r)}=\ell^{(r)}_{0}],\qquad v=e,s.$
+Hence, (5.4) follows from (A.5) because $\mathbb{E}[\widehat{T}^{(r)}_{s}]$ converges to $\mu^{-1}$ as $r\downarrow 0$. Thus, Lemma 5.1 is proved.
+We next prove Lemma 5.2. By Schwartz’s inequality,
+$\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0}]\big{\}}^{2}=\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3/2}(\widehat{R}^{(r)}_{e})^{1/2}1^{2}(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}^{2}$
+$\quad\leq\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})].$
+Since $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded in $r$ by Corollary 4.4, (5.4) implies (5.5). This complete the proof of Lemma 5.2.
 Acknowledgement
-I thank the editor and referees for their helpful comments and careful reading.
 This study is motivated by the BAR approach, coined by Jim Dai (e.g., see [4]).
-I am grateful to him for his grand design about the BAR approach, and to Rami Atar for providing his recent paper [1].
+I am grateful to him for his grand design about it, and to Rami Atar for providing his recent papers [1, 2].
 I also benefit from Evsey Morozov for his comments on this paper.

### v8

@@
 Abstract
-We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue.
-This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed.
+We consider a single server queue which has a threshold to change its arrival process and service speed by its queue length, which is referred to as a two-level single server queue.
+This model is motivated by an energy saving problem for a single server queue whose arrival process and service speed are controlled.
 To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic.
-This limit corresponds to its diffusion approximation.
+Except for a special case, this limit corresponds to its diffusion approximation.
 It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters and generally distributed inter-arrival times and workloads brought by customers.
@@
 1 Introduction
-We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length.
-Namely, customers arrive subject to a renewal process as long as its queue length (including a customer being served) is less than threshold $\ell_{0}$, while it is changed to another renewal process otherwise.
-The service rate is changed similarly but in a slightly different way.
-That is, the service rate is constant when the queue length is not greater than the threshold, while it is another constant otherwise.
+We consider a single server queue which has a threshold to change its arrival process and g by its queue length.
+Namely, customers arrive subject to a renewal process as long as its queue length (including a customer being served) is less than threshold $\ell_{1}$, while it is changed to another renewal process otherwise.
+The service speed is changed similarly but in a slightly different way.
+That is, the service speed is constant when the queue length is not greater than the threshold, while it is another constant otherwise.
 We assume that customers are served in the FCFS manner, and the amounts of the work which are brought by them are independent and identically distributed and independent of the arrival process.
 We refer to this model as a 2-level $GI/G/1$ queue.
-This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service rate.
-Here, the system is an internet server, and its energy consumption depends on service rate.
-Then, the problem is to find an optimal threshold to minimize the operating energy of the server and customer’s waiting cost.
+This 2-level $GI/G/1$ queue is motivated by an energy saving problem to control the arrival process and service speed.
+Here, the system is an internet server, and its energy consumption depends on service speed.
+Then, the problem is to find an optimal threshold to minimize the total cost of the operating energy of the server and customer’s waiting.
 So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length.
-However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured in this exponential case.
-Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution.
-This limiting distribution corresponds to the stationary distribution of a one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold, which is recently studied in [11, 13].
+However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available, but the variability of the arrival times and service amounts are not well captured in this exponential case.
+Because of these facts, an approximation by a reflected diffusion is helpful, where a reflected diffusion means a one-dimensional diffusion reflected at the origin, in which the term “diffusion” is used for a one-dimensional continuous Markov process on the real line which has a finite and positive second moment at each time.
+So we study the diffusion approximation through the limit of the stationary distributions of the scaled queue length of this 2-level $GI/G/1$ queue in heavy traffic, where the scaling factor and threshold level are chosen so that the limit is well obtained as a probability distribution.
+Except for a special case, this limiting distribution corresponds to the stationary distribution of a reflected diffusion whose drift and variance are discontinuously changed at the level corresponding to the threshold (see Section 6.1).
+The stationary distribution of this reflected diffusion is recently studied in [11, 13].
 Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
 There is another motivation of this study.
-In general, theoretical study of diffusion approximation is hard for state-dependent queues.
-For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see [8, 15]), while the arrival process and service rate discontinuously change in the present model.
+In general, it is hard to theoretically study diffusion approximation for state-dependent queues.
+For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see [8, 15]), while the arrival process and service speed discontinuously change in the present model.
 Thus, theoretical study for its diffusion approximation is challenging.
@@
 Furthermore, [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room.
-Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see [6]).
-Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it.
+Both distributions agree with the stationary distributions of the corresponding reflected diffusions (see [6]).
+Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential or uniform below the threshold level and exponential above it.
 This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional.
 Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [3, 4, 10, 12].
-This verification, Theorem 3.1, requires a mild limitation on the parameters of the arrival processes (see (3.f)).
+This verification, Theorem 3.1, requires a mild limitation on the parameters of the arrival process (see (3.f) in Section 3.1).
 Its proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
-However, the proof has a complicated structure, so we split it into many components so that they are orthogonal as long as possible.
+However, the proof is complicated, so we split it into several components so that they are orthogonal as long as possible.
 This is the reason why we spend two sections for the proof.
-In this paper, we only study the limit of the scaled stationary distribution, but, as our results are related to the diffusion approximation, it is interesting to identify a reflected Brownian motion as a process limit for the 2-level $GI/G/1$ queue.
-This is a hard question because this Brownian motion discontinuously changes its drift and variance at the fixed state.
+In this paper, we only study the limit of the scaled stationary distribution, but, as our results are related to the diffusion approximation, it is interesting to identify a reflected diffusion as a process limit for the 2-level $GI/G/1$ queue.
+This is a hard question because this diffusion discontinuously changes its drift and variance at the fixed state.
 We intuitively discuss what can be expected about this process limit in Section 6.1.
@@
 Let $L(t)$, $R_{e}(t)$ and $R_{s}(t)$ be the queue length, the remaining arrive time and the remaining workload of a customer being served, respectively, at time $t$.
-Note that the remaining service time cannot be used for $R_{s}(t)$ here because the service rate may change in future.
+Note that the remaining service time cannot be used for $R_{s}(t)$ here because the service speed may change in future.
 Define
@@
 We denote those point processes by $N_{e}$ and $N_{s}$ respectively.
-According to [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
+According to [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\Theta_{\bullet}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
 (2.a) Measurability: $A\in{\mathcal{F}}_{t}\;\mbox{ implies }\;\Theta_{s}^{-1}A\in{\mathcal{F}}_{s+t},\qquad s,t\geq 0,$
@@
 Since this sequence can be countable, one may take $r=1/\sqrt{n}$ for positive integer $n$ so that the scaling is identical with the standard diffusion approximation, but we keep $r$ just for convenience.
+Let $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper.
 3.1 Assumptions
-For the $r$-th model, let $\ell^{(r)}_{0}\geq 2$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$.
+For the $r$-th model, let $\ell^{(r)}_{1}\geq 1$ be the threshold level, and let $L^{(r)}(t)$ be the number of customers in the system at time $t$.
 We refer to this number as the queue length.
@@
 We assume that $T_{e}^{(r)}(n)\equiv t^{(r)}_{e,n}-t^{(r)}_{e,n-1}$ is independent of $\{t^{(r)}_{e,n^{\prime}};0\leq n^{\prime}\leq n-1\}$ and everything else for all $n\geq 1$, where $t^{(r)}_{e,0}=0$.
-Thus, the inter-arrival times are independent, but they may have two different distributions depending on the queue length.
-Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{0}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{0}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are independent random variables which may have different distributions.
+Thus, the inter-arrival times are independent, but their distributions depend on the queue length.
+Namely, $T_{e}^{(r)}(n)$ has the same distribution as $T^{(r)}_{e,1}$ if $L(t^{(r)}_{e,n-1}-)<\ell^{(r)}_{1}$ while its distribution agrees with that of $T_{e,2}^{(r)}$ if $L(t^{(r)}_{e,n-1}-)\geq\ell^{(r)}_{1}$, where $T^{(r)}_{e,1}$ and $T^{(r)}_{e,2}$ are independent random variables which may have different distributions.
 As for service, let $T^{(r)}_{s}(n)$ be the workload of the $n$-th arriving customers.
-We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variable.
+We assume that $\{T^{(r)}_{s}(n);n\geq 1\}$ is a sequence of $i.i.d.$ random variables.
 We denote a random variable subject to the common distribution of $T^{(r)}_{s}(n)$ by $T^{(r)}_{s}$.
-Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{0}$ and otherwise, respectively.
+Let $c^{(r)}_{1}$ and $c^{(r)}_{2}$ be the service speeds when the queue length is not greater than $\ell^{(r)}_{1}$ and otherwise, respectively.
 Namely, as long as $R^{(r)}_{s}(t)$ is differentiable at time $t$,
-$\frac{d}{dt}R^{(r)}_{s}(t)=\left\{\begin{array}[]{ll}-c^{(r)}_{1},&L^{(r)}(t)\leq\ell^{(r)}_{0},\\ -c^{(r)}_{2},&L^{(r)}(t)>\ell^{(r)}_{0}.\end{array}\right.$
-These dynamics is depicted in the sample path of $L^{(r)}(t)$ below.
-Figure 3.1: Sample path of the queue length process
+$\frac{d}{dt}R^{(r)}_{s}(t)=-c^{(r)}_{1}1(1\leq L^{(r)}(t)\leq\ell^{(r)}_{1})-c^{(r)}_{2}1(L^{(r)}(t)>\ell^{(r)}_{1}),$
+where $1(\cdot)$ is the indicator function of the statement $\cdot$”. These dynamics is depicted in the sample path of $L^{(r)}(t)$ below.
+Figure 3.1: Sample path of the queue length process of the $r$-th model
 We will use the following notations.
-$m^{(r)}_{e,i}=\mathbb{E}(T^{(r)}_{e,i}),\qquad\lambda^{(r)}_{i}=1/m^{(r)}_{e,i},\qquad\sigma^{(r)}_{e,i}=\mathbb{E}[(T^{(r)}_{e,i}-m^{(r)}_{e,i})^{2}],\qquad i=1,2,$
-$m^{(r)}_{s}=\mathbb{E}(T^{(r)}_{s}),\qquad\mu^{(r)}=1/m^{(r)}_{s},\qquad\sigma^{(r)}_{s}=\mathbb{E}[(T^{(r)}_{s}-m^{(r)}_{s})^{2}],$
+$m^{(r)}_{e,i}=\mathbb{E}(T^{(r)}_{e,i}),\qquad\lambda^{(r)}_{i}=1/m^{(r)}_{e,i},\qquad\sigma^{(r)}_{e,i}=(\mathbb{E}[(T^{(r)}_{e,i}-m^{(r)}_{e,i})^{2}]^{-\frac{1}{2}},\qquad i=1,2,$
+$m^{(r)}_{s}=\mathbb{E}(T^{(r)}_{s}),\qquad\mu^{(r)}=1/m^{(r)}_{s},\qquad\sigma^{(r)}_{s}=(\mathbb{E}[(T^{(r)}_{s}-m^{(r)}_{s})^{2}])^{-\frac{1}{2}},$
 $\rho^{(r)}_{i}=\lambda^{(r)}_{i}/(c^{(r)}_{i}\mu^{(r)}),\qquad i=1,2.$
-For the $r$-th model, at time $t$, let $R^{(r)}_{e}(t)$ be the remaining arrival time, and let $R^{(r)}_{s}(t)$ be the remaining workload of a customer being served.
+For the $r$-th model, let $R^{(r)}_{e}(t)$ and $R^{(r)}_{s}(t)$ be the remaining arrival time and the remaining workload of a customer being served, respectively, at time $t$.
 Define $X^{(r)}(\cdot)\equiv\{X^{(r)}(t);t\geq 0\}$ by
@@
 and denote the counting processes of the arrivals and service completions by $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$, respectively.
-Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov.
-Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with the time-shift operator $\Psi$.
+Note that $X^{(r)}(\cdot)$ is a piecewise deterministic Markov process, so it is strong Markov (e.g., see [5]).
+Obviously, $X^{(r)}(\cdot)$, $N^{(r)}_{e}(\cdot)$ and $N^{(r)}_{s}(\cdot)$ are consistent with the time-shift operator $\Theta_{\bullet}$.
 We assume the following assumptions.
@@
   • (3.b)
-    For $i=1,2$, $m^{(r)}_{e,i},\lambda^{(r)}_{i},\sigma^{(r)}_{e,i},c^{(r)}_{i}$ converge to $m_{e,i},\lambda_{i},\sigma_{e,i},c_{i}>0$, respectively, as $r\downarrow 0$.
-Similarly, $m^{(r)}_{s},\mu^{(r)},\sigma^{(r)}_{s}$ converge to $m_{s},\mu,\sigma_{s}$, respectively, as $r\downarrow 0$.
+    For $i=1,2$, $m^{(r)}_{e,i},\lambda^{(r)}_{i},\sigma^{(r)}_{e,i},c^{(r)}_{i}$ and $m^{(r)}_{s},\mu^{(r)},\sigma^{(r)}_{s}$ are finite and positive, and converge to $m_{e,i},\lambda_{i},\sigma_{e,i},c_{i}>0$ and $m_{s},\mu,\sigma_{s}>0$, respectively, as $r\downarrow 0$.
   • (3.c)
@@
   • (3.d)
-    There is a constant $\ell_{0}$ such that $r\ell^{(r)}_{0}=\ell_{0}+o(r)$ as $r\downarrow 0$.
+    There is a constant $\ell_{1}$ such that $r\ell^{(r)}_{1}=\ell_{1}+o(r)$ as $r\downarrow 0$.
   • (3.e)
     $\rho^{(r)}_{2}<1$.
-Here, for positive valued functions $f,g$ of $r$, we write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$.
+Here, for functions $f,g:(0,1]\to(0,\infty)$, we write $f(r)=o(g(r))$ as $r\downarrow 0$ if $\lim_{r\downarrow 0}f(r)/g(r)=0$.
 Similarly, $f(r)=O(g(r))$ if $\limsup_{r\downarrow 0}f(r)/g(r)<\infty$.
@@
 Remark 3.1.
-The condition (3.f) comes out from technical requirements for the proof of Theorem 3.1.
-It is a mild limitation because no condition is needed for $\lambda_{1}\not=\lambda_{2}$, but we think that it is likely not needed.
-However, removing (3.f) seems to be beyond the framework of this paper.
+This is a mild limitation because no condition is needed for $\lambda_{1}\not=\lambda_{2}$ and it is automatically satisfied when $T_{e,1}$ and $T_{e,2}$ are identically distributed.
+This condition is likely not needed but hard to remove.
 See Remark 5.1 for its details.
@@
 For $i=1,2$, let $N^{(r)}_{e,i}$ be the stationary point process obtained from the renewal process whose inter-counting times have the same distribution as that of $T^{(r)}_{e,i}$.
-Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$, and let $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$, which will be used throughout the paper.
-We note the following simple but important facts.
+Similarly, let $N^{(r)}_{s}$ be the stationary point process for the service completion instants.
+Define $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ and $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]$.
+At present, we cannot compute $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$, but the following lemma is intuitively clear because $X^{(r)}(\cdot)$ is a stationary process by (3.e).
+Since this lemma is a key for our analysis, we formally prove it in Appendix A.1.
 Lemma 3.1.
-Assume the assumptions (3.a)–(3.e) and the stationarity of $X^{(r)}(\cdot)$, then $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
-Proof.
-Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$.
-Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework.
-Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$.
-Similarly, let $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1}))^{-1}T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1})^{-1})T_{s}]<\infty$.
-∎
+Assume the assumptions (3.b) and (3.e), then $\alpha^{(r)}_{e}$ is finite and positive, and $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$.
 3.2 Main results
-We are now ready to present a main result, which will be proved in Section 5.
+We now present a main result, which will be proved in Section 5.
 Recall that $L^{(r)}$ is a random variable subject to the stationary distribution of the $r$-th 2-level $GI/G/1$ queue.
@@
 We refer to this $\nu^{(r)}$ as a scaled stationary distribution.
-We decompose $\nu^{(r)}$ using finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$, which are defined as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
-$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;\ell^{(r)}_{0}<L^{(r)}),$
-$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|\ell^{(r)}_{0}<L^{(r)}).$
+Define finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as
+$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{1}),\qquad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;\ell^{(r)}_{1}<L^{(r)}),$
+$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{1}),\qquad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|\ell^{(r)}_{1}<L^{(r)}),$
+for $B\in{\mathcal{B}}(\mathbb{R}_{+})$. Obviously, $\nu^{(r)}(B)=\nu^{(r)}_{1}(B)+\nu^{(r)}_{2}(B)$.
 Theorem 3.1.
@@
 For each $r\in(0,1]$, define $\widetilde{\nu}_{i|cd}$ as the probability measures on ${\mathcal{B}}(\mathbb{R}_{+})$ which have the density functions $h_{i|cd}$ for $i=1,2$ such that
-$h_{1|cd}(x)=1(0\leq x\leq\ell_{0})\times\left\{\begin{array}[]{ll}\displaystyle 1/\ell_{0},&b_{1}=0,\\ \displaystyle\frac{e^{\beta_{1}\ell_{0}}\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}e^{-\beta_{1}x},&b_{1}\not=0,\end{array}\right.$ (3.4)
-$h_{2|cd}(x)=1(x>\ell_{0})\beta_{2}e^{-\beta_{2}(x-\ell_{0})}.$ (3.5)
+$h_{1|cd}(x)=1(0\leq x\leq\ell_{1})\times\left\{\begin{array}[]{ll}\displaystyle 1/\ell_{1},&b_{1}=0,\\ \displaystyle\frac{e^{\beta_{1}\ell_{1}}\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}e^{-\beta_{1}x},&b_{1}\not=0,\end{array}\right.$ (3.4)
+$h_{2|cd}(x)=1(x>\ell_{1})\beta_{2}e^{-\beta_{2}(x-\ell_{1})}.$ (3.5)
   • (i)
@@
     where $\widetilde{A}_{i}$ for $i=1,2$ are given by
-    $\widetilde{A}_{1}=\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)},\;&b_{1}\not=0,\end{cases}\qquad\widetilde{A}_{2}=1-\widetilde{A}_{1}.$ (3.10)
+    $\widetilde{A}_{1}=\begin{cases}\displaystyle\frac{2b_{2}\ell_{1}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{1}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)},\;&b_{1}\not=0,\end{cases}\qquad\widetilde{A}_{2}=1-\widetilde{A}_{1}.$ (3.10)
     If either $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$, equivalently, $c^{(r)}_{1}-c^{(r)}_{2}=O(r)$, or $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,1]$, then
@@
 In diffusion applications of a single server queue, the mean queue length is estimated through $(1-\rho^{(r)})\mathbb{E}[L^{(r)}]$ as $r\downarrow 0$.
-The following corollary computes this limit, whose details are given in Appendix A.1.
+This limit is computed in the following corollary, which is proved in Appendix A.2.
 Corollary 3.1.
@@
 $\lim_{r\downarrow 0}(1-\rho^{(r)})\mathbb{E}[L^{(r)}]=\lim_{r\downarrow 0}\mathbb{P}(L^{(r)}=0)\mathbb{E}[L^{(r)}]$
-=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.15)
+=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{1}(b_{2}\ell_{1}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{1}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)+2b_{1}b_{2}\ell_{1}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.15)
 Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed.
@@
 Under the assumptions (3.a)–(3.e), the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
-$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{0}}{1+\beta_{2}\ell_{0}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.16)
+$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{1}}{1+\beta_{2}\ell_{1}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.16)
 and $\widetilde{A}_{\exp,2}=1-\widetilde{A}_{\exp,1}$.
@@
 However, for a double check, we give its direct proof in Appendix A.3, which is helpful to convince (ii) of Theorem 3.1.
-This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of greater than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
+This proof is not obvious because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that an arrival find the queue length less than $\ell_{1}$ when the previous arrival found the queue length greater than $\ell_{1}$ but multiple departures occurred (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1).
 This Markov chain will be detailed in Appendix A.3.
@@
 In this section, we prepare auxiliary notations and results for proving Theorem 3.1.
-A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$.
+A major difficulty for this proof comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain, and we have to work with $X^{(r)}(\cdot)$.
 Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below.
@@
 Recently, this approach is further studied in [12].
-Throughout this section, we assume (3.a)–(3.e), but (3.f) is not assumed. The proof of Theorem 3.1 will be done in Section 5.
+Throughout this section, we assume (3.a)–(3.e) unless stated otherwise, but (3.f) is not assumed. The proof of Theorem 3.1 will be done in Section 5.
 4.1 Palm distribution and BAR
@@
 Then, for each $r\in(0,1]$, $X^{(r)}(\cdot)$ and the increments of $N^{(r)}_{e}$ and $N^{(r)}_{s}$ are jointly stationary.
-Since $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite by Lemma 3.1, we can define probability distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ on $(\Omega,{\mathcal{F}})$ by
+Since $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite and positive by Lemma 3.1, we can define probability distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ on $(\Omega,{\mathcal{F}})$ by
 $\mathbb{P}^{(r)}_{v}(A)=(\alpha^{(r)}_{v})^{-1}\mathbb{E}\left[\int_{(0,1]}1_{\Theta_{u}^{-1}A}N^{(r)}_{v}(du)\right],\qquad A\in{\mathcal{F}},v=e,s,$ (4.1)
@@
 For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$.
-Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded and continuous functions from $S$ to $\mathbb{R}_{+}$ which are partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right.
+Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded and continuous functions from $S$ to $\mathbb{R}_{+}$ which are partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}f$ is the $i$-th partial derivative from the right.
 Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
-${\mathcal{H}}f(z,{\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}})-(c_{1}1(1\leq z\leq\ell_{0})+c_{2}1(z>\ell_{0}))\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}}),\quad(z,{\boldsymbol{x}})\in S,$
-where ${\boldsymbol{x}}=(x_{1},x_{2})\in\mathbb{R}_{+}^{2}$.
+${\mathcal{H}}f({\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f({\boldsymbol{x}})-(c_{1}1(1\leq x_{1}\leq\ell_{1})+c_{2}1(x_{1}>\ell_{1}))\frac{\partial^{+}}{\partial x_{3}}f({\boldsymbol{x}}),\qquad{\boldsymbol{x}}\in S,$
+where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})\in\mathbb{R}_{+}^{2}$.
 In what follows, we often use a difference operator $\Delta$, which is defined as $\Delta g(t)=g(t)-g(t-)$ for function $g$ from $\mathbb{R}_{+}$ to $\mathbb{R}$ which is right continuous with the left-limit.
 For $t\geq 0$, define $\mathbb{F}$-adapted process $X^{(r)}_{1}(\cdot)$ and $L^{(r)}_{1}(\cdot)$ as
-$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
+$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{1},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{1},\end{cases}$
 $L^{(r)}_{1}(t)=L^{(r)}(t-)+1(\Delta N^{(r)}_{e}(t)=1),\qquad R^{(r)}_{1}(t)=(R^{(r)}_{e}(t),R^{(r)}_{s}(t-)),$
-where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$.
+where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{1})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{1})$.
 We call $X^{(r)}_{1}(t)$ an intermediate state.
-Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$ when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
+Note that
+$X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$
+when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
 In these definitions, $X^{(r)}(0-)$ is required, but formally not defined.
@@
 For convenience, we define another difference operators $\Delta_{e}$ and $\Delta_{s}$ as
-$\Delta_{e}f(X^{(r)}(t))=f(X^{(r)}_{1}(t))-f(X^{(r)}(t-)),$
-$\Delta_{s}f(X^{(r)}(t))=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)).$
-Then, from the elementary observation of the time evolution of $X^{(r)}(t)$, we have
-$f(X^{(r)}(t)=f(X^{(r)}(0))+\int_{0}^{t}{\mathcal{H}}f(X^{(r)}(u))du+\sum_{u\in(0,t]}[\Delta_{e}f(X^{(r)}(u))+\Delta_{s}f(X^{(r)}(u))],$
-where the summation is well defined because $\Delta_{e}f(X^{(r)}(u))$ and $\Delta_{s}f(X^{(r)}(u))$ vanish except for finitely many $u\in(0,t]$.
-Take the expectation of the above equation for $t=1$ by $\mathbb{P}$, then the stationarity of $X(\cdot)$ and the definition of the Palm distribution yield
+$\Delta_{e}f(X^{(r)})(t)=f(X^{(r)}_{1}(t))-f(X^{(r)}(t-)),$
+$\Delta_{s}f(X^{(r)})(t)=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)).$
+Further, let $f(X)(t)=f(X(t))$. Then, from the time evolution of $X^{(r)}(t)$, we have
 […]
-$\mathbb{E}\left[\sum_{u\in(0,1]}\Delta_{v}f(X^{(r)}(u))]\right]=\alpha^{(r)}_{v}\mathbb{E}^{(r)}_{v}\left[\Delta_{v}f(X^{(r)}(0))\right],\qquad v=e,s.$
+$+\sum_{u\in(0,t]}[\Delta_{e}f(X^{(r)})(u)+\Delta_{s}f(X^{(r)})(u)],$ (4.2)
+where the summation is well defined because $\Delta_{e}f(X^{(r)})(u)$ and $\Delta_{s}f(X^{(r)})(u)$ vanish except for finitely many $u\in(0,t]$.
+Take the expectation of (4.1) for $t=1$ by $\mathbb{P}$, then the stationarity of $X(\cdot)$ and the definition of the Palm distribution yield
+[…]
+$\mathbb{E}\left[\sum_{u\in(0,1]}\Delta_{v}f(X^{(r)})(u)]\right]=\alpha^{(r)}_{v}\mathbb{E}^{(r)}_{v}\left[\Delta_{v}f(X^{(r)})(0)\right],\qquad v=e,s.$
 Hence, we have a stationary equation:
 $\mathbb{E}\left[{\mathcal{H}}f(X^{(r)}(0))\right]$
-$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]=0,\qquad f\in C_{b}^{p},$ (4.2)
+$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)})(0)\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)})(0)\right]=0,\qquad f\in C_{b}^{p},$ (4.3)
 The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short, then we call analysis based on this BAR as a BAR approach.
 A major difficulty in this BAR approach is to evaluate the expectations under the Palm distributions (the last two terms in (4.1)).
-This evaluation will be done for the 2-level $GI/G/1$ queue in Section 4.3.
+Their evaluations will be done for the 2-level $GI/G/1$ queue in Section 4.3.
 Remark 4.1.
-In using Palm distributions, we must be careful about the following two points.
-  • (i)
-    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$.
-So, $X^{(r)}(0-)$ is a random variable well defined under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$.
-Hence, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions.
-If we define the stationary process $X^{(r)}(\cdot)$ on the time axis (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ are well defined even under $\mathbb{P}$.
-  • (ii)
-    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00.
+We give two remarks on Palm distributions.
+(i) $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$.
+So, $X^{(r)}(0-)$ is a random variable which can be defined under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$.
+Hence, $\Delta_{e}f(X^{(r)})(0)$ and $\Delta_{s}f(X^{(r)})(0)$ in (4.1) are well defined under the Palm distributions.
+(ii) Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00.
 This means that we take a so called arrival first framework.
 As for the intermediate state, see [12] for its details.
-We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$.
+We consider $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$.
 This prevents to directly refer to the intermediate state $X^{(r)}_{1}(0)$ in our arguments.
@@
 Lemma 4.1.
-Under conditions (3.a) and (3.c), $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$, and
+Under conditions (3.b) and (3.e),
 $\mathbb{P}^{(r)}_{e}[(L^{(r)}(0-)=k]=\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=k],\qquad k\geq 0,$ (4.4)
-$\alpha^{(r)}_{e}=\frac{\lambda^{(r)}_{1}\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)<\ell^{(r)}_{0})+\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)\geq\ell^{(r)}_{0})},$ (4.4)
-$\alpha^{(r)}_{e}=c^{(r)}_{1}\mu^{(r)}\mathbb{P}(1\leq L^{(r)}\leq\ell^{(r)}_{0})+c^{(r)}_{2}\mu^{(r)}\mathbb{P}(L^{(r)}>\ell^{(r)}_{0}).$ (4.5)
+$\alpha^{(r)}_{e}=\frac{\lambda^{(r)}_{1}\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)<\ell^{(r)}_{1})+\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)\geq\ell^{(r)}_{1})},$ (4.5)
+$\alpha^{(r)}_{e}=c^{(r)}_{1}\mu^{(r)}\mathbb{P}(1\leq L^{(r)}\leq\ell^{(r)}_{1})+c^{(r)}_{2}\mu^{(r)}\mathbb{P}(L^{(r)}>\ell^{(r)}_{1}).$ (4.6)
 Proof.
 For each integer $k\geq 0$, let $f(z,{\boldsymbol{x}})=z\wedge(k+1)$ for $(z,{\boldsymbol{x}})\in S$, then obviously $f\in C_{b}^{p}(S)$.
-Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
-$\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]$
+Since $\Delta_{e}f(X^{(r)})(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)})(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
+$\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)})(0)\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)})(0)\right]$
 $\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{P}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
@@
 $\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{P}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
-Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$.
-Because of this equality, the above formula also yields (4.3).
-We next let $f(z,{\boldsymbol{x}})=x_{1}$, and apply (4.1), then
-$-1+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[T^{(r)}_{e,1}1(L^{(r)}(0-)<\ell^{(r)}_{0})+T^{(r)}_{e,2}1(L^{(r)}(0-)\geq\ell^{(r)}_{0})\right]=0$
+Since $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$ by Lemma 3.1, the above formula yields (4.4). We next let $f(z,{\boldsymbol{x}})=x_{1}$, and apply (4.1), then
+$-1+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[T^{(r)}_{e,1}1(L^{(r)}(0-)<\ell^{(r)}_{1})+T^{(r)}_{e,2}1(L^{(r)}(0-)\geq\ell^{(r)}_{1})\right]=0$
 because $(R^{(r)}_{e})^{\prime}(t)=-1$. This proves (4.5). Similarly, let $f(z,{\boldsymbol{x}})=x_{2}$, then
-$(R^{(r)}_{s})^{\prime}(t)=-c_{1}^{(r)}1(1\leq L^{(r)}(0)\leq\ell^{(r)}_{0})-c_{2}^{(r)}1(L^{(r)}(0)>\ell^{(r)}_{0}),$ (4.6)
+$(R^{(r)}_{s})^{\prime}(t)=-c_{1}^{(r)}1(1\leq L^{(r)}(0)\leq\ell^{(r)}_{1})-c_{2}^{(r)}1(L^{(r)}(0)>\ell^{(r)}_{1}),$ (4.7)
 and $\mathbb{E}^{(r)}_{s}[\Delta_{s}R^{(r)}_{s}(0)]=\mathbb{E}[T^{(r)}_{s}]$. Hence, (4.1) yields (4.6). ∎
-Since it follows from (4.4) that
-$\min(\lambda^{(r)}_{1},\lambda^{(r)}_{2})\leq\alpha^{(r)}_{e}\leq\max(\lambda^{(r)}_{1},\lambda^{(r)}_{2}),\qquad r\in(0,1],$ (4.7)
+Since it follows from (4.5) and (3.b) that
+$0<\lambda^{(r)}_{1}\wedge\lambda^{(r)}_{2}\leq\alpha^{(r)}_{e}\leq\lambda^{(r)}_{1}\vee\lambda^{(r)}_{2}<\infty,\qquad r\in(0,1],$ (4.8)
 we have the following corollary.
 Corollary 4.1.
-$\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
-To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1.
+Under conditions (3.b) and (3.e), $\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
+To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$ because $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$ by Lemma 3.1.
 We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 5.
@@
 Namely, for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$,
-$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(\ell^{(r)}_{0}<L^{(r)})],$
-$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|\ell^{(r)}_{0}<L^{(r)}].$
+$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{1})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(\ell^{(r)}_{1}<L^{(r)})],$
+$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{1}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|\ell^{(r)}_{1}<L^{(r)}].$
 To this end, we introduce test functions for the BAR (4.1). We first define functions $w^{(r)}_{i}$ and $g^{(r)}_{i,\theta}$ for $i=1,2$ and $\theta\in\mathbb{R}$ as
-$w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{0}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{0}),\qquad z\in\mathbb{Z}_{+},$
+$w^{(r)}_{1}(z)=1(z\leq\ell^{(r)}_{1}),\qquad w^{(r)}_{2}(z)=1(z>\ell^{(r)}_{1}),\qquad z\in\mathbb{Z}_{+},$
 $g^{(r)}_{i,\theta}(x_{1},x_{2})=\exp\left(-\eta^{(r)}_{i}(\theta)(x_{1}\wedge 1/r)-\zeta^{(r)}(\theta)(x_{2}\wedge 1/r)\right),\qquad(x_{1},x_{2})\in\mathbb{R}_{+}^{2},$
@@
 We will take this convention throughout the paper as long as it is possibe.
-By this lemma, $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ is bounded by $e^{|\theta_{i}|(d_{e,i}+d_{s})}$ for sufficiently small $r$ for each $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, and therefore, we have the following facts.
+Recall that (3.a)–(3.e) are always assumed in Section 4 if particular assumptions among them are specified.
+By Lemma 4.2, $g^{(r)}_{i,r\theta_{i}}(R^{(r)})$ is bounded by $e^{|\theta_{i}|(d_{e,i}+d_{s})}$ for sufficiently small $r$ for each $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, so we have the following facts.
 Corollary 4.2.
@@
 $\psi^{(r)}_{i}(\theta_{i})<\infty,\quad\mathbb{E}_{u}\left[f^{(r)}_{r{\boldsymbol{\theta}}}(L^{(r)},R^{(r)})\right]<\infty,\quad u=e,s,\quad\mbox{uniformly in $r\in(0,1]$},$ (4.17)
-because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$.
+because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{1})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$.
 These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
@@
 For $\theta_{1}\in\mathbb{R}_{+}$ and $\theta_{2}\leq 0$
-$\lim_{r\downarrow 0}\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|=0,\quad i=1,2.$ (4.17)
+$\lim_{r\downarrow 0}\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i})\right|=0,\quad i=1,2.$ (4.18)
 Proof.
 By Corollary 4.2, we have
-$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{0})]\right|$
-$\leq e^{|\theta_{1}|r\ell^{(r)}_{0}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
+$\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i})\right|$ $=\left|\mathbb{E}[e^{r\theta_{i}L^{(r)}}(g_{1,\theta_{i}}(R^{(r)}_{e})-1)1(L^{(r)}\leq\ell^{(r)}_{1})]\right|$
+$\leq e^{|\theta_{1}|r\ell^{(r)}_{1}}\left|\mathbb{E}[g_{i,\theta_{i}}(R^{(r)}_{e})-1]\right|\to 0,\qquad r\downarrow 0.$
 This proves (4.18). ∎
@@
 $\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s-}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$ (4.22)
-Since $g^{(r)}_{i,\theta_{i}}$ changes only when $L^{(r)}(0-)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{0}$ under $\mathbb{P}^{(r)}_{s}$, (4.8) implies that
-$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0-)\not=\ell^{(r)}_{0})\right]=0,$
-$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))1(L^{(r)}(0)\not=\ell^{(r)}_{0})\right]=0.$
-Hence, from the definitions of $X^{(r)}_{1}(0)$ and $\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))$, we have
-$\mathbb{E}^{(r)}_{e}[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{e}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad=\mathbb{E}^{(r)}_{e}\Big{[}e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0))1(L^{(r)}_{1}(0)=\ell^{(r)}_{0}+1)$
-$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{0})\Big{]}$
-$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
-$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.22)
+Since $g^{(r)}_{i,\theta_{i}}(R^{(r)}(t))$ changes at $t=0$ only when $L^{(r)}(0-)=\ell^{(r)}_{1}$ under $\mathbb{P}^{(r)}_{e}$ or $L^{(r)}(0)=\ell^{(r)}_{1}$ under $\mathbb{P}^{(r)}_{s}$, (4.9) implies that
+$\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)1(L^{(r)}(0-)\not=\ell^{(r)}_{1})\right]=0,$
+$\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)1(L^{(r)}(0)\not=\ell^{(r)}_{1})\right]=0.$
+Hence, from the definitions of $X^{(r)}_{1}(0)$ and $\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)$, we have
+$\mathbb{E}^{(r)}_{e}[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)]=\mathbb{E}^{(r)}_{e}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad=\mathbb{E}^{(r)}_{e}\Big{[}e^{r\theta_{2}(\ell^{(r)}_{1}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0))1(L^{(r)}_{1}(0)=\ell^{(r)}_{1}+1)$
+$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{1}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0-)))1(L^{(r)}(0-)=\ell^{(r)}_{1})\Big{]}$
+$\quad=e^{r\theta_{2}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1}))],$
+$\quad=(e^{r\theta_{2}\ell^{(r)}_{1}}-e^{r\theta_{1}\ell^{(r)}_{1}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\qquad+e^{r\theta_{1}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{1}))],$ (4.23)
 where the third equality is obtained by applying (4.9). Similarly, we have
-$\mathbb{E}^{(r)}_{s}[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))]=\mathbb{E}^{(r)}_{s}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0)))1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0-)=\ell^{(r)}_{0}+1)\Big{]}$
-$\quad=e^{r\theta_{1}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad=(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\qquad+e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.23)
-Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))\right]\right)$.
+$\mathbb{E}^{(r)}_{s}[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)]=\mathbb{E}^{(r)}_{s}[(f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}(0))-f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)}_{1}(0)))1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{1}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{1})$
+$\qquad-e^{r\theta_{2}(\ell^{(r)}_{1}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0-)=\ell^{(r)}_{1}+1)\Big{]}$
+$\quad=e^{r\theta_{1}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\qquad-e^{r\theta_{2}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\quad=(e^{r\theta_{1}(\ell^{(r)}_{1}+1)}-e^{r\theta_{2}(\ell^{(r)}_{1}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\qquad+e^{r\theta_{2}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.24)
+Define $D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}\left(\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)\right]+\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f^{(r)}_{r{\boldsymbol{\theta}}}(X^{(r)})(0)\right]\right)$.
 Then, by (4.2) and (4.2),
-$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}(e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}e^{r\theta_{2}(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.24)
+$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{1}}-e^{r\theta_{1}\ell^{(r)}_{1}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-})}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad+\alpha^{(r)}_{e}(e^{r\theta_{1}(\ell^{(r)}_{1}+1)}-e^{r\theta_{2}(\ell^{(r)}_{1}+1)})\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\quad+\alpha^{(r)}_{e}e^{r\theta_{2}(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta_{1})\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta_{2})\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.25)
 Thus, applying (4.2), (4.2) and (4.2) to (4.1), we have a pre-limit BAR for the 2-level $GI/G/1$ queue.
@@
 For each fixed $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$, we have, as $r\downarrow 0$,
-$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
+$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta_{i}r)^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
 $\quad-c_{1}\left(\mu r\theta_{1}-\mu^{3}\sigma_{s}^{2}r^{2}\theta_{1}^{2}/2+o((\theta_{1}r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})]=D^{(r)}(\theta_{1},\theta_{2}).$ (4.26)
@@
 (Step 1)   In this step, we show that the moments of $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ are well behaved as $r\downarrow 0$.
-This is also examined for $\widehat{R}^{(r)}_{s-}1(L^{(r)}=\ell^{(r)}_{0})$ and $\widehat{R}^{(r)}_{e}1(L^{(r)}=\ell^{(r)}_{0-})$ under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$, respectively.
+This is also examined for $\widehat{R}^{(r)}_{s-}1(L^{(r)}=\ell^{(r)}_{1})$ and $\widehat{R}^{(r)}_{e}1(L^{(r)}=\ell^{(r)}_{0-})$ under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$, respectively.
 These are basis for the expansion of the BAR.
@@
 Lemma 4.4.
-Under the assumption (3.a), for $k=0,1,2$ and $r\in(0,1]$,
+For $k=0,1,2$ and $r\in(0,1]$,
 $(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}],$ (4.27)
@@
 Lemma 4.5.
-Under the assumption (3.a), we have, for $k=1,2,3$ as $r\downarrow 0$,
-$c^{(r)}_{1}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(1\leq L^{(r)}\leq\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.30)
-$c^{(r)}_{2}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.31)
-$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}\leq\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,1})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]\Big{)},$ (4.32)
-$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{0})]$
-$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)}.$ (4.33)
+For $k=1,2,3$ and $r\in(0,1]$,
+$c^{(r)}_{1}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(1\leq L^{(r)}\leq\ell^{(r)}_{1})]$
+$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{1}]\Big{)},$ (4.31)
+$c^{(r)}_{2}k\mathbb{E}[(R^{(r)}_{s})^{k-1}(1-1(R^{(r)}_{s}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{1})]$
+$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{1}]\Big{)},$ (4.32)
+$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}\leq\ell^{(r)}_{1})]$
+$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,1})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{1}]\Big{)},$ (4.33)
+$k\mathbb{E}[(R^{(r)}_{e})^{k-1}(1-1(R^{(r)}_{e}\geq 1/r))1(L^{(r)}>\ell^{(r)}_{1})]$
+$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{1})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{1}]\Big{)}.$ (4.34)
 Proof.
-For $k\geq 1$, we first prove (4.5). We apply $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
-$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,1\leq L^{(r)}(t)\leq\ell^{(r)}_{0}).$
-Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ and $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, respectively.
+For $k\geq 1$, we first prove (4.5). We apply $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{1})$ to (4.1), then we have (4.5) because
+$\frac{d}{dt}f(X^{(r)}(t))=k(R^{(r)}_{s}(t))^{k-1}1(R^{(r)}_{s}<1/r,1\leq L^{(r)}(t)\leq\ell^{(r)}_{1}).$
+Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z>\ell^{(r)}_{1})$, $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{1})$ and $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z>\ell^{(r)}_{1})$, respectively.
 ∎
@@
 Corollary 4.4.
-For each $k=1,2,3$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
+For each $k=1,2,3$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{1})]$ are uniformly bounded for $r\in(0,1]$.
 Note that the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{2}(\theta_{2})$ should be obtained from (4.3).
@@
 Define, for $i=1,2$,
-$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}].$ (4.34)
+$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad+\lambda^{(r)}_{i}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}].$ (4.35)
 Then, $\Delta^{(r)}_{i}$ is well defined and finite by Corollary 4.4, and we have
 Lemma 4.6.
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.35)
-$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]+o(r).$ (4.36)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{1}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.36)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\mathbb{P}[L^{(r)}>\ell^{(r)}_{1}]+o(r).$ (4.37)
 Proof.
 For $k=1$, multiply both sides of (4.5) and (4.5) by $\mu^{(r)}$ and $\lambda^{(r)}_{1}$, respectively, and take their difference, then, noting that $c^{(r)}_{1}\mu^{(r)}-\lambda^{(r)}_{1}=r\mu^{(r)}b_{1}+o(r)$, we have
-$(r\mu^{(r)}b_{1}+o(r))\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]-c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]$
-$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
-$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
-because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r)=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r)=o(r^{2})$ by (4.29).
+$(r\mu^{(r)}b_{1}+o(r))\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{1}]-c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]$
+$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]$
+$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{1})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1}]\Big{)}+o(r),$
+because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{1}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{1}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r)=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r)=o(r^{2})$ by (4.30).
 This proves (4.36).
 Similarly, (4.37) is obtained from (4.5) and (4.5), ∎
-We will use Lemma 4.6 to determine the limits of $\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]$ and $\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]$, equivalently, $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, as $r\downarrow 0$ under some conditions.
+We will use Lemma 4.6 to determine the limits of $\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{1}]$ and $\mathbb{P}[L^{(r)}>\ell^{(r)}_{1}]$, equivalently, $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, as $r\downarrow 0$ under some conditions.
 (Step 2)   In this step, we expand $D^{(r)}(\theta,0)$ and $D^{(r)}(0,\theta)$ concerning $r$ as $r\downarrow 0$.
@@
 For each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$,
-$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}(1+o(\theta_{i}r)),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r+o(\theta_{i}r)),\quad i=1,2,$ (4.37)
-$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)),$ (4.38)
-$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}$
-$\quad=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)).$ (4.39)
+$e^{r\theta_{i}\ell^{(r)}_{1}}=e^{\theta_{i}\ell_{1}}(1+o(\theta_{i}r)),\quad e^{r\theta_{i}(\ell^{(r)}_{1}+1)}=e^{\theta_{i}\ell_{1}}(1+\theta_{i}r+o(\theta_{i}r)),\quad i=1,2,$ (4.38)
+$e^{r\theta_{1}\ell^{(r)}_{1}}-e^{r\theta_{2}\ell^{(r)}_{1}}=e^{\theta_{1}\ell_{1}}-e^{\theta_{2}\ell_{1}}+e^{(\theta_{1}\vee\theta_{2})\ell_{1}}(o(\theta_{1}r)+o(\theta_{2}r)),$ (4.39)
+$e^{r\theta_{1}(\ell^{(r)}_{1}+1)}-e^{r\theta_{2}(\ell^{(r)}_{1}+1)}$
+$\quad=e^{\theta_{1}\ell_{1}}-e^{\theta_{2}\ell_{1}}+r(e^{\theta_{1}\ell_{1}}\theta_{1}-e^{\theta_{2}\ell_{1}}\theta_{2})+e^{(\theta_{1}\vee\theta_{2})\ell_{1}}(o(\theta_{1}r)+o(\theta_{2}r)).$ (4.40)
 Lemma 4.8.
@@
 $\qquad+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}).$ (4.44)
-These lemmas are proved in Appendix A.4.
-We first compute $D^{(r)}(\theta,0)$.
-$D^{(r)}(\theta,0)$ $=\alpha^{(r)}_{e}(1-e^{r\theta\ell^{(r)}_{0}})\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
-$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(1-e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\qquad+\alpha^{(r)}_{e}(e^{r\theta(\ell^{(r)}_{0}+1)}-1)\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-1)1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$=-\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.44)
-Define ${\mathcal{E}}^{(r)}_{1}$ is defined as
-${\mathcal{E}}^{(r)}_{1}$ $=2^{-1}\mathbb{E}^{(r)}_{e}[\{(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.45)
+These lemmas are proved in Appendix A.5. We first compute $D^{(r)}(\theta,0)$.
+$D^{(r)}(\theta,0)$ $=\alpha^{(r)}_{e}(1-e^{r\theta\ell^{(r)}_{1}})\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]$
+$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[(1-e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\qquad+\alpha^{(r)}_{e}(e^{r\theta(\ell^{(r)}_{1}+1)}-1)\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-1)1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$=-\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.45)
+Define ${\mathcal{E}}^{(r)}_{1}$ as
+${\mathcal{E}}^{(r)}_{1}$ $=2^{-1}\mathbb{E}^{(r)}_{e}[\{(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.46)
 Applying (4.8) and
@@
 to (4.3) and recalling the definition of $\Delta^{(r)}_{1}$, we have
-$D^{(r)}(\theta,0)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(-r\theta\Delta^{(r)}_{1}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{1})+o((r\theta)^{2}),$ (4.46)
-because $\mathbb{E}^{(r)}_{s}[S_{3}(\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
+$D^{(r)}(\theta,0)=\alpha^{(r)}_{e}e^{\theta\ell_{1}}(-r\theta\Delta^{(r)}_{1}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{1})+o((r\theta)^{2}),$ (4.47)
+because $\mathbb{E}^{(r)}_{s}[S_{3}(\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{1})]$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
 Similarly, define ${\mathcal{E}}^{(r)}_{2}$ as
-${\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{e}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.47)
+${\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{e}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.48)
 Then, $D^{(r)}(0,\theta)$ is computed from (4.2) as
-$D^{(r)}(0,\theta)=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}-1)\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
-$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
-$\quad+\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(1-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}$
-$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[1+r\theta\mu^{(r)}\widehat{R}^{(r)}_{s-}-2^{-1}(r\theta)^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
+$D^{(r)}(0,\theta)=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{1}}-1)\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
+$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{1})]+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{1}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{1}]$
+$\quad+\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(1-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}\big{\{}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]\big{\}}$
+$=\alpha^{(r)}_{e}e^{\theta\ell_{1}}\big{\{}\mathbb{E}^{(r)}_{e}[1+r\theta\mu^{(r)}\widehat{R}^{(r)}_{s-}$
+$\qquad-2^{-1}(r\theta)^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
 $\quad-(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}[1-r\theta(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})$
-$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}]1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}+o((r\theta)^{2})$
-$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r\theta\Delta^{(r)}_{2}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{2})+o((r\theta)^{2}).$ (4.48)
+$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}]1(L^{(r)}(0)=\ell^{(r)}_{1})]\big{\}}+o((r\theta)^{2})$
+$=\alpha^{(r)}_{e}e^{\theta\ell_{1}}(r\theta\Delta^{(r)}_{2}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{2})+o((r\theta)^{2}).$ (4.49)
 (Step 3)   In this step, we derive asymptotic expansions for $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$. First, note that (4.36) and (4.37) can be written as
@@
 Then, substituting (4.47) into the asymptotic BAR (4.3) with $\theta_{2}=0$ and applying (4.50), we have
-$\frac{1}{2}rc_{1}\mu\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-r\theta_{1}{\mathcal{E}}^{(r)}_{1})+\theta_{1}o(r),$
-$\quad=c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0]+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+\theta_{1}o(r),$ (4.51)
+$\frac{1}{2}rc_{1}\mu\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-e^{\theta_{1}\ell_{1}}\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-r\theta_{1}{\mathcal{E}}^{(r)}_{1})+\theta_{1}o(r),$
+$\quad=c_{1}\mu(1-e^{\theta_{1}\ell_{1}})\mathbb{P}[L^{(r)}=0]+re^{\theta_{1}\ell_{1}}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+\theta_{1}o(r),$ (4.52)
 Similarly, substituting (4.3) into (4.3) with $\theta_{1}=0$ and applying (4.51), we have
-$\frac{1}{2}rc_{2}\mu\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\big{(}\Delta^{(r)}_{2}+r\theta_{2}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r)$
-$\quad=re^{\theta_{2}\ell_{0}}\big{(}\mu b_{2}\varphi^{(r)}_{2}(0)+\theta_{2}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r).$ (4.52)
+$\frac{1}{2}rc_{2}\mu\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{1}}\big{(}\Delta^{(r)}_{2}+r\theta_{2}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r)$
+$\quad=re^{\theta_{2}\ell_{1}}\big{(}\mu b_{2}\varphi^{(r)}_{2}(0)+\theta_{2}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r).$ (4.53)
 We next compute $\mathbb{P}[L^{(r)}=0]$ for (4.3).
 Lemma 4.9.
-$\mathbb{P}[L^{(r)}=0]=\begin{cases}\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b=0,\\ \frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),&b\not=0.\end{cases}$ (4.53)
+$\mathbb{P}[L^{(r)}=0]=\begin{cases}\frac{r}{2c_{1}\mu\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b=0,\\ \frac{re^{\beta_{1}\ell_{1}}}{c_{1}\mu(e^{\beta_{1}\ell_{1}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),&b\not=0.\end{cases}$ (4.54)
 Proof.
 From (4.3), we have
-$\mathbb{P}[L^{(r)}=0]$ $=r\frac{c_{1}\mu\sigma_{1}^{2}\big{(}\theta_{1}-\beta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}(b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r).$ (4.54)
+$\mathbb{P}[L^{(r)}=0]$ $=r\frac{c_{1}\mu\sigma_{1}^{2}\big{(}\theta_{1}-\beta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{1}}(b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{1}}-1)}+o(r).$ (4.55)
 If $b_{1}=0$, then $\beta_{1}=b_{1}=0$, and therefore (4.55) yields
-$\mathbb{P}[L^{(r)}=0]$ $=r\lim_{\theta_{1}\to 0}\frac{\theta_{1}(c_{1}\mu\sigma_{1}^{2}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r)$
-$=\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),$
+$\mathbb{P}[L^{(r)}=0]$ $=r\lim_{\theta_{1}\to 0}\frac{\theta_{1}(c_{1}\mu\sigma_{1}^{2}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{1}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{1}}-1)}+o(r)$
+$=\frac{r}{2c_{1}\mu\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),$
 which proves (4.54) for $b_{1}=0$. If $b_{1}\not=0$, then substituting $\theta_{1}=\beta_{1}$ into (4.55) yields
-$\mathbb{P}[L^{(r)}=0]=\frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),$
+$\mathbb{P}[L^{(r)}=0]=\frac{re^{\beta_{1}\ell_{1}}}{c_{1}\mu(e^{\beta_{1}\ell_{1}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),$
 which is (4.54) for $b_{1}\not=0$. ∎
@@
 Lemma 4.10.
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}\Big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\frac{1}{c_{1}\mu}\beta_{1}e^{\beta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\Big{)}+o(r),&b\not=0.\end{cases}$ (4.55)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{1}}-1}\Big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\frac{1}{c_{1}\mu}\beta_{1}e^{\beta_{1}\ell_{1}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\Big{)}+o(r),&b\not=0.\end{cases}$ (4.56)
 This lemma is immediate from substituting $\mathbb{P}^{(r)}_{e}[L^{(r)}=0]$ of Lemma 4.9 into (4.50).
@@
 Lemma 4.11.
-$\psi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}\not=0,\end{cases}$ (4.56)
-$\psi^{(r)}_{2}(\theta_{2})$ $=\frac{\beta_{2}e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\varphi^{(r)}_{2}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}e^{\theta_{2}\ell_{0}}+o(1).$ (4.57)
+$\psi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{1}}-1}{\ell_{1}\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}}+o(1),&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}}+o(1),&b_{1}\not=0,\end{cases}$ (4.57)
+$\psi^{(r)}_{2}(\theta_{2})$ $=\frac{\beta_{2}e^{\theta_{2}\ell_{1}}}{\beta_{2}-\theta_{2}}\Big{(}\varphi^{(r)}_{2}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}e^{\theta_{2}\ell_{1}}+o(1).$ (4.58)
 Proof.
@@
 $-rc_{1}\mu\sigma_{1}^{2}\theta_{1}\psi^{(r)}(\theta_{1})$
-$\quad=(1-e^{\theta_{1}\ell_{0}})\frac{r}{\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+r2e^{\theta_{1}\ell_{0}}\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r),$
+$\quad=(1-e^{\theta_{1}\ell_{1}})\frac{r}{\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+r2e^{\theta_{1}\ell_{1}}\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r),$
 which proves (4.57) for $b_{1}=0$. Next assume that $b_{1}\not=0$. Similarly to the case of $b_{1}=0$, it follows from Lemma 4.9 and (4.3) for $b_{1}\not=0$ that
 $\frac{r}{2}c_{1}\mu\sigma_{1}^{2}(\beta_{1}-\theta_{1})\psi^{(r)}(\theta_{1})$
-$\quad=\frac{re^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})}{(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r)$
-$\quad=\frac{rb_{1}\mu\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\big{(}e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+r\frac{\beta_{1}e^{\beta_{1}\ell_{0}}(1-e^{\theta_{1}\ell_{0}})+\theta_{1}e^{\theta_{1}\ell_{0}}(e^{\beta_{1}\ell_{0}}-1)}{(e^{\beta_{1}\ell_{0}}-1)}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r).$
+$\quad=\frac{re^{\beta_{1}\ell_{1}}(1-e^{\theta_{1}\ell_{1}})}{(e^{\beta_{1}\ell_{1}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+re^{\theta_{1}\ell_{1}}\big{(}b_{1}\mu\varphi^{(r)}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r)$
+$\quad=\frac{rb_{1}\mu\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{1}}-1}\big{(}e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}})+r\frac{\beta_{1}e^{\beta_{1}\ell_{1}}(1-e^{\theta_{1}\ell_{1}})+\theta_{1}e^{\theta_{1}\ell_{1}}(e^{\beta_{1}\ell_{1}}-1)}{(e^{\beta_{1}\ell_{1}}-1)}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r).$
 Hence, we have
-$\psi^{(r)}_{1}(\theta_{1})$ $=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}-\theta_{1}e^{\theta_{1}\ell_{0}}+(\theta_{1}-\beta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
-$=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}$
-$\quad+\frac{\beta_{1}(e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}})+(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}-(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{0}}e^{\beta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
-$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2e^{\theta_{1}\ell_{0}}}{c_{1}\mu\sigma_{1}^{2}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
+$\psi^{(r)}_{1}(\theta_{1})$ $=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}$
+$\quad+\frac{\beta_{1}e^{\beta_{1}\ell_{1}}-\theta_{1}e^{\theta_{1}\ell_{1}}+(\theta_{1}-\beta_{1})e^{\theta_{1}\ell_{1}}e^{\beta_{1}\ell_{1}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
+$=\frac{\beta_{1}\varphi^{(r)}(0)}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}$
+$\quad+\frac{\beta_{1}(e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}})+(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{1}}-(\beta_{1}-\theta_{1})e^{\theta_{1}\ell_{1}}e^{\beta_{1}\ell_{1}}}{c_{1}\mu\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)(\beta_{1}-\theta_{1})}2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
+$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2e^{\theta_{1}\ell_{1}}}{c_{1}\mu\sigma_{1}^{2}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}+o(r)$
 Thus, we have (4.57) for $b_{1}\not=0$. We next compute $\psi^{(r)}_{2}(\theta_{2})$. From (4.3), we have
-$\psi^{(r)}_{2}(\theta_{2})=\frac{e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\beta_{2}\varphi^{(r)}_{2}(0)+(\beta_{2}-(\beta_{2}-\theta_{2}))\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}+\theta_{2}o(r),$
+$\psi^{(r)}_{2}(\theta_{2})=\frac{e^{\theta_{2}\ell_{1}}}{\beta_{2}-\theta_{2}}\Big{(}\beta_{2}\varphi^{(r)}_{2}(0)+(\beta_{2}-(\beta_{2}-\theta_{2}))\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}+\theta_{2}o(r),$
 which proves (4.58). ∎
@@
 This suggests that (5.1) must be a key condition for Theorem 3.1.
-Taking this into account, we separate the proof of Theorem 3.1 in two steps by two propositions.
+Taking this into account, we separate the proof of Theorem 3.1 in two steps.
 Proposition 5.1.
-In Theorem 3.1, the assumption (3.f) can be replaced by (5.1).
+If (5.1) holds, then Theorem 3.1 holds without the assumption (3.f).
 Proposition 5.2.
-Under the assumptions (3.a)–(3.e), (5.1) holds if
-$\lim_{r\downarrow 0}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})=0.$ (5.2)
-Furthermore, (5.2) holds if (3.f) holds. Hence, (3.f) implies (5.1).
-In the next two subsections, we prove these propositions. For their proofs, we will use the following lemmas, which are proved in Appendix A.5, respectively.
+Under the assumptions (3.a)–(3.e), if
+$\lim_{r\downarrow 0}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})=0,$ (5.2)
+then (5.1) holds. In particular, (3.f) implies (5.2) and therefore (5.1).
+These propositions are proved in the next two subsections. Obviously, they prove Theorem 3.1. To prove Proposition 5.2, we will use the following lemma, which is proved in Appendix A.6.
 Lemma 5.1.
-Under the assumptions (3.a)–(3.e), if
-$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,$ (5.3)
-then
-$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.4)
+Under the assumptions (3.a)–(3.e), we have
+$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{i}1(L^{(r)}(0)=\ell^{(r)}_{1})]=0,\qquad i=1,2,$ (5.3)
+if either $\lambda_{2}\not=\lambda_{2}$ or
+$\lim_{r\downarrow 0}\mathbb{P}^{(r)}[L^{(r)}=\ell^{(r)}_{1}]=0.$ (5.4)
 Lemma 5.2.
-Under the assumptions (3.a)–(3.e), if (5.4) holds, then
-$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0.$ (5.5)
+Under the assumptions (3.a)–(3.e), (5.3) implies (5.2).
 Remark 5.1.
-From the proof of Proposition 5.2, we can see that (5.4) and (5.5) imply (5.2), and therefore (5.3) is sufficient for (5.2) and (5.1).
-We think that (5.3) is likely true because $L^{(r)}(\cdot)$ would weakly converge to a reflected diffusion and the stationary distribution of the reflected diffusion has a density.
-However, the verification of (5.3) would need a process limit for $L^{(r)}(\cdot)$, so this verification is beyond the present framework.
+By Lemmas 5.1 and 5.2 and Proposition 5.2, (5.4) is sufficient for (5.1).
+We conjecture that (5.4) is true because $L^{(r)}(\cdot)$ would weakly converge to a reflected diffusion whose stationary distribution has a density.
+However, the verification of (5.4) would need a process limit for $L^{(r)}(\cdot)$, so it is beyond the present framework.
 5.1 Proof of Proposition 5.1
-Throughout this proof, the part (i), (ii) and (iii) of Theorem 3.1 under the condition (5.1) instead of (3.f) are denoted by the part (i)’, (ii)’ and (iii)’, respectively.
-We first prove the part (i)’. Since the condition in the part (i)’ is given by (3.6), Corollary 4.3 implies that
+Throughout this proof, the parts (i), (ii) and (iii) of Theorem 3.1 under the condition (5.1) instead of (3.f) are denoted by the part (i)’, (ii)’ and (iii)’, respectively.
+We first prove the part (i)’. Since (3.6) holds, Corollary 4.3 implies that
 $\lim_{r\downarrow 0}|\psi^{(r)}_{i}(\theta_{i})/\varphi^{(r)}_{i}(0)-\varphi^{(r)}_{i|cd}(\theta_{i})|=0.$ (5.5)
 By this formula, we can immediately see that Lemma 4.11 and (5.1) yield
-$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (5.7)
-$\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{0}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (5.8)
+$\lim_{r\downarrow 0}\varphi^{(r)}_{1|cd}(\theta_{1})=\widetilde{\varphi}_{1|cd}(\theta_{1})\equiv\begin{cases}\displaystyle\frac{e^{\theta_{1}\ell_{1}}-1}{\ell_{1}\theta_{1}},&b_{1}=0,\\ \displaystyle\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}},\quad&b_{1}\not=0,\end{cases}$ (5.6)
+$\lim_{r\downarrow 0}\varphi^{(r)}_{2|cd}(\theta_{2})=\widetilde{\varphi}_{2|cd}(\theta_{2})\equiv e^{\theta_{2}\ell_{1}}\frac{\beta_{2}}{\beta_{2}-\theta_{2}}.$ (5.7)
 Since $\widetilde{\varphi}_{i|cd}$ is the moment generating function of $\widetilde{\nu}_{i|cd}$, (5.6) and (5.7) are equivalent to that $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$ for $i=1,2$, respectively.
@@
 First consider the case that $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,t]$.
-In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$, so, by Lemma 4.4 for $k=1$,
-$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)\;=\ell^{(r)}_{0})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]+o(r).$
+In this case, $R^{(r)}_{s}(0-)$ has the same distribution as $T^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{e}$ by Lemma 2.1 and Theorem 3.1 of [12], so, by Lemma 4.4 for $k=1$,
+$\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)\;=\ell^{(r)}_{1})]$ $\;=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]+o(r).$
 Applying this fact to the definitions of $\Delta^{(r)}_{i}$ yields
 $\lambda^{(r)}_{2}\Delta^{(r)}_{1}-\lambda^{(r)}_{1}\Delta^{(r)}_{2}$
-$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])=o(r).$ (5.9)
+$\quad=(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})(\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{1})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}])=o(r).$ (5.8)
 Compute $\alpha^{(r)}_{e}\Delta^{(r)}_{1}$ by Lemma 4.10 and (5.1), and recall (4.51), then
-$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}b_{1}\mu\varphi^{(r)}_{1}(0)+o(r),&b\not=0,\\ \end{cases}$ (5.10)
+$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{1}}c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{1}}-1}b_{1}\mu\varphi^{(r)}_{1}(0)+o(r),&b\not=0,\\ \end{cases}$ (5.9)
 $\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (5.10)
 Then, for $b_{1}=0$, (5.1) implies that
-$\lambda_{2}\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
+$\lambda_{2}\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
 Since $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$ and $\mu c_{i}=\lambda_{i}$ for $i=1,2$, this can be written as
-$\frac{\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$ (5.12)
+$\frac{\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(r)}_{1}(0)=\frac{b_{2}}{c_{2}}(1-\varphi^{(r)}_{1}(0))+o(1).$ (5.11)
 Hence,
-$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}\sigma_{1}^{2}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.13)
+$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{2b_{2}\ell_{1}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{1}}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}\sigma_{1}^{2}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{1}}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.12)
 Similarly, for $b_{1}\not=0$, from (5.9) and (5.10), we have
-$\frac{\lambda_{2}b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$ (5.14)
+$\frac{\lambda_{2}b_{1}\mu}{e^{\beta_{1}\ell_{1}}-1}\varphi^{(r)}_{1}(0)-\lambda_{1}\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$ (5.13)
 Hence, $\mu c_{i}=\lambda_{i}$ for $i=1,2$ yields
-$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.15)
+$\lim_{r\downarrow 0}\varphi^{(r)}_{i}(0)=\widetilde{\varphi}_{i}(0)\equiv\begin{cases}\displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}=\widetilde{A}_{1}>0,&i=1,\\ \displaystyle\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}=\widetilde{A}_{2}>0,&i=2.\end{cases}$ (5.14)
 We next consider the case that $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$. Define $\overline{\varphi}_{1}(0)$ and $\underline{\varphi}_{1}(0)$ as
@@
 To see this, we first consider $\overline{\varphi}_{1}(0)$. Since $\varphi^{(\overline{r}_{n})}_{1}(0)$ converges to $\overline{\varphi}_{1}(0)$, applying (5.1) and (4.18) to Lemma 4.11, we have
-$\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(\theta_{1})=\widetilde{\varphi}_{1}(\theta_{1})\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(\theta_{1})\overline{\varphi}_{1}(0).$ (5.16)
-Since the distribution $\widetilde{\nu}_{1}$ has the bounded density $h_{1|cd}$ on $[0,\ell_{0}]$, this implies that
-$\lim_{n\to\infty}\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]\leq\lim_{n\to\infty}\mathbb{P}[\overline{r}_{n}L^{(\overline{r}_{n})}\in[\ell_{0}-\varepsilon,\overline{r}_{n}\ell^{(\overline{r}_{n})}_{0}]]$
+$\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(\theta_{1})=\lim_{n\to\infty}\psi^{(\overline{r}_{n})}_{1}(\theta_{1})=\widetilde{\varphi}_{1}(\theta_{1})\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(\theta_{1})\overline{\varphi}_{1}(0).$ (5.15)
+Since the distribution $\widetilde{\nu}_{1}$ has the bounded density $h_{1|cd}$ on $[0,\ell_{1}]$, this implies that
+$\lim_{n\to\infty}\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]\leq\lim_{n\to\infty}\mathbb{P}\left[\overline{r}_{n}L^{(\overline{r}_{n})}\in[\ell_{1}-\varepsilon,\overline{r}_{n}\ell^{(\overline{r}_{n})}_{0}]\right]$
 […]
@@
 If $b_{1}=0$, then (4.51) and Lemma 4.10 yield
-$\alpha^{(\overline{r}_{n})}_{e}(\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2})=\frac{rc_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)-r\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)+o(\overline{r}_{n}),$
+$\alpha^{(\overline{r}_{n})}_{e}(\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2})=\frac{rc_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(\overline{r}_{n})}_{1}(0)-r\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)+o(\overline{r}_{n}),$
 Hence, by (5.18) and $\varphi^{(\overline{r}_{n})}_{1}(0)+\varphi^{(\overline{r}_{n})}_{2}(0)=1$,
-$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)=\mu b_{2}(1-\varphi^{(\overline{r}_{n})}_{1}(0))+o(1).$
+$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(\overline{r}_{n})}_{1}(0)=\mu b_{2}(1-\varphi^{(\overline{r}_{n})}_{1}(0))+o(1).$
 This is identical with (5.11) for $r=\overline{r}_{n}$ because $c_{1}=c_{2}$ by $\lambda_{1}=\lambda_{2}$.
@@
 Similarly, if $b_{1}\not=0$, then, again by (4.51), (5.10) and (5.18) for $r=\overline{r}_{n}$, we have
-$\frac{b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(\overline{r}_{n})}_{1}(0)-\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)=o(1).$
+$\frac{b_{1}\mu}{e^{\beta_{1}\ell_{1}}-1}\varphi^{(\overline{r}_{n})}_{1}(0)-\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)=o(1).$
 This is identical with (5.13) for $r=\overline{r}_{n}$ by $\lambda_{1}=\lambda_{2}$. Hence, we have (5.14) for $r=\overline{r}_{n}$ and $b_{1}\not=0$. Thus, we have
@@
 Finally, we prove the part (iii)’. Assume that $\lambda_{1}>\lambda_{2}$, then $\Delta^{(r)}_{1}>\Delta^{(r)}_{2}$ by (5.18). For $b=0$, this yields
-$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
+$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{1}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
 and therefore $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$.
@@
 5.2 Proof of Proposition 5.2
-Using (5.2), we prove (5.1) by reductio ad absurdum.
+We prove that (5.2) implies (5.1) by reductio ad absurdum.
 First, assume that ${\mathcal{E}}^{(r)}_{1}$ does not vanish as $r\downarrow 0$.
@@
 Then, by Lemma 4.11, we have
-$\lim_{n\to\infty}\varphi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}\not=0.\end{cases}$
-This shows that $\nu^{(r)}_{1}$ weakly converges to a finite measure on $[0,\ell_{0}]$, but this measure cannot have a negative mass at $\ell_{0}$, so we must have that $\delta_{1}<0$ because $\delta_{1}\not=0$.
+$\lim_{n\to\infty}\varphi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{1}}-1}{\ell_{1}\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}},&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}},&b_{1}\not=0.\end{cases}$
+This shows that $\nu^{(r)}_{1}$ weakly converges to a finite measure on $[0,\ell_{1}]$, but this measure cannot have a negative mass at $\ell_{1}$, so we must have that $\delta_{1}<0$ because $\delta_{1}\not=0$.
 We next consider the sequence $\{{\mathcal{E}}^{(r^{\prime}_{n})}_{2};n\geq 1\}$.
@@
 We now prove that (3.f) implies (5.2). By the definitions of ${\mathcal{E}}^{(r)}_{1}$ and ${\mathcal{E}}^{(r)}_{2}$, we have
-${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
-$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.20)
+${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}=-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
+$\qquad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (5.19)
 We separately consider the two cases that $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\not=\lambda_{2}$.
@@
 Hence, (5.2) implies (5.2).
-Next consider the case that $\lambda_{1}\not=\lambda_{2}$. In this case, we consider the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$. From (4.2), we have
-$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.21)
-Applying (4.37), (4.8) and Corollary 4.4 to this equation, we have
-$D^{(r)}(\theta,\theta)$ $=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}\big{[}(r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}$
-$\quad+2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}\lambda^{(r)}_{2}-(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}\lambda^{(r)}_{1}\}\widehat{R}^{(r)}_{e}$
-$\quad-2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2})^{2}-(\lambda^{(r)}_{1})^{2}\}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+o((r\theta)^{2}),$
-$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+(r\theta)^{2}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})\big{\}}+o((r\theta)^{2}).$
-This implies that
-$D^{(r)}(\theta,\theta)$ $=r\theta\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+O((r\theta)^{2}).$ (5.22)
-Substitute this formula into the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$, which is given by
-$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta)$
-$\quad=c_{1}\left(\mu r\theta-\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o((\theta r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta).$ (5.23)
-Then, picking up the terms of order $r\theta$, we have
-$c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r\theta).$
-Since $\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=O(r)$ by Corollary 4.2 and Lemma 4.9, this yields
-$(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r).$ (5.24)
-Since $\lambda_{1}\not=\lambda_{2}$ by (f.1), (5.24) implies (5.4), which further implies (5.5). Hence,
-$\lim_{r\downarrow 0}\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,$ (5.25)
-and therefore we have (5.2) by (5.2) and Lemma 5.2. Thus, the proof of Proposition 5.2 is completed.
+Next consider the case that $\lambda_{1}\not=\lambda_{2}$.
+In this case, we have (5.3) by Lemma 5.1.
+Applying (5.3) to (5.2), we have (5.2) because $\lambda^{(r)}_{i}$ and $\sigma^{(r)}_{e,i}$ converge to constants $\lambda_{i}$ and $\sigma_{e,i}$, respectively, as $r\downarrow 0$ for $i=1,2$.
+This completes the proof of Proposition 5.2.
 6 Concluding remarks
@@
 Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic.
-From the assumptions (3.a)–(3.f), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
+From the assumptions (3.a)–(3.f), one can guess this limiting distribution must be the stationary distribution of the reflected diffusion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{1}$.
 Recall that, if one of the conditions in (ii) of Theorem 3.1 is satisfied, then, for $b_{1}\not=0$, the stationary distribution $\widetilde{\nu}$ is characterized by the stationary equation, that is, BAR:
-$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
-$\qquad+\beta_{2}e^{\theta_{2}\ell_{0}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (6.1)
-which can be obtained from (4.3) and (4.2) similarly to $D^{(r)}(\theta,\theta)$ from (5.22) and (5.2), recall that $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
+$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\frac{\beta_{1}e^{\beta_{1}\ell_{1}}}{e^{\beta_{1}\ell_{1}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)-\frac{\beta_{1}e^{\theta_{1}\ell_{1}}}{e^{\beta_{1}\ell_{1}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
+$\qquad+\beta_{2}e^{\theta_{2}\ell_{1}}\theta_{2}\frac{c_{2}\sigma^{2}_{2}}{2}\widetilde{\varphi}_{2}(0)=0,\qquad\theta\equiv(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}_{-},$ (6.1)
+which can be obtained from (4.3) and (4.2) similarly to $D^{(r)}(\theta,\theta)$ from (A.39) and (A.6), recall that $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$.
 The BAR for $b_{1}=0$ is obtained from (6.1) by letting $b_{1}\downarrow 0$.
@@
 To this end, define real-valued functions $\sigma(x)$ and $b(x)$ for $x\in\mathbb{R}_{+}$ as
-$b(x)=b_{1}\mu 1(x\leq\ell_{0})+b_{2}\mu 1(x>\ell_{0}),$
-$\sigma(x)=\sqrt{c_{1}}\mu\sigma_{1}1(x\leq\ell_{0})+\sqrt{c_{2}}\mu\sigma_{2}1(x>\ell_{0}),$
+$b(x)=b_{1}\mu 1(x\leq\ell_{1})+b_{2}\mu 1(x>\ell_{1}),$
+$\sigma(x)=\sqrt{c_{1}}\mu\sigma_{1}1(x\leq\ell_{1})+\sqrt{c_{2}}\mu\sigma_{2}1(x>\ell_{1}),$
 then one may expect that $Z(\cdot)$ satisfies the following stochastic integral equation.
@@
 Assume that $Z(\cdot)$ is stationary, then, applying the Ito formula to (6.2) for test function $f(x)=e^{\theta x}$, we have
-$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\leq 0.$ (6.3)
-where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$.
+$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{1}}}{e^{\beta_{1}\ell_{1}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\leq 0.$ (6.3)
+where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{1})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{1})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$.
 (6.3) is the stationary equation for $Z(\cdot)$.
@@
 First assume that $b_{1}\not=0$.
-Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{0}$ from below or above.
-Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{0}$ at time $\tau^{(r)}_{2}$ after time $B^{(r)}_{2}$.
-Since the relative frequency of moving up at level $\ell^{(r)}_{0}$ is guessed to be $f^{(r)}_{\rm up}\equiv\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, the mean sojourn time for $L^{(r)}(t)$ to be above $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
+Under the stationary setting, let $\tau^{(r)}_{1}$ be the first time for $L^{(r)}(t)$ to get into $\ell^{(r)}_{1}$ from below or above.
+Suppose that $L^{(r)}(t)$ first moves to $\ell^{(r)}_{1}+1$ after time $\tau^{(r)}_{1}$, and returns to $\ell^{(r)}_{1}$ at time $\tau^{(r)}_{2}$ after time $B^{(r)}_{2}$.
+Since the relative frequency of moving up at level $\ell^{(r)}_{1}$ is guessed to be $f^{(r)}_{\rm up}\equiv\lambda^{(r)}_{2}/(\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)})$ and $B^{(r)}_{2}$ is stochastically almost identical with the busy period of the $GI/G/1$ queue with inter-arrival time $T^{(r)}_{e,2}$ and service time $T^{(r)}_{s}/c^{(r)}_{2}$, the mean sojourn time for $L^{(r)}(t)$ to be above $\ell^{(r)}_{1}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
 $f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]\sim\frac{\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{c^{(r)}_{2}\mu^{(r)}(1-\rho^{(r)}_{2})}=\frac{c_{2}}{c_{1}+c_{2}}\times\frac{1}{rb_{2}\mu}+o(1),$ (6.4)
 where $\sim$” stands for the same order as $r\downarrow 0$, that is, for non-zero valued functions $f_{1}$ and $f_{2}$, $f_{1}(r)\sim f_{2}(r)$ means that $\lim_{r\downarrow 0}f_{1}(r)/f_{2}(r)=1$.
-On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{0}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{0}$ from below.
-Let $f^{(r)}_{\rm down}=1-f^{(r)}_{\rm up}$, then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that the mean sojourn time for $L^{(r)}(t)$ to be below $\ell^{(r)}_{0}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
-$f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]$ $\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]}$
-$=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{0}}-1}{rb_{1}\mu}+o(1).$ (6.5)
-Since $f^{(r)}_{\rm up}B^{(r)}_{2}+f^{(r)}_{\rm down}B^{(r)}_{1}$ is the mean cycle time to return to $\ell^{(r)}_{0}$, it follows from the cycle formula of a regenerative process that
-$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]}{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]+f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}=\widetilde{A}_{1}.$
+On the other hand, when $L^{(r)}(t)$ first moves to $\ell^{(r)}_{1}-1$ after time $\tau^{(r)}_{1}$, let $B^{(r)}_{1}$ be the time to returns to $\ell^{(r)}_{1}$ from below.
+Let $f^{(r)}_{\rm down}=1-f^{(r)}_{\rm up}$, then, similarly to $B^{(r)}_{2}$ but, taking $b_{1}\not=0$ and the reflecting boundary at level 00 into account, it follows from (4.25) of [12] that the mean sojourn time for $L^{(r)}(t)$ to be below $\ell^{(r)}_{1}$ in the cycle $(\tau^{(r)}_{1},\tau^{(r)}_{2}]$ is
+$f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]$ $\sim\frac{c^{(r)}_{1}\mu^{(r)}}{\lambda^{(r)}_{2}+c^{(r)}_{1}\mu^{(r)}}\times\frac{1}{\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{1}]}$
+$=\frac{c_{1}}{c_{1}+c_{2}}\times\frac{e^{\beta_{1}\ell_{1}}-1}{rb_{1}\mu}+o(1).$ (6.5)
+Since $f^{(r)}_{\rm up}B^{(r)}_{2}+f^{(r)}_{\rm down}B^{(r)}_{1}$ is the mean cycle time to return to $\ell^{(r)}_{1}$, it follows from the cycle formula of a regenerative process that
+$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})=\lim_{r\downarrow 0}\frac{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]}{f^{(r)}_{\rm down}\mathbb{E}[B^{(r)}_{1}]+f^{(r)}_{\rm up}\mathbb{E}[B^{(r)}_{2}]}=\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}=\widetilde{A}_{1}.$
 Thus, for $b_{1}\not=0$, (3.13) is intuitively obtained for $i=1$, and it is similarly obtained for $i=2$. The case for $b_{1}=0$ is similarly derived. Thus, we have support of (3.13).
 Appendix
-A.1 Proofs of Corollary 3.1
+A.1 Proof of Lemma 3.1
+Let $N^{(r)}_{e,\wedge}(\cdot)$ and $N^{(r)}_{e,\vee}(\cdot)$ be the renewal processes whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ and $T_{e,1}(n)\vee T_{e,2}(n)$, respectively, for $n\geq 1$.
+Obviously, $0<1/\mathbb{E}[T_{e,1}\vee T_{e,2}]=\mathbb{E}[N^{(r)}_{e,\vee}(t)]\leq\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,\wedge}(t)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$ under the stationary framework.
+Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]$ is finite and positive.
+To prove $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$, consider the stationary process $X^{(r)}_{a}(t)\equiv L^{(r)}(t)\wedge a$ for each constant $a>0$.
+Then, $X^{(r)}_{a}(t)$ is piecewise constant, has the increment $X^{(r)}_{a}(u)-X^{(r)}_{a}(u-)=1(L^{(r)}(u-)\leq a-1)$ at the counting instant $u$ of $N^{(r)}_{e}$, and has the increment $X^{(r)}_{a}(u)-X^{(r)}_{a}(u-)=-1(L^{(r)}(u)\leq a)$ at the counting instant $u$ of $N^{(r)}_{s}$.
+Hence, from an elementary calculus similar to (4.1), we have
+$X^{(r)}_{a}(1)-X^{(r)}_{a}(0)$ $=\int_{0}^{1}1(L^{(r)}(u-)\leq a-1)N^{(r)}_{e}(du)-\int_{0}^{1}1(L^{(r)}(u)\leq a)N^{(r)}_{s}(du).$
+Since $X^{(r)}_{a}(t)$ is stationary by (3.e) and bounded by $a$, this equation yields
+$\mathbb{E}\Big{[}\int_{0}^{1}1(L^{(r)}(u-)\leq a-1)N^{(r)}_{e}(du)\Big{]}=\mathbb{E}\Big{[}\int_{0}^{1}1(L^{(r)}(u)\leq a)N^{(r)}_{s}(du)\Big{]}.$
+Letting $a\to\infty$, the left-hand side of this equation converges to $\mathbb{E}[N^{(r)}_{e}(1)]=\alpha^{(r)}_{e}$, while it right-hand side converges to $\mathbb{E}[N^{(r)}_{s}(1)]=\alpha^{(r)}_{s}$.
+Hence, we have $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$.
+A.2 Proofs of Corollary 3.1
 We only verify (3.1) for $b_{1}\not=0$ because it is similarly proved for $b_{1}=0$. Since ${\mathcal{E}}^{(r)}_{1}=o(1)$ by Proposition 5.2, it follows from (4.54) and (3.10) that
-$\frac{1}{r}\mathbb{P}[L^{(r)}=0]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
-$=\frac{b_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1).$
+$\frac{1}{r}\mathbb{P}[L^{(r)}=0]$ $=\frac{b_{1}e^{\beta_{1}\ell_{1}}}{c_{1}(e^{\beta_{1}\ell_{1}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}+o(1)$
+$=\frac{b_{1}b_{2}e^{\beta_{1}\ell_{1}}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}+o(1).$
 On the other hand, from (3.8),
 $\lim_{r\downarrow 0}\mathbb{E}[rL^{(r)}]$
-$\quad=\frac{c_{1}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1-\beta_{1}\ell_{0})}{2b_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+\left(\ell_{0}+\frac{c_{2}\sigma_{2}^{2}}{2b_{2}}\right)\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}$
-$\quad=\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))}.$
+$\quad=\frac{c_{1}\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1-\beta_{1}\ell_{1})}{2b_{1}(e^{\beta_{1}\ell_{1}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}+\left(\ell_{1}+\frac{c_{2}\sigma_{2}^{2}}{2b_{2}}\right)\frac{c_{2}b_{1}}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}$
+$\quad=\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)+2b_{1}b_{2}\ell_{1}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1))}.$
 Hence, we have (3.1).
-A.2 Proof of Proposition 3.1
+A.3 Proof of Proposition 3.1
 We prove Proposition 3.1 only for $b_{1}\not=0$ because the case for $b_{1}=0$ can be similarly proved.
@@
 Taking this into account, we describe this model by the continuous time Markov chain with state space
-$S=\{(j,\ell);j=1,2,\ell=0,1,\ldots,\ell_{0}\}\cup\{(3,\ell);\ell\geq\ell_{0}+1\},$
+$S=\{(j,\ell);j=1,2,\ell=0,1,\ldots,\ell_{1}\}\cup\{(3,\ell);\ell\geq\ell_{1}+1\},$
 and its transition diagram is depicted in Figure A.1. Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$.
@@
 We partition the state space $S$ of $X_{\exp}(\cdot)$ into the following three disjoint subsets.
-$S_{j}=\{(j,\ell)\in S;\ell\in[0,\ell_{0}]\},\quad j=1,2,\qquad S_{3}=\{(3,\ell):\ell\geq\ell_{0}+1\}.$
+$S_{j}=\{(j,\ell)\in S;\ell\in[0,\ell_{1}]\},\quad j=1,2,\qquad S_{3}=\{(3,\ell):\ell\geq\ell_{1}+1\}.$
 By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{j,\ell}$ as
@@
 $\lambda_{1}p_{1,0}=c_{1}\mu p_{1,1},$ (A.1)
-$(\lambda_{1}+c_{1}\mu)p_{1,\ell}=\lambda_{1}p_{1,\ell-1}+c_{1}\mu p_{1,\ell+1}+\lambda_{2}p_{2,\ell-1},\quad\ell\in[1,\ell_{0}-1],$ (A.2)
-$(\lambda_{1}+c_{1}\mu)p_{1,\ell_{0}}=\lambda_{1}p_{1,\ell_{0}-1}+\lambda_{2}p_{2,\ell_{0}-1},$ (A.3)
+$(\lambda_{1}+c_{1}\mu)p_{1,\ell}=\lambda_{1}p_{1,\ell-1}+c_{1}\mu p_{1,\ell+1}+\lambda_{2}p_{2,\ell-1},\quad\ell\in[1,\ell_{1}-1],$ (A.2)
+$(\lambda_{1}+c_{1}\mu)p_{1,\ell_{1}}=\lambda_{1}p_{1,\ell_{1}-1}+\lambda_{2}p_{2,\ell_{1}-1},$ (A.3)
 and, for states in $S_{2}$,
 $\lambda_{2}p_{2,0}=c_{1}\mu p_{2,1},$ (A.4)
-$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{0}-1],$ (A.5)
-$(\lambda_{2}+c_{1}\mu)p_{2,\ell_{0}}=c_{2}\mu p_{3,\ell_{0}+1},$ (A.6)
-For states in $S_{3}$, it follows from the equation for $(3,\ell_{0}+1)$ and the local balance that
-$(\lambda_{2}+c_{2}\mu)p_{3,\ell_{0}+1}=\lambda_{1}p_{1,\ell_{0}}+\lambda_{2}p_{2,\ell_{0}}+c_{2}\mu p_{\ell_{0}+2},$ (A.7)
-$\lambda_{2}p_{3,\ell}=c_{2}\mu p_{3,\ell+1},\qquad\ell\geq\ell_{0}+1.$ (A.8)
+$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,\ell+1},\qquad\ell\in[1,\ell_{1}-1],$ (A.5)
+$(\lambda_{2}+c_{1}\mu)p_{2,\ell_{1}}=c_{2}\mu p_{3,\ell_{1}+1},$ (A.6)
+For states in $S_{3}$, it follows from the equation for $(3,\ell_{1}+1)$ and the local balance that
+$(\lambda_{2}+c_{2}\mu)p_{3,\ell_{1}+1}=\lambda_{1}p_{1,\ell_{1}}+\lambda_{2}p_{2,\ell_{1}}+c_{2}\mu p_{\ell_{1}+2},$ (A.7)
+$\lambda_{2}p_{3,\ell}=c_{2}\mu p_{3,\ell+1},\qquad\ell\geq\ell_{1}+1.$ (A.8)
 In what follows, we use the following notations.
@@
 We first note that, from (A.8),
-$p_{3,\ell}=\rho_{22}^{\ell-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad\ell\geq\ell_{0}+1.$ (A.9)
-Thus, $p_{3,\ell}$ for $\ell\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
-We next consider $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
+$p_{3,\ell}=\rho_{22}^{\ell-(\ell_{1}+1)}p_{3,\ell_{1}+1},\qquad\ell\geq\ell_{1}+1.$ (A.9)
+Thus, $p_{3,\ell}$ for $\ell\geq\ell_{1}+1$ is simply determined by $p_{3,\ell_{1}+1}$.
+We next consider $p_{2,\ell}$ for $\ell\in[0,\ell_{1}-1]$. From (A.4) and (A.5), we have
 $p_{2,0}=\frac{c_{1}\mu}{\lambda_{2}}p_{2,1}=\rho_{21}^{-1}p_{2,1},$ (A.10)
-$p_{2,\ell}=\gamma_{12}p_{2,\ell+1}=\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}},\quad\ell\in[1,\ell_{0}-1].$ (A.11)
-Thus, $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written by (A.4) as
-$p_{2,\ell}=\gamma_{12}^{-1}p_{2,\ell-1}=\gamma_{12}^{-\ell+1}p_{2,1}=\gamma_{12}^{-\ell+1}\rho_{21}p_{2,0},\quad\ell\in[1,\ell_{0}-1].$ (A.12)
-Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{j}$ for $j=1,2,3$.
+$p_{2,\ell}=\gamma_{12}p_{2,\ell+1}=\gamma_{12}^{\ell_{1}-\ell}p_{2,\ell_{1}},\quad\ell\in[1,\ell_{1}-1].$ (A.11)
+Thus, $p_{2,\ell}$ for $\ell\in[0,\ell_{1}-1]$ is determined by $p_{2,\ell_{1}}$. Furthermore, (A.11) can be written by (A.4) as
+$p_{2,\ell}=\gamma_{12}^{-1}p_{2,\ell-1}=\gamma_{12}^{-\ell+1}p_{2,1}=\gamma_{12}^{-\ell+1}\rho_{21}p_{2,0},\quad\ell\in[1,\ell_{1}-1].$ (A.12)
+Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{1}}$, $p_{2,\ell_{1}}$ and $p_{3,\ell_{1}+1}$ at the boundaries states of $S_{j}$ for $j=1,2,3$.
 We call them boundary probabilities.
-There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{0}}$ for large $\ell_{0}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}-1}=\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$ by (A.12) and (A.11), so we do not consider it.
+There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{1}}$ for large $\ell_{1}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}-1}=\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}$ by (A.12) and (A.11), so we do not consider it.
 We first note that, from (A.6),
-$p_{2,\ell_{0}}=\gamma_{22}p_{3,\ell_{0}+1}.$ (A.13)
-Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $\ell=\ell_{0}+1$, (A.7) can be written as
-$p_{3,\ell_{0}+1}=\rho_{12}p_{1,\ell_{0}}+\rho_{22}p_{2,\ell_{0}},$ (A.14)
-and, substituting $p_{3,\ell_{0}+1}=\gamma_{22}^{-1}p_{2,\ell_{0}}$ of (A.13) to this equation, we have
-$(\gamma_{22}^{-1}-\rho_{22})p_{2,\ell_{0}}=\rho_{12}p_{1,\ell_{0}}.$
+$p_{2,\ell_{1}}=\gamma_{22}p_{3,\ell_{1}+1}.$ (A.13)
+Since $c_{2}\mu p_{\ell_{1}+2}=\lambda_{2}p_{3,\ell_{1}+1}$ by (A.8) for $\ell=\ell_{1}+1$, (A.7) can be written as
+$p_{3,\ell_{1}+1}=\rho_{12}p_{1,\ell_{1}}+\rho_{22}p_{2,\ell_{1}},$ (A.14)
+and, substituting $p_{3,\ell_{1}+1}=\gamma_{22}^{-1}p_{2,\ell_{1}}$ of (A.13) to this equation, we have
+$(\gamma_{22}^{-1}-\rho_{22})p_{2,\ell_{1}}=\rho_{12}p_{1,\ell_{1}}.$
 Since $\gamma_{22}^{-1}-\rho_{22}=\frac{c_{1}\mu+\lambda_{2}}{c_{2}\mu}-\frac{\lambda_{2}}{c_{2}\mu}=\frac{c_{1}}{c_{2}}=\rho_{12}\rho_{11}^{-1}$, this equation yields
-$p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}.$ (A.15)
-Thus, the boundary probabilities $p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ are explicitly related by (A.13) and (A.15).
-We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{0}}$, but this is not easy.
-To resolve this problem, we consider $p_{1,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.2),
-$\lambda_{1}p_{1,\ell}-c_{1}\mu p_{1,\ell+1}=\lambda_{1}p_{1,\ell-1}-c_{1}\mu p_{1,\ell}+\lambda_{2}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1],$
+$p_{2,\ell_{1}}=\rho_{11}p_{1,\ell_{1}}.$ (A.15)
+Thus, the boundary probabilities $p_{1,\ell_{1}}$, $p_{2,\ell_{1}}$ and $p_{3,\ell_{1}+1}$ are explicitly related by (A.13) and (A.15).
+We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{1}}$, but this is not easy.
+To resolve this problem, we consider $p_{1,\ell}$ for $\ell\in[0,\ell_{1}-1]$. From (A.2),
+$\lambda_{1}p_{1,\ell}-c_{1}\mu p_{1,\ell+1}=\lambda_{1}p_{1,\ell-1}-c_{1}\mu p_{1,\ell}+\lambda_{2}p_{2,\ell-1},\qquad\ell\in[1,\ell_{1}-1],$
 then this equation can be written as
-$\rho_{11}p_{1,\ell}-p_{1,\ell+1}=\rho_{11}p_{1,\ell-1}-p_{1,\ell}+\rho_{21}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1].$
-Summing up both sides of this equation for $\ell\in[1,m]$ for $m\in[1,\ell_{0}-1]$, we have
-$\rho_{11}p_{1,m}-p_{1,m+1}=\rho_{21}\sum_{\ell=1}^{m}p_{2,\ell-1},\qquad m\in[0,\ell_{0}-1],$ (A.16)
+$\rho_{11}p_{1,\ell}-p_{1,\ell+1}=\rho_{11}p_{1,\ell-1}-p_{1,\ell}+\rho_{21}p_{2,\ell-1},\qquad\ell\in[1,\ell_{1}-1].$
+Summing up both sides of this equation for $\ell\in[1,m]$ for $m\in[1,\ell_{1}-1]$, we have
+$\rho_{11}p_{1,m}-p_{1,m+1}=\rho_{21}\sum_{\ell=1}^{m}p_{2,\ell-1},\qquad m\in[0,\ell_{1}-1],$ (A.16)
 because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $m=0$ by the convention that the empty sum vanishes, which we always take.
 Hence, applying (A.10) and (A.11) to (A.16), we have
-$\rho_{11}p_{1,m}-p_{1,m+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{\ell=1}^{m-1}\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}}\right)$
-$=\left(1+\rho_{21}\sum_{\ell=1}^{m-1}\gamma_{12}^{-(\ell-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
-$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[1,\ell_{0}-1].$
+$\rho_{11}p_{1,m}-p_{1,m+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{\ell=1}^{m-1}\gamma_{12}^{\ell_{1}-\ell}p_{2,\ell_{1}}\right)$
+$=\left(1+\rho_{21}\sum_{\ell=1}^{m-1}\gamma_{12}^{-(\ell-1)}\right)\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}$
+$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}},\qquad m\in[1,\ell_{1}-1].$
 Multiplying both sides by $\rho_{11}^{-(m+1)}$,
 $\rho_{11}^{-m}p_{1,m}-\rho_{11}^{-(m+1)}p_{1,m+1}$
-$\quad=\rho_{11}^{-(m+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(m\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[0,\ell_{0}-1].$
-Summing up both sides of this equation from $m=0$ to $m=k-1\leq\ell_{0}-1$, we have
-$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{m=0}^{k-1}\rho_{11}^{-(m+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{m=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(m-1)}-\rho_{11}^{-(m-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
-$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}}$
-$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}\rho_{11}^{-2}-\frac{\rho_{11}^{-k}-\rho_{11}^{-1}}{1-\rho_{11}}\right)\frac{\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}},\quad k\in[1,\ell_{0}].$
-Hence, multiplying both sides of this equation by $\rho_{11}^{k}$, we have
-$p_{1,k}$ $=\rho_{11}^{k}p_{1,0}-\frac{1-\rho_{11}^{k}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\frac{1-\rho_{11}^{k-1}}{1-\rho_{11}}\frac{\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
-$\quad-\frac{\gamma_{12}^{-(k-1)}-\rho_{11}^{k-1}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}},\qquad k\in[1,\ell_{0}].$ (A.17)
-In particular, for $k=\ell_{0}$, substituting $p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}$ of (A.15), we have
-$\rho_{11}^{\ell_{0}}\frac{p_{1,0}}{p_{1,\ell_{0}}}$ $=1+\frac{1-\rho_{11}^{\ell_{0}}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}\rho_{11}-\frac{1-\rho_{11}^{\ell_{0}-1}}{1-\rho_{11}}\frac{\rho_{11}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}$
-$\quad+\frac{\rho_{11}\gamma_{12}-\rho_{11}^{\ell_{0}}\gamma_{12}^{\ell_{0}}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}.$ (A.18)
-Thus, $p_{1,0}$ is explicitly related to $p_{1,\ell_{0}}$.
+$\quad=\rho_{11}^{-(m+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(m\geq 1)\right)\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}},\qquad m\in[0,\ell_{1}-1].$
+Summing up both sides of this equation from $m=0$ to $m=k-1\leq\ell_{1}-1$, we have
+$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{m=0}^{k-1}\rho_{11}^{-(m+1)}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\sum_{m=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(m-1)}-\rho_{11}^{-(m-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}$
+$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}}{1-\gamma_{12}}$
+$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}\rho_{11}^{-2}-\frac{\rho_{11}^{-k}-\rho_{11}^{-1}}{1-\rho_{11}}\right)\frac{\rho_{21}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}}{1-\gamma_{12}}$
+for $k\in[1,\ell_{1}]$. Hence, multiplying both sides of this equation by $\rho_{11}^{k}$, we have
+$p_{1,k}$ $=\rho_{11}^{k}p_{1,0}-\frac{1-\rho_{11}^{k}}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\frac{1-\rho_{11}^{k-1}}{1-\rho_{11}}\frac{\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}$
+$\quad-\frac{\gamma_{12}^{-(k-1)}-\rho_{11}^{k-1}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}},\qquad k\in[1,\ell_{1}].$ (A.17)
+In particular, for $k=\ell_{1}$, substituting $p_{2,\ell_{1}}=\rho_{11}p_{1,\ell_{1}}$ of (A.15), we have
+$\rho_{11}^{\ell_{1}}\frac{p_{1,0}}{p_{1,\ell_{1}}}$ $=1+\frac{1-\rho_{11}^{\ell_{1}}}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}\rho_{11}-\frac{1-\rho_{11}^{\ell_{1}-1}}{1-\rho_{11}}\frac{\rho_{11}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{1}}$
+$\quad+\frac{\rho_{11}\gamma_{12}-\rho_{11}^{\ell_{1}}\gamma_{12}^{\ell_{1}}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}.$ (A.18)
+Thus, $p_{1,0}$ is explicitly related to $p_{1,\ell_{1}}$.
 We now apply these computations to the $r$-th 2-level $M/M/1$ queue, and compute the limit of its scaled stationary distribution through moment generating functions.
@@
 Define the moment generating functions $g^{(r)}_{j}(\theta)$ for $j=1,2,3$ as
-$g^{(r)}_{j}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{j,k},\quad\theta\in\mathbb{R},j=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
+$g^{(r)}_{j}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}}e^{r\theta k}p^{(r)}_{j,k},\quad\theta\in\mathbb{R},j=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{1}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
 and define $g^{(r)}(\theta)=\sum_{j=1}^{3}g^{(r)}_{j}(\theta)$.
@@
 We first compute $g^{(r)}_{j}(\theta)$ for $j=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
-$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{\ell=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-\ell)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-\ell}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{0}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}},$ (A.19)
-$g^{(r)}_{3}(\theta)$ $=\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}.$ (A.20)
-On the other hand, computing $g^{(r)}_{1}(\theta)$ is a bit complicated. To ease it, we introduce the following functions.
-$f^{(r)}_{1}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}e^{r\theta k},\qquad f^{(r)}_{\rho_{1}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}\rho^{(r)}_{1})^{k},\qquad f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k},$
-then
-$f^{(r)}_{1}(\theta)=\frac{1-(e^{r\theta})^{\ell^{(r)}_{0}}}{1-e^{r\theta}}1(\theta\not=0)+\ell^{(r)}_{0}1(\theta=0),\qquad f^{(r)}_{\rho_{1}}(\theta)=\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-e^{r\theta}\rho^{(r)}_{1}},$
-$f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\frac{e^{r\theta\ell^{(r)}_{0}}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{-\ell^{(r)}_{0}+1}.$
-Using these function, we compute $g^{(r)}_{1}(\theta)$ from (A.2) as given below.
-$g^{(r)}_{1}(\theta)$ $=\sum_{k=0}^{\ell^{(r)}_{0}}(e^{r\theta}\rho^{(r)}_{1})^{k}p^{(r)}_{1,0}-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta k}-(e^{r\theta}\rho^{(r)}_{1})^{k}}{1-\rho^{(r)}_{1}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad+\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta(k-1)}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{1-\rho^{(r)}_{1}}\frac{e^{r\theta}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k-1}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell_{0}}$
-$=\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(\theta)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad+\left(f^{(r)}_{1}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$ (A.21)
+$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{1}r\theta}\sum_{\ell=1}^{\ell^{(r)}_{1}}e^{-(\ell^{(r)}_{1}-\ell)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-\ell}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{1}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{1}},$ (A.19)
+$g^{(r)}_{3}(\theta)$ $=\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{1}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{1}+1}.$ (A.20)
+On the other hand, we compute $g^{(r)}_{1}(\theta)$ from (A.3) as given below.
+$g^{(r)}_{1}(\theta)$ $=\sum_{k=0}^{\ell^{(r)}_{1}}(e^{r\theta}\rho^{(r)}_{1})^{k}p^{(r)}_{1,0}-\sum_{k=1}^{\ell^{(r)}_{1}}\frac{e^{r\theta k}-(e^{r\theta}\rho^{(r)}_{1})^{k}}{1-\rho^{(r)}_{1}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad+\sum_{k=1}^{\ell^{(r)}_{1}}\frac{e^{r\theta(k-1)}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{1-\rho^{(r)}_{1}}\frac{e^{r\theta}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad-\sum_{k=1}^{\ell^{(r)}_{1}}\frac{(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k-1}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell_{1}}.$ (A.21)
+To ease further compuations, we introduce the following functions.
+$f^{(r)}_{1}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}-1}e^{r\theta k},\qquad f^{(r)}_{\rho_{1}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}-1}(e^{r\theta}\rho^{(r)}_{1})^{k},\qquad f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}-1}(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k},$
+which are computed as
+$f^{(r)}_{1}(\theta)=\frac{1-(e^{r\theta})^{\ell^{(r)}_{1}}}{1-e^{r\theta}}1(\theta\not=0)+\ell^{(r)}_{1}1(\theta=0),\qquad f^{(r)}_{\rho_{1}}(\theta)=\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}}}{1-e^{r\theta}\rho^{(r)}_{1}},$
+$f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\frac{e^{r\theta\ell^{(r)}_{1}}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{e^{r\theta}-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{-\ell^{(r)}_{1}+1}.$
+Then, using these function, (A.3) becomes
+$g^{(r)}_{1}(\theta)$ $=\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(\theta)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad+\left(f^{(r)}_{1}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$ (A.22)
 We now compute the limits of $g^{(r)}_{j}(0)$ for $j=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
-$(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=O(\gamma_{12}^{\ell_{0}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.22)
-Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}$ decays exponentially fast as the function of $1/r$. On the other hand,
+$(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}=O(\gamma_{12}^{\ell_{1}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.23)
+Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}$ decays exponentially fast as the function of $1/r$. On the other hand,
 $\lim_{r\downarrow 0}(1-\rho^{(r)}_{\ell})/r=\beta_{\ell},\qquad\ell=1,2,$ (A.24)
-$\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{0}/r+o(r)}=e^{-\beta_{1}\ell_{0}},$ (A.24)
+$\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{1}/r+o(r)}=e^{-\beta_{1}\ell_{1}},$ (A.25)
 where recall that $\beta_{\ell}=b_{\ell}/c_{\ell}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.
@@
 For the sequence of the 2-level $M/M/1$ queue satisfying (3.b)–(3.e),
-$\lim_{r\downarrow 0}\frac{p^{(r)}_{2,\ell_{0}}}{p^{(r)}_{1,\ell_{0}}}=1,\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{3,\ell_{0}+1}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}},\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}.$ (A.25)
+$\lim_{r\downarrow 0}\frac{p^{(r)}_{2,\ell_{1}}}{p^{(r)}_{1,\ell_{1}}}=1,\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{3,\ell_{1}+1}}{p^{(r)}_{1,\ell_{1}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}},\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{1}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{1}}.$ (A.26)
 Proof.
@@
 For the third limit, we have, from (A.3) for the $r$-th system,
-$\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}$ $=\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{-\ell^{(r)}_{0}}\left(1+\frac{\rho^{(r)}_{1}\gamma^{(r)}_{12}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}\right).$
+$\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{1}}}$ $=\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{-\ell^{(r)}_{1}}\left(1+\frac{\rho^{(r)}_{1}\gamma^{(r)}_{12}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}\right).$
 Hence, applying (A.25), we have the third limit of (A.26). ∎
-We now tentatively assume that there exists finite $\widetilde{p}_{1,\ell_{0}}$ such that
-$\widetilde{p}_{1,\ell_{0}}=\lim_{r\downarrow 0}\frac{p^{(r)}_{1,\ell^{(r)}_{0}}}{r}.$ (A.26)
-The existence of this $\widetilde{p}_{1,\ell_{0}}$ will be verified at the final step (see (A.31)). Applying these asymptotic properties, it follows from (A.2) and (A.20) with help of (A.26) that
-$\lim_{r\downarrow 0}g_{2}^{(r)}(0)$ $=\lim_{r\downarrow 0}\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}}$
-$=\frac{c_{1}\mu+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{1,\ell^{(r)}_{0}}=0,$ (A.27)
-$\lim_{r\downarrow 0}g^{(r)}_{3}(0)$ $=\lim_{r\downarrow 0}\frac{1}{1-\rho^{(r)}_{2}}p^{(r)}_{3,\ell^{(r)}_{0}+1}=\frac{c_{2}}{b_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
-$=\beta_{2}^{-1}\gamma_{22}^{-1}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{2}^{-1}\widetilde{p}_{1,\ell_{0}}.$ (A.28)
+We now tentatively assume that there exists finite $\widetilde{p}_{1,\ell_{1}}$ such that
+$\widetilde{p}_{1,\ell_{1}}=\lim_{r\downarrow 0}\frac{p^{(r)}_{1,\ell^{(r)}_{1}}}{r}.$ (A.27)
+The existence of this $\widetilde{p}_{1,\ell_{1}}$ will be verified at the final step (see (A.32)). Applying these asymptotic properties, it follows from (A.3) and (A.20) with help of (A.27) that
+$\lim_{r\downarrow 0}g_{2}^{(r)}(0)$ $=\lim_{r\downarrow 0}\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}+\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{1-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{1}}$
+$=\frac{c_{1}\mu+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{2,\ell^{(r)}_{1}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{1,\ell^{(r)}_{1}}=0,$ (A.28)
+$\lim_{r\downarrow 0}g^{(r)}_{3}(0)$ $=\lim_{r\downarrow 0}\frac{1}{1-\rho^{(r)}_{2}}p^{(r)}_{3,\ell^{(r)}_{1}+1}=\frac{c_{2}}{b_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{1}+1}$
+$=\beta_{2}^{-1}\gamma_{22}^{-1}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{2,\ell^{(r)}_{1}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{2}^{-1}\widetilde{p}_{1,\ell_{1}}.$ (A.29)
 To compute the limit of $g^{(r)}_{1}(0)$, we first compute it from (A.3) as
-$g^{(r)}_{1}(0)$ $=\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad+\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}.$ (A.29)
+$g^{(r)}_{1}(0)$ $=\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad+\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}.$ (A.30)
 To proceed to further compute, we note that
-$\lim_{r\downarrow 0}rf^{(r)}_{1}(0)=\lim_{r\downarrow 0}r\ell^{(r)}_{0}=\ell_{0},$
-$\lim_{r\downarrow 0}rf^{(r)}_{\rho_{1}}(0)=\lim_{r\downarrow 0}r\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-\rho^{(r)}_{1}}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}}),$
-$\lim_{r\downarrow 0}f^{(r)}_{\gamma_{12}^{-1}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\gamma^{(r)}_{12}=\frac{c_{1}\mu}{\lambda_{2}}=\rho_{21}^{-1}.$
+$\lim_{r\downarrow 0}rf^{(r)}_{1}(0)=\lim_{r\downarrow 0}r\ell^{(r)}_{1}=\ell_{1},$
+$\lim_{r\downarrow 0}rf^{(r)}_{\rho_{1}}(0)=\lim_{r\downarrow 0}r\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}}{1-\rho^{(r)}_{1}}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{1}}),$
+$\lim_{r\downarrow 0}f^{(r)}_{\gamma_{12}^{-1}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}=\lim_{r\downarrow 0}\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{1-\gamma^{(r)}_{12}}\gamma^{(r)}_{12}=\frac{c_{1}\mu}{\lambda_{2}}=\rho_{21}^{-1}.$
 Then, we have, using Lemma A.1,
-$\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}})\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,0}$
-$\quad=\beta_{1}^{-1}\left(e^{\beta_{1}\ell_{0}}-1\right)\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\widetilde{p}_{1,\ell_{0}},$
-$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}=0,$
-$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}=0,$
+$\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{1}})\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,0}$
+$\quad=\beta_{1}^{-1}\left(e^{\beta_{1}\ell_{1}}-1\right)\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\widetilde{p}_{1,\ell_{1}},$
+$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}=0,$
+$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}=0,$
 and
-$\lim_{r\downarrow 0}\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
-$\quad=\lim_{r\downarrow 0}f^{(r)}_{\gamma^{-1}_{12}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}p^{(r)}_{2,\ell^{(r)}_{0}}=0.$
+$\lim_{r\downarrow 0}\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
+$\quad=\lim_{r\downarrow 0}f^{(r)}_{\gamma^{-1}_{12}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}p^{(r)}_{2,\ell^{(r)}_{1}}=0.$
 Hence, from (A.3)
-$\lim_{r\downarrow 0}g^{(r)}_{1}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{1}^{-1}(e^{\beta_{1}\ell_{0}}-1)\widetilde{p}_{1,\ell_{0}}.$ (A.30)
+$\lim_{r\downarrow 0}g^{(r)}_{1}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{1}^{-1}(e^{\beta_{1}\ell_{1}}-1)\widetilde{p}_{1,\ell_{1}}.$ (A.31)
 Hence, by (A.3) and (A.3), we have
-$1=\lim_{r\downarrow 0}g^{(r)}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\left(\beta_{2}^{-1}+\beta_{1}^{-1}(e^{\beta_{1}\ell_{0}}-1)\right)\widetilde{p}_{1,\ell_{0}}.$ (A.31)
-This proves that $\widetilde{p}_{1,\ell_{0}}$ indeed exists and is finite, and given by
-$\widetilde{p}_{1,\ell_{0}}=\frac{\lambda_{2}\beta_{1}\beta_{2}}{(\lambda_{1}+\lambda_{2})(\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1))}.$ (A.32)
-We are now in a final step to find $\widetilde{g}(\theta)\equiv\lim_{r\downarrow 0}g^{(r)}(\theta)$ for $\theta<0$.
-This can be obtained from (A.2), (A.2) and (A.20) in the following way.
-From (A.20) and Lemma A.1,
-$\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$ $=\lim_{r\downarrow 0}\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
-$=\lim_{r\downarrow 0}\frac{r}{1-(1+r\theta+o(r\theta))(1-r\beta_{2}+o(r))}e^{\ell^{(r)}_{0}\theta}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
-$=\frac{e^{\ell_{0}\theta}}{\beta_{2}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,\ell^{(r)}_{0}}$
-$=\frac{e^{\ell_{0}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},$
+$1=\lim_{r\downarrow 0}g^{(r)}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\left(\beta_{2}^{-1}+\beta_{1}^{-1}(e^{\beta_{1}\ell_{1}}-1)\right)\widetilde{p}_{1,\ell_{1}}.$ (A.32)
+This proves that $\widetilde{p}_{1,\ell_{1}}$ indeed exists and is finite, and given by
+$\widetilde{p}_{1,\ell_{1}}=\frac{\lambda_{2}\beta_{1}\beta_{2}}{(\lambda_{1}+\lambda_{2})(\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1))}.$ (A.33)
+We are now in a final step to find $\widetilde{g}(\theta)\equiv\lim_{r\downarrow 0}g^{(r)}(\theta)$ for $\theta<0$. This can be obtained in the following way. From (A.20) and Lemma A.1,
+$\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$ $=\lim_{r\downarrow 0}\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{1}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{1}+1}$
+$=\lim_{r\downarrow 0}\frac{r}{1-(1+r\theta+o(r\theta))(1-r\beta_{2}+o(r))}e^{\ell^{(r)}_{1}\theta}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{1}+1}$
+$=\frac{e^{\ell_{1}\theta}}{\beta_{2}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,\ell^{(r)}_{1}}$
+$=\frac{e^{\ell_{1}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)},$
 while, from (A.3) and Lemma A.1,
-$\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)$ $=\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\lim_{r\downarrow 0}\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}+1}}{1-e^{r\theta}\rho^{(r)}_{1}}p^{(r)}_{1,0}$
-$=\frac{1-e^{\ell_{0}(\theta-\beta_{1})}}{\beta_{1}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}\widetilde{p}_{1,\ell_{0}}$
-$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\ell_{0}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}$
-Thus, we have, for $\theta\leq 0$ satisfying $\theta\not=\beta_{1}$,
+$\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)$ $=\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}=\lim_{r\downarrow 0}\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}+1}}{1-e^{r\theta}\rho^{(r)}_{1}}p^{(r)}_{1,0}$
+$=\frac{1-e^{\ell_{1}(\theta-\beta_{1})}}{\beta_{1}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{1}}\widetilde{p}_{1,\ell_{1}}$
+$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\ell_{1}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}.$
+Further, $\lim_{r\downarrow 0}g^{(r)}_{2}(\theta)=0$ by (A.3). Hence, we have, for $\theta\leq 0$ satisfying $\theta\not=\beta_{1}$,
 $\lim_{r\downarrow 0}g^{(r)}(\theta)$ $=\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)+\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$
-$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\ell_{0}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}+\frac{e^{\ell^{(r)}_{0}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}.$
+$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\ell_{1}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}+\frac{e^{\ell^{(r)}_{1}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}.$
 This completes the proof of Proposition 3.1.
-A.3 Proofs of Lemma 4.4
+A.4 Proofs of Lemma 4.4
 We first prove (4.27) and (4.28). Apply $f(z,{\boldsymbol{x}})=x_{1}^{k+1}$ to (4.1), then
-$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
-$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
+$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{1}]$
+$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{1}],$
 where (4.1) is applicable since the right-hand side of this equation is finite. Hence, (4.27) is obtained. Similarly, we have (4.28) applying $f(z,{\boldsymbol{x}})=x_{2}^{k+1}$ to (4.1) because
@@
 We omit the proof for $v=s$ because it is similar to the case of $v=e$ except for (4.27) to be used instead of (4.28).
-A.4 Proofs of Lemmas 4.7 and 4.8
+A.5 Proofs of Lemmas 4.7 and 4.8
 Proof of Lemma 4.7.
@@
 Applying $f(x)=e^{-x}$ to this inequality, we have $|e^{-x}-1+x|\leq\frac{1}{2}|x|^{2}e^{|x|}$. Hence, for each fixed $\theta\in\mathbb{R}$, as $r\downarrow 0$
-$e^{r\theta\ell^{(r)}_{0}}=e^{\theta(\ell_{0}+o(r))}=e^{\theta\ell_{0}}(1+\theta o(r))+o(\theta o(r))=e^{\theta\ell_{0}}(1+o(\theta r)),$
-$e^{r\theta(\ell^{(r)}_{0}+1)}=e^{\theta r\ell^{(r)}_{0}}e^{r\theta}=e^{\theta\ell_{0}}(1+o(\theta r))(1+\theta r+o((\theta r)^{2}))=e^{\theta\ell_{0}}(1+\theta r+o(\theta r)).$
+$e^{r\theta\ell^{(r)}_{1}}=e^{\theta(\ell_{1}+o(r))}=e^{\theta\ell_{1}}(1+\theta o(r))+o(\theta o(r))=e^{\theta\ell_{1}}(1+o(\theta r)),$
+$e^{r\theta(\ell^{(r)}_{1}+1)}=e^{\theta r\ell^{(r)}_{1}}e^{r\theta}=e^{\theta\ell_{1}}(1+o(\theta r))(1+\theta r+o((\theta r)^{2}))=e^{\theta\ell_{1}}(1+\theta r+o(\theta r)).$
 Thus, (4.38) is obtained. (4.39) and (4.7) are immediate from these asymptotic expansions. ∎
@@
 Since $\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}$ and $\eta^{(r)}_{i}(r\theta)\widehat{R}^{(r)}_{e}$ are uniformly bounded in $r<a$ for a sufficiently small $a>0$ and vanish as $r\downarrow 0$, it follows from Lemma 4.2 and (A.5) that
-$e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}=1-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}+\frac{1}{2}(\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-})^{2}+O\big{(}|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
+$e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}$
+$\quad=1-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}+\frac{1}{2}(\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-})^{2}+O\big{(}|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
 $\quad=1-\left(-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}(r\theta)^{2}+o((r\theta)^{2})\right)\widehat{R}^{(r)}_{s-}$
@@
 where recall that $S_{3}(\widehat{R}^{(r)}_{s-})=\widehat{R}^{(r)}_{s-}+(\widehat{R}^{(r)}_{s-})^{2}+(\widehat{R}^{(r)}_{s-})^{3}$. This proves (4.8), and (4.8) and (4.8) are similarly proved. ∎
-A.5 Proofs of Lemmas 5.1 and 5.2
-We first prove Lemma 5.1, namely, that (5.3) implies (5.4). To prove this, we use the BAR (4.1) for $f(z,x_{1},x_{2})=(x_{1}\wedge 1/r)(x_{2}\wedge 1/r)1(z=\ell^{(r)}_{0})$, then we have
+A.6 Proof of Lemma 5.1
+We first assume $\lambda_{1}\not=\lambda_{2}$, and prove (5.3) for $i=1$. Consider the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$. From (4.2), we have
+$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (A.38)
+zz Applying (4.38), (4.8) and Corollary 4.4 to this equation, we have
+$D^{(r)}(\theta,\theta)$ $=\alpha^{(r)}_{e}e^{\theta\ell_{1}}(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}\big{[}(r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}$
+$\quad+2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}\lambda^{(r)}_{2}-(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}\lambda^{(r)}_{1}\}\widehat{R}^{(r)}_{e}$
+$\quad-2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2})^{2}-(\lambda^{(r)}_{1})^{2}\}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{1})\big{]}+o((r\theta)^{2}),$
+$=\alpha^{(r)}_{e}e^{\theta\ell_{1}}\big{\{}r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})\big{]}+(r\theta)^{2}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})\big{\}}$
+$\quad+o((r\theta)^{2}).$
+Since ${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}$ is uniformly bounded in $r\in(0,1]$, this implies that
+$D^{(r)}(\theta,\theta)$ $=r\theta\alpha^{(r)}_{e}e^{\theta\ell_{1}}(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})\big{]}+O((r\theta)^{2}).$ (A.39)
+Substitute this formula into the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$, which is given by
+$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta)$
+$\quad=c_{1}\left(\mu r\theta-\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o((\theta r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta).$ (A.40)
+Then, picking up the terms of order $r\theta$, we have
+$c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{\theta\ell_{1}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})]=O(r\theta).$
+Since $\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=O(r)$ by Corollary 4.2 and Lemma 4.9, this yields
+$(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})]=O(r).$ (A.41)
+Since $\lambda_{1}\not=\lambda_{2}$, (A.41) implies (5.3) for $i=1$. To prove (5.3) for $i=2$, we derive the following formula by Schwartz’s inequality.
+$\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{1}]\big{\}}^{2}=\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3/2}(\widehat{R}^{(r)}_{e})^{1/2}1^{2}(L^{(r)}(0)=\ell^{(r)}_{1})]\big{\}}^{2}$
+$\quad\leq\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{1})]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (A.42)
+Since $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{1})]$ is uniformly bounded in $r$ by Corollary 4.4, (5.3) for $i=1$ implies that for $i=2$. Thus, (5.3) is proved.
+We next prove that (5.4) implies (5.3). For this proof, we use the BAR (4.1) for $f(z,x_{1},x_{2})=(x_{1}\wedge 1/r)(x_{2}\wedge 1/r)1(z=\ell^{(r)}_{0})$, then we have
 $\mathbb{E}[R^{(r)}_{e}1(R^{(r)}_{e}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]+\mathbb{E}[R^{(r)}_{s}1(R^{(r)}_{s}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]$
@@
 $\big{(}\mathbb{E}[R^{(r)}_{v}1(R^{(r)}_{v}\leq 1/r,L^{(r)}=\ell^{(r)}_{0})]\big{)}^{2}\leq\mathbb{E}[(R^{(r)}_{v})^{2}]\mathbb{P}[L^{(r)}=\ell^{(r)}_{0}],\qquad v=e,s.$
-Hence, (5.4) follows from (A.5) because $\mathbb{E}[\widehat{T}^{(r)}_{s}]$ converges to $\mu^{-1}$ as $r\downarrow 0$. Thus, Lemma 5.1 is proved.
-We next prove Lemma 5.2. By Schwartz’s inequality,
-$\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0}]\big{\}}^{2}=\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3/2}(\widehat{R}^{(r)}_{e})^{1/2}1^{2}(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}^{2}$
-$\quad\leq\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})].$
-Since $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded in $r$ by Corollary 4.4, (5.4) implies (5.5). This complete the proof of Lemma 5.2.
+Hence, we have (5.3) for $i=1$. It for $i=2$ follows from (A.6) as we have argued.
+A.7 Proof of Lemma 5.2
+In the proof of Proposition 5.2, under the assumption $\lambda_{1}\not=\lambda_{2}$, (5.2) is obtained applying (5.3) to (5.2).
+Obviously, if (5.3) holds, then this proof is also valid without any extra assumption including $\lambda_{1}\not=\lambda_{2}$.
+Hence, the lemma is proved.
 Acknowledgement